A Complete Bibliography of Publications in SysAdmin:

The Journal for UNIX Systems Administrators

Nelson H. F. Beebe Center for Scientific Computing

University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

05 July 2016 Version 1.63

Title word cross-reference

SA = f(x) [Gar97]. /etc/hosts [Ess98a]. /proc [Kla98]. 1 [Smi01b]. 100Mb [CY98]. 100Mb/s [CY98]. 101 [Hon98a, Sei00a]. 1998 [Nol98]. 2.2 [BB99]. 2.5.1 [Dia00]. 2.X [BR99]. 2000 [Wit01]. 3 [Koe98]. 39 [Kie99b]. 4.0 [Joh97, Kni92, Tis98].

6bone [LA01]. 7.0 [Dul00a]. 8 [Gal02c, LA01, Wen01]. 8.9 [Sch99b]. 8.x [OY01]. 9.x [DeJ02]. 90s [Hun92d]. 95 [Don95b]. 98 [Gol01]. = [Gom94]. ABCs [Bec99]. Abyss [WF01]. Access [Dav00e, DB02, EK00, Hel99c, Hon98b, Isa94b, Jon02, LB96, Sal98, Smi02a, Vin00, Wad96]. According [McC00l]. Accounting [Haz99b, Man01a, Rie94]. Accounts [Bel98, Sal97b]. Achieving [EK00, Oli01]. ACLs [Oli01]. Acquiring [Tan98a]. Across [Ram01, Bak95, BL99]. Activity [Gen93a, Har93c, Wyn97]. Acts [Sch00g]. Add [Bro97]. Add-on [Bro97]. Adding [Fla97]. Additional [McK00a]. Address [McC00g]. Addresses [Rez94d]. Admin [Bun96, Plo97, Rau99, Vog99, Ano96-47, Ano98-68, Ano98-69, Bar98m, Bec00, Isa96a, Plo98]. Administering [LE97, Har93g, Nap02]. Administration [Ben98, De 97a, Dij00, FPA99, Gal02b, Gal02e, Gar97, Gen94c, Gen94d, Gen95, Hun92c, Hun93a, Hun93b, Kol99, Rug00, Sar02, Tar96]. Administrative [Bye99]. Administrator [App00, Dyk96, Hil98, Hun92a, Luc00b, Whi96, WC97]. Administrators [Man01b, Pra01b, Pra01c]. Admins [Che97]. Advantage [Wyn94]. Advertiser [Ano92a, Ano92b, Ano92c, Ano92d, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano95a, Ano95b, Ano95c, Ano95d, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano00m, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano02a, Ano02b]. Advertiser [Ano02c]. Advisory [She98]. Affinity [Fer01b]. afind [GN01a]. Age [Har98]. Agent [Vog99, dRV01]. Aging [McK98a]. agrep [GN01a]. AIT [Swi98]. AIX [Bol00, Can99, Car01b, Cou92, Cou93a, Cou93b, Cou93c, Daw01, Gen94a, Hun92b, Mar01a, Mor95, Pet95, Ric94a, Ric95, Ric96b, San98, Skl01, Woo01]. Alarms [AS96]. Alias [Coo00]. Aliases [Pet01]. Alike [Che97]. alp [Jac99a]. Alternate [Mar01a, McK98b]. Alternative [Bar99a]. Alternatives [McC00i, Vad98]. Amanda [Hal02, Smi02b]. Analysis [AF00, Gag99b]. Analyzing [Ber98b]. Answers [McK99b, McK99c, McK00b, McK00c, McK00d, McK00a, McK00e, McK00f, McK00g, McK00h, McK00i, McK00j, McK00k, MR00, MR01a, MR01b, MR01c, MR01d, MR01e, MR01f, MR01g, MR01h, Ric01a, Ric01b, Ric01c, Ric01d, Ric01e, Ric02a, Ric02b, Ric02c, Ric02d, Ric02e, Sat92a, Sat92b, Sat92c, Sat93a, Sat93b, Sat93c, Sat93d, Sat93e, Sat94a, Sat94b, Sat94c, Sat94d, Sat94e, Sat94f, Sat95b, Sat95c, Sat95d, Sat95e, Sat95f, Sat96a, Sat96b, Sat96c, Sat96d, Sat96e, Sat96f, Sat96g, Sat96h, Sat96i, Sat96j, Sat97a, Sat97b, Sat97c, Sat97d, Sat97e, Sat97f, Sat97g, Sat97h, Sat97i, Sat97j, Sat98a, Sat98b, Sat98c, Sat98d, Sat98e, Sat98f, Sat98g, Sat98h, Sat98i, Sat98j, Sat98k, Sat99a, Sat99b, Sat99c, Sat99d, Sat99e, Sat99f, Sat99g, Sat99h, Sat99i]. Antivirus [Par01]. Any [Hil93]. Anyway [Sch01f]. Apache [Bal99, Bus99d, Den99b, GW00, Gou01, Ram01, Rib02, Vog00c]. Application [Bus01, GW00, Jac99a, Man01a]. Applications [Cay94a, Gag99a, Gag02, Har96, Joh96a, Nak01, Rez95d, RB01b]. Approach [FPA99, Mic99, Nas98a, Nas98b]. Approaches [Tee00]. Appropriate [Ron01]. archie [Bus94a]. Architecture [Pur01]. Archives [Vel95a]. Archiving [NS94]. Area [EK00, Sch00d, Sul00]. ARP [McC00l]. Art [BL00, Don98a, KP99, PK99, Vog01]. Artecon [van97a]. ASCII [Pol01]. Assessing [Ber98a]. Assignment [JK94]. Associated [Wat01b]. Assorted [Don96a]. Asynchronous [Har94a, Vel98a]. ATM [Rus96]. Attaching [Joh96c]. Attack [Hel01a, KP99, PK99]. Attacks [O’B96]. Attributes [Pet95]. Auditing [May97, WM00]. Authenticating [Hel98a]. Authentication [Bal99, Sei00d]. Authorization [Fre02]. Auto [Mor96b]. Auto-Locking [Mor96b]. Automated [Cal01, Dyk96, Gom94, Mal01, McD95, Tis00, VW00, Win99]. Automatic [Isa93c, Mar01b]. Automatically [Dav00a, Ho01, San93a]. Automating [Che96a, Con97a, Dav00b, Don96b, GS00, Hil00, KR94, McC98, Moo97, Sch99g, Tus95, Wyn97]. Automounted [All96]. Automounter [Bak95]. Availability [Ano92q, Ano92r, Ano92s, Ano93-27, Ano93-28, Ano93-29, Ano93-30, Ano93-31, Ano93-32, Ano94-31, Ano94-33, Ano94-32, Ano97y, Ano98-37, Ano98-38, Ano98-39, Ano98-40, Ano98-41, Ano98-42, Ano98-43, Ano98-44, Ano98-45, Ano98-46, Ano98-47, Ano98-48, Ano99-37, Ano99-38, Ano99-39, Ano99-40, Ano99-41, Ano99-42, Ano99-43, Ano99-44, Ano99-45, Ano99-46, Ano99-47, Ano99-48, Ano00-28, Ano00-29, Ano00-30, Ano00-31, Ano00-32, Ano00-33, Ano00-34, Ano00-35, Ano00-36, Ano00-37, Ano00-38, Ano00-39, Ano01u, Ano01v, Ano01w, Ano01x, Ano01y, Ano01z, Ano01-27, Ano01-28, Ano01-29, Ano01-30, Ano01-31, BN01, Jac99b, Koe98, Reh01a, Smi01b]. Available [Don99b]. Away [Ols00]. awk [Naz01].

Babysitting [Fel97a]. Back [BC99, Coo01]. Background [Zol92a]. Backing [Jon98b]. Backup [AS93, Bea96, Cas01a, Gil00, Goo98, Goo99, Hal02, Hel97e, Hel98d, Hu98a, Jon99, Man01b, Nas97b, Pet94, Pre96, Pre98, Pre01a, Pre01e, Pre01h, Sch02a, Sch00c, Str02, Mor95]. Backups [Abl93a, Ane96, Bay01, Che96a, Dav00a, Dav00b, HD97, Pre00b, Pre01f, Ric98, Smi02a, Wak02]. Bad [Dul94c, McC00c, Rea93]. Balancing [Bye99, Gou01]. Bandwidth [Hel97a]. Bar [Sch01a]. Based [Ali99, Bro97, Hel00h, Jon99, Kin95, Lib02, Lym02, Mac97, McC00b, McK01, Rau99, Sch95, Sha97, Smi98, SN95, Sto01, Vel94c, Vin00, Wol00, GW00, Haz97, Ric97, Don96f, Mag96]. Bash [Sch02b]. Basic [Wyn97]. Basics [Bal99, Bri00, Bus99a, Gar98a, Hel98c]. Batch [Dul96b]. BBWARN [Cro98]. Becoming [Hil98]. Been [Hug01]. Begin [Jos01]. Begin/Finish [Jos01]. Best [Gal02e]. Between [Bla99, Joh96a, Liu96, Dul96a, Joh95, Pan94, SL95a, Wei92b]. Beware [Abl92a]. Beyond [Dul96b, Kra00, Rea93]. Big [Mac97, Cro98, Har93a, Kha01]. BigAdmin [WT00]. Bin [Sto01]. Binaries [Rez95b]. Binary [Isa96b]. BIND [Con99, DeJ02]. Black [Sch02b]. Blocks [Pre01g]. BNU [Har93g]. Book [Zin97a, Zin97b, Zin97c]. Books [Zin92a, Zin92b, Zin92c, Zin93a, Zin93b, Zin93c, Zin93d, Zin93e, Zin93f, Zin94a, Zin94b, Zin94c, Zin94d, Zin94e, Zin94f, Zin95a, Zin95b, Zin95c, Zin95d, Zin95e, Zin95f, Zin96a, Zin96b, Zin96c, Zin96d, Zin96e, Zin96f, Zin96g, Zin96h, Zin96i, Zin96j, Zin96k, Zin96l, Zin97d, Zin97e, Zin97f, Zin97g, Zin97h, Zin97i, Zin97j, Zin98a, Zin98b, Zin98c, Zin98d, Zin98e, Zin98f, Zin98g, Zin98h, Zin98i, Zin98j, Zin98k, Zin98l, Zin99a, Zin99b, Zin99c, Zin99d, Zin99e, Zin99f, Zin99g, Zin99h, Zin99i, Zin99j, Zin99k, Zin99l, Zin00a, Zin00b, Zin00c, Zin00d, Zin00e, Zin00f, Zin00g, Zin00h, Zin00i, Zin00j, Zin00k, Zin00l, Zin00m, Zin01a, Zin01b, Zin01c, Zin01d, Zin01e, Zin01f, Zin01g, Zin01h, Zin01i, Zin01j, Zin01k, Zin01l]. Boot [McK98b, Vel94b, Win99]. Boot-Disk [Vel94b]. Booting [Gol01, Smi01b]. Box [Sch02b]. Brief [Gar98b]. Brother [Mac97, Cro98, Har93a, Kha01]. Browser [CC01, Don96h]. BSD [Bro95, Lee95a, O’G01]. BSDs [Luc99]. Budget [Pre01a]. Bug [Kie99a]. Build [McC00a, Mey01]. Building [Don96c, Don96d, Don00, Fel96a, Fel98, GW00, Pre01b, Pre01c, Pre01g, Sch00c, vdW00, Zol93a]. Built [Kom00a, Wes01]. Built-In [Kom00a]. Bus [Swe00b]. Bust [Zol92c]. Buy [Don99c]. Buyer [Bar98m]. Buying [And01].

C [New02, Rez92a, Sch93]. C2 [Har93a, Lam99b, WM00]. Cache [Won02]. Caching [McC99c]. Calendar [Isa95]. Call [Ano92e, Ano92f, Ano92g, Ano92h, Ano93g, Ano93h, Ano93i, Ano93j, Ano93k, Ano93l, Ano94g, Ano94h, Ano94i, Ano94j, Ano94k, Ano94l, Ano95e, Ano95f, Ano95g, Ano95h, Ano96k, Ano96l, Ano96m, Ano96n, Ano96o, Ano96p, Ano96r, Ano96q, Ano96s, Ano96t, Ano97i, Ano97j, Ano97k, Ano97l, Ano97m, Ano97n, Ano97o, Ano97p, Ano98m, Ano98n, Ano98o, Ano98p, Ano98q, Ano98r, Ano98s, Ano98t, Ano98u, Ano98v, Ano98w, Ano98x, Ano99m, Ano99n, Ano99o, Ano99p, Ano99q, Ano99r, Ano99s, Ano99t, Ano99u, Ano99v, Ano99w, Ano99x, Ano00n, Ano00o, Ano00p, Ano00q, Ano00r, Ano00s, Ano00t, Ano00u, Ano00v, Ano00w, Ano00x, Ano00y, Ano00z, Ano01l, Ano01m, Ano01n, Ano01o, Ano01p, Ano01q, Ano01r, Ano01s, Ano02d, Ano02e, Ano02f]. Can [Kle97, Lam99b]. Capabilities [Bro92b]. Capacity [Vel95a, Zwi98, Zwi99c]. Cards [Gar96]. Career [Ano96u, Ano96v, Ano96w, Ano96x, Ano96y, Ano96z, Ano96-27, Ano96-28, Ano97q, Ano97r, Ano97s, Ano97t, Ano97u, Ano97v, Ano97w, Ano97x, Ano98y, Ano98z, Ano98-27, Ano98-28, Ano98-29, Ano98-30, Ano98-31, Ano98-32, Ano98-33, Ano98-34, Ano98-35, Ano98-36, Ano99y, Ano99z, Ano99-27, Ano99-28, Ano99-29, Ano99-30, Ano99-31, Ano99-32, Ano99-33, Ano99-34, Ano99-35, Ano99-36, Ano00-27, Ano01t]. Careers [Dij00]. Case [Kie99b, Soh01, Wei92b]. CD [Car97, Kuo98, Smi02a]. CD-ROM [Car97]. CD-ROMs [Kuo98]. Centaur [LMBMR98]. Center [Kin01, McC00h]. Centralize [Hel01d]. Centralized [Cas01a, Par01, Wei98]. Centralizing [Agu01]. centric [De 97b]. Certificate [Tan98a]. Certification [Dul00b]. Cfengine [McC98]. CGI [End96, LB96, WB01]. Chains [BB99]. Change [Bru94]. Channel [Nel99, Pre00c, Pre01d, Sch98]. Character [Sch97a]. Cheapskates [Sch02c]. Check [Cla94, Mer98b]. checkcron [Isa93a]. Checker [Kin94]. Checking [Hel99a, Isa93a, Mer98a, Naz01, Pra01a, Rez94a]. Choices [Bro93, Hon98a]. CIFS [All98]. Cisco [Fre02, Lin99, McC99e, Sam01, Tis01, Zha98]. Class [Cla01]. Cleaning [San93a]. Client [May96]. Client/Server [May96]. Clients [Jon98b, Kos96]. Clone [Kni92]. Cloning [Daw01]. closeit [Eve95]. Closer [McC00f]. Clue [Dav99b]. Cluster [Gou01, Haz97]. Clusters [De 98a, Fer01a]. CODA [Lym01d, Lym01a]. Code [Ano92q, Ano92r, Ano92s, Ano93-27, Ano93-28, Ano93-29, Ano93-30, Ano93-31, Ano93-32, Ano94-31, Ano94-33, Ano94-29, Ano94-30, Ano94-32, Ano97y, Ano98-37, Ano98-38, Ano98-39, Ano98-40, Ano98-41, Ano98-42, Ano98-43, Ano98-44, Ano98-45, Ano98-46, Ano98-47, Ano98-48, Ano99-37, Ano99-38, Ano99-39, Ano99-40, Ano99-41, Ano99-42, Ano99-43, Ano99-44, Ano99-45, Ano99-46, Ano99-47, Ano99-48, Ano00-28, Ano00-29, Ano00-30, Ano00-31, Ano00-32, Ano00-33, Ano00-34, Ano00-35, Ano00-36, Ano00-37, Ano00-38, Ano00-39, Ano01u, Ano01v, Ano01w, Ano01x, Ano01y, Ano01z, Ano01-27, Ano01-28, Ano01-29, Ano01-30, Ano01-31, Sch01a]. Coherent [Kni92]. Cold [Wat01a]. Collide [Dav98a]. Colon [Cou92]. Color [Hof97]. Combining [WT00]. Come [Har92e]. Coming [Har98]. Command [Pru02, Sch00a]. Command-Line [Pru02, Sch00a]. Commands [Abl93a, Wee99, Wei01a]. Commerce [Kla00]. Common [Bar99j]. Communicating [Dul94a, Sch01a]. Communication [Joh96a]. Communications [Gol98, Hel97d]. Community [WT00, Zol92b]. Community-Style [Zol92b]. Companion [Gal01b, Gal02b, Gal02c, Gal02a]. Company [Zol93d]. Comparing [Vel95a]. Comparison [Iba97, Swi98]. Compiling [Sch99h]. Complete [Kin01]. Complex [Goo99]. Comprehensive [Mic99]. Computer [Jam94, Kin95]. Computers [Hal01, Nie00]. Computing [Haz97, SL95b]. Concept [BB99]. Conference [Nol98]. Configs [McC98]. Configurable [SN95]. Configuration [Ess96, Gar98a, Goo98, Goo99, Sat95a, Tus95, dRV01]. Configure [Lin99]. Configuring [Coo96, DB02, Gan94b, Har93b, Les00, Mes94, Reh01a, Smi02b, Zha98]. Connect [Bun97]. Connected [Bla99]. Connection [Dav99b, Fel94, Fel96b, Hel96a, Liu96]. Connections [Ano97-27, Her02, Muh97b]. Connectivity [Hun92d, Joh97]. Considerations [McC99a, McC99d, McK99a]. Console [Bou99]. Consolidating [Hil96c]. Construction [Har95a]. Consumption [Hel01c]. Content [Ols01a, RK97]. Contents [Ano96-47]. Context [Sch01c]. Continued [McC99c]. Control [Dav00e, Dil02, Har93e, Hu98b, Plo98, Sal98, Smi98, Vin00]. Controlled [Cox94]. Controlling [Con97b, RK97, War99]. Controls [Har92c, Har01a, Har01b]. Convergence [Hel00a, Hel00b, Hel00c]. Converging [Ano00-40]. Converting [SL95a, SB96]. Cookies [Hel98e]. Coordinating [Rez95a]. Copies [Vel95b]. Copying [Sch94, Sch00e]. Core [Har92e]. Corner [Gal01c, Gal01d, Gal02f, Gal02e, Gal02d].

Corporate [Ris00]. Correction

[Ano97z, Ano97-27]. Cost [Liu96, Vad98]. Costs [Won02]. CPU [Ber98a]. Cracker [Ric99]. Crackers [Ols00]. crash [Kon93]. Crawl [BC99]. Create [Kuo98]. Created [Ped02]. Creating [Bel98, Bro95, Cox94, Don96e, End96, End00, Hal97, Har95d, Hil95, Mel96, Orw01, Raa01, Rez95b, Sch02b, Tes96, Wat01a]. cron [Abl92b, NS94, Rez93e]. crontab [Rez93e]. Crontabs [Hel01d]. Cross [Agu00]. Crossword [Ano98-49, Ano98-50, Ano98-51, Ano98-52, Ano98-53, Ano98-68, Ano98-69]. Crypto [Sei00a]. Cryptography [WC97]. CTAR [Zol93d]. Cure [Bro92a]. Customer [Swe99]. Customizing [Bro96, Isa95, Mag97]. CVS [Kie99b].

Daemon [Sol02]. Dance [Nas97a]. Dangers [Aki01, Mor96a]. Data [Cer01b, Cou93a, EK00, Fel97b, Hag01, Hel00c, Lym01c, Plo00, SL95a]. Database [Che96a, Con97a, LE97, Gen94c, Gen94d, Gen95, McK99a, Pre96, Pru02, AS96, Tes96, Vog00c]. Databases [Lib02, Mob98]. Date [Sch95]. Date/Time [Sch95]. Dates [MS99]. Day [Rez93g, Sch00h]. DBA [Tay99]. DBAs [BL99]. De-Spamming [Sch99b]. Debugger [Har92d]. de}code [Bur01a]. Deep [Sch00e]. Defender [Coo00]. Defending [PK99]. Defense [Hel00h]. Definition [Rez92b]. Defunct [DS00]. Dejan [Ano97-27]. Deleted [BL02]. Deleting [Ped02]. Delivering [Zha00]. Delivery [D’A96]. Dependent [Jay98]. Deploying [End00, Wid00]. Description [Tho95]. Design [Bou99, Kra99, McC99a, McC99d]. Designing [BF99, Jos00]. Desk [Mel96]. Desktop [Bar98m, Kni92]. Desktops [LMBMR98]. Detailed [Sch99f]. Detecting [Dav96, End98]. Detection [Ano01-32, Cha01, Don98a, Dub01, Hor93b, Iba97, Kue01, McC99d, McC00b, Wes00a]. Determining [Hel96a]. Develop [Hil96b]. Developing [Sch01b, Sol02]. Development [Bau96, Wil01b]. Device [Don97d, Smi01b]. Devices [Joh96c, McK98b, Ris00]. Devouring [Leo01]. DHCP [Ben98, Sat95a, VW00]. Diagnosing [RL98]. Diagnostic [BR99, Dul94b]. Diagramming [Bic02]. Dial [Dia00, Kue00]. Dial-Up [Dia00, Kue00]. Did [Abl93c, And01, Har92e]. Difference [Sch00i]. Digital [Che96b]. Directories [San93a, Sch94]. Directory [Bri00, Bru94, De 97a, Gut96, Zol93c]. Dirty [DW01]. Disabled [BB01]. Disaster [BF99, Cer01a, Dea92, Hil97, Ric98]. Disconnectable [Lym01a]. Discussion [Mag96]. Disk [Gar00, Lei96, Mar01a, Mil00a, Rez95c, Rya94, Vel94b, Wil93, Win99]. Disks [Mur01]. Dispatcher [AS96, Wil01a]. Disruptions [Sal01]. Distributed [Cha01, Cox95, Gag00a, Gag98, Har96, Hil96b, Kin94, Reh97, SL95b, Tay01]. Distributing [Abl93b]. Distribution [Mal01, Pet01]. DLT [Swi98]. DNS [Gan94b, Jac00]. Do [Lym01b, Lym01c, Lym01d, McC00a, Sch99c, Won02]. Documentation [Kie98, Mar01b]. Does [Don99c]. Doesn’t [Lym01b, Lym01c, Lym01d]. Doing [Sch02d]. Domain [McC99a]. Don’t [Won02]. Door [Bur01b]. DOS/UNIX [Hun92d]. Down [Hu00, Wel98]. Downside [Tho98a]. Downtime [Bri94, Hor93b]. Drive [Swi98]. Driven [Sch96]. Drivers [Don97d]. Drives [AS93]. DSL [Zha99a]. Dump [Ash93, Hu98a, Nas97b]. dump/ restore [Hu98a]. Duo [Hel97a]. Duplicate [Nie00]. Duplicating [Lym01c]. Dynamic [Hel97a, Leg01].

E-Class [Cla01]. E-Commerce [Kla00]. E-Mail [Hil93, WG93]. E250 [Wat01a].

Easy [Ano97-27, Cli97, Isa96b, Kom00b, Kon93, Muh97b, Pre00a, Sch00j, Smi02a]. Economics [Hel00a]. Edge [Don99a]. Editor [Ank99b, Ank99a, Ank00a, Ank00b, Ank00c, Ank00d, Ank00e, Ank00f, Ank00g, Ank00h, Ank00i, Ank00j, Ank00k, Ank00l, Ank00m, Ank01a, Ank01b, Ank01c, Ank01d, Ank01e, Ank01f, Ank01g, Ank01i, Ank01h, Ank02a, Ank02b, Bar96a, Bar96b, Bar96c, Bar96d, Bar97a, Bar97b, Bar97c, Bar97d, Bar97e, Bar97f, Bar97g, Bar97h, Bar97i, Bar97j, Bar97k, Bar98a, Bar98b, Bar98c, Bar98d, Bar98e, Bar98f, Bar98g, Bar98h, Bar98i, Bar98j, Bar98k, Bar98l, Bar99h, Bar99i, Bar99j, Bar99b, Bar99c, Bar99d, Bar99e, Bar99f, Bar99g, War94a, War94b, War95a, War95b, War95c, War96a, War96b, War96c, War96d, War96e, War96f, Ank01k]. Education [Ano01-58, Ano01-59, Ano01-60, Ano02m, Ano02n]. Effects [Sal01]. Efficiency [Nak01]. Efficient [Dav01a]. Either [Dav01a]. Electronic [Abl92c]. ELM [Wei92b]. Email [Dil02, Gag00b, Gar02, Hel96c, Liu96, Mag95, Ris00, Sch97b]. Emergence [Hel00b]. Emulators [Bro92b]. Enabled [Isa98, Lib02]. Enabling [Gag99a]. Encrypted [Sta02]. Encryption [Cou93b, Sch99a, Zol93a, Hon98a]. End [Ric97, Swe99]. Energy [Hel01c]. Enforcer [Bot93]. Engine [Car01a, AS96]. Enhance [Don96f]. Enhanced [Che96b]. Enhancement [Zol93b]. Enhancing [Bus99b]. EnlightenDSM [Gag00a]. Enough [CY98]. Enterprise [Kha01, Sch99b, Kom00a]. Environment [Cox94, Don95a, Don98e, End96, Gau98, Hel96b, Hel97c, Hof97, Len94, Sef93, Smi96, Tay01, WM00, Wen01]. Environments [Cli97, Har96, Raj01, Sal98]. Equipment [Hel97d]. Equivalency [Har93d]. Error [Ric94a, Skl01]. Eschewing [Fel97b]. etc [Ess98a]. Ethereal [Les01]. Ethernet [Hel99b, Lee96]. Evaluating [Dul00a]. Event [Agu99, Agu00, Agu01]. Everything [Ron01]. Exabyte [Swi98]. Examining [Dul00b]. Exclusive [Bus99a]. Executable [RK97]. Exercise [Pet01]. Expect [Lin99, Sch99g]. Expiration [Hu98b]. Expiring [Hil00]. Exploits [Lau99]. Exploring [Mag96]. Exporting [Cli97]. Expressions [Rez92c, Sch99h]. Extension [Cro98]. External [Fel97b].

Facilities [Rie93, Rie94]. Facility [Skl01, WG93, Zol93c]. fack [Cla94]. Fail [Win99]. Failed [Cou93c]. Failover [Sch02c]. Fast [AS93, CY98, Dav01a, Gar00, McC00a]. Fastest [RB01a, RB01b]. Favorite [Fie92]. fcd [Bru94]. fchange [Isa96a, Isa93b]. FCheck [McC00b]. Features [Sam01]. Fibre [Nel99, Pre00c, Pre01d, Sch98]. Field [Ric99]. Figuring [Sch99i]. File [Agu99, All98, Ber98b, BN01, Bla99, Cal01, Cas01b, Cay94b, Cla94, Cox95, Har92d, Har92e, Har94b, Hu98a, Isa93b, Isa93c, Isa94b, Jon99, Kin94, Kla98, Lee95c, Lei95b, Lym01a, Mal01, May95, McK98c, Mer98a, Oli01, Plo98, Pre01e, Rez93b, Smi02a, SA95, WF01]. File-by-File [Jon99]. Filenames [SB96]. Files [Bal99, BL02, Cou92, Isa96b, Mic99, Mil00a, Ped02, Pol01, Rez95a, Sch00b, Tot00]. Fileserver [Kos96]. Filesystem [All96, Bol00, Gen93a, Hil96b, Lib02, Woo96]. Filesystem-Based [Lib02]. Filesystems [Dul95a, Mor96a, Vel94b]. Filter [McC00d, Muh97a, Nap01]. Filtering [Don98b, McC99e]. Filters [Bro95]. find [Rez92a]. Finding [Rez95c]. Firewall [BB99, Don96c, Fev01, Har95a, Luc00a, Mey01, Pra01a, Sus00]. Firewalls [Gan98, Hel97b, Iba97, Mur02, Tho98a]. First [Ada97, Qui01b]. Flash [How01]. Flexibility [Bye99]. Flexible [Cas01a, Oli01, Par01]. Floppy [Mey01]. Follow [RB01a]. Follow-Up [RB01a]. Foreign [Mor96a]. Forensic [Apt02]. Forget [Bry97a]. Form [Hel01a, San93b, WG93]. Forum [Ank99b, Ank99a, Ank00a, Ank00b, Ank00c, Ank00d, Ank00e, Ank00f, Ank00g, Ank00h, Ank00i, Ank00j, Ank00k, Ank00l, Ank00m, Ank01a, Ank01b, Ank01c, Ank01d, Ank01e, Ank01f, Ank01g, Ank01h, Ank02a, Ank02b, Bar96a, Bar96b, Bar96c, Bar96d, Bar97a, Bar97b, Bar97c, Bar97d, Bar97e, Bar97f, Bar97g, Bar97h, Bar97i, Bar97j, Bar97k, Bar98a, Bar98b, Bar98c, Bar98d, Bar98e, Bar98f, Bar98g, Bar98h, Bar98i, Bar98j, Bar98k, Bar98l, Bar99h, Bar99i, Bar99j, Bar99b, Bar99c, Bar99d, Bar99e, Bar99f, Bar99g, Rez95e, War92a, War92b, War92c, War92d, War93a, War93b, War93c, War93d, War93e, War93f, War94a, War94b, War94c, War94d, War94e, War94f, War95a, War95b, War95c, War96a, War96b, War96c, War96d, War96e, War96f, Ank01k]. Forwarding [Dav99a, Sch97b]. Frame [Coo98]. Framework [Cal01, Jos01]. Free [Bro97, Lau02]. FreeBSD [DB02, Hel98g, Luc00a, Luc00c, Luc00b, Luc00d, Pla00, RB01a, Sar01]. FreeS [Nap00]. FreeS/WAN [Nap00]. Freeware [Ali99, Bah01, BmC01, Dub01]. Freeware-Based [Ali99]. Friendly [Sal97b]. Front [Bur01b, Ric97]. FrontPage [Wid00]. Frozen [Bro92a]. fsck [Woo96]. FSDB [Har92d]. FTP [CC01, Gre01, Mes94, Hil00, Wat01b]. Fun [Sch00h]. Functionality [Dul96b, Mag96]. Functions [Raa01, Sch02b]. Fuzzy [GN01a]. fvwm [Bro96].

Gate [Mag95]. Gatekeeper [Sch99e]. Gateway [Gag00c]. Generation [KR94, Nap01, Sat95a, Zha00]. Generator [Wil93]. Gentle [Bot93]. Gentler [Dea92]. Geographic [Sch02c]. Get [Abl93c, Hil93]. GetHost [She02]. Gets [Pre00a]. Gettemp [Kom00a]. Getting [Bee01, Don97a, Har92a, Har92b, LA01, Ram01, Sch00f]. Ghost [Tho98b]. Gigabit [Hel99b]. Glasses [Don98d]. Global [Raa01]. GNATS [Plo97, Sal97a]. GnuPG [Hag01]. Go [Dul94c]. Golden [Gal02b]. Good [McC00c]. gopher [Bus94b]. Grail [Haz99a]. Graphical [Bau96, Bun96, Hal97, Ric97]. Graphing [Mic98]. Ground [Bar99j]. Group [Mor96c, Rez95a]. GroupWise [Fel96b]. GroupWise/SMTP [Fel96b]. Growth [Zwi98]. GUI [Bro00, Don96f, Don96h, Pla00]. Guide [Bar98m, End97, Hun92a, May97].

Hack [Sei00b, Sei00c, Sei01]. Hacked [Hug01]. Halted [Mur02]. Handling [Dul96a]. Handy [Cou93c]. Hard [Dea93]. Hardening [Qui01b, Zwi99b]. Haunted [Tho98b]. Header [Rea93]. Health [Dav99b, Hel99a, Lee94a]. Heartbeat [BN01]. Help [Har95d, Mel96, Tho98c]. Helpdesk [Sha97]. Helping [BB01]. Heterogeneous [Fer01a, JK94]. Hidden [Mor96a]. Hiding [Rez95d]. Hierarchical [Tho99]. High [Bar97l, BN01, Gar96, Jac99b, Reh01a, Rez93a, RB01b, Smi01b, Vad98, Vel95a]. High-Availability [BN01, Reh01a]. High-Capacity [Vel95a]. High-Performance [RB01b]. High-Speed [Bar97l, Gar96, Rez93a]. Hogs [Rez95c]. Holy [Haz99a]. Home [Lee96]. Homebrew [Kue01, Smi01b]. Homegrown [Str02]. Honey [End00]. Horse [Abl92a]. Host [Ho01, Hor93b, Lee94a, McC00b, McK01, Sat95a, Zwi99b]. Host-Based [McC00b, McK01]. Hosted [Coo98]. Hosting [Gre01, Soh01]. Hosts [DeJ02, Rib02, Ess98a]. Hot [Pre01f]. HP [Bro97, Kle97, Rez95h]. HP-UX [Kle97, Rez95h]. HTML [Bun96, Mur00]. HTTP [Fre02]. Hub [Rez95h]. Human [Sal97b]. Hung [Bro92a]. Hybrid [LMBMR98]. Hype [Tho98c].

I/O [Har96, McK99a, New02, Ric96a]. IBM [Den99a, Wil01a]. ICMP [McC00c]. I’d [Plo00]. Idle [Rez93c]. II [McC00f, McC01d, Pre01c]. Illegal [End98]. Illusion [Hel99b]. Image [Jon99]. Image-Based [Jon99]. Implementation [Don95b, McK98d, Zha99a]. Implementing [Car97, Dav00c, Gan95, Hel00c, Lam99a, Smi01a, WM00, Wen01, Zha99b]. Improved [Ric94a]. Improving [De 98a, Nak01]. In-line [Zol92c]. Increasing [Apt02]. Independent [Gil00]. Index [Ano92a, Ano92b, Ano92c, Ano92d, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano95a, Ano95b, Ano95c, Ano95d, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano00m, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano02a, Ano02b]. Index [Ano02c]. Inexpensive [Fel97c]. Info [Har92a, Har92b]. Information [Abl93b, LE97, Hel96c, Kin01, SD96]. Informix [NS94]. Inodes [Dul94c]. Input [Zol92c]. Ins [Don98b]. Install [Woo01]. Installation [Fer01a, Mar01a, Moo97, Rez94b, dRV01]. Installations [HD97]. Installing [Cay94a, Hun92b, Les00, Mur01, Tan98a]. Instant [Jos02]. InstantInfo [Ano93m, Ano93n, Ano94m, Ano94n, Ano94o, Ano94p]. Instead [And01]. Integrating [All98, Bak95, Hal01, Hal02, Hel97c, Kha01, Sch98, Wit01]. Integrit [Cas01b]. Integrity [Vel94a]. Interactions [BL99]. Interest [McC99c]. Interesting [Sch01e]. Interface [Bau96, Bay00, Bun96, Luc00c, Rez93b, Sal97b, Sch96]. Interfaces [Jay98, Sch97c]. Interlinking [Joh96b]. Intermec [Sch01a]. Internet [Bun97, Bus94a, Bus94b, Bus99d, Dea93, Don96g, Fel97a, Fel97b, Gag00c, Hel96a, Her02, Mag95, SD96, Wid01, Won02]. Interoperating [Wei01b]. Interpreting [Dav97]. Interprocess [Joh96a]. Interrupts [Vel98a]. Intranet [Hon98b, Fel97a]. Intranets [De 97b]. Introducing [Nap00]. Introduction [FM99, How01, May96, Sei00b, Wil01a]. Intruders [Dav96]. Intrusion [Ano01-32, Cha01, Dub01, Iba97, Kue01, McC99d, McC00b, Wes00a]. IOS [Fre02, Sam01]. iostat [Gen94b]. IP [BB99, Bun97, Dav99a, Don98c, Jac00, Jon02, McC00d, McC00e, Rus96]. IPSec [FM99, Nap02, Nap00, Wen01]. IPTables [Nap01]. IPTables/NetFilter [Nap01]. iptool [Ess98a]. IPv6 [Gar98b]. isascii [Isa96b]. ISDN [Don97b]. Islands [Don96e]. Isn’t [CY98, Ron01]. Isolating [Nic00]. Issues [LS97].

Jabber [Jos02]. Jail [Sar01]. Jailed [Wid01]. Java [Don96f, GW00, Ron01, WM97]. Job [Bro93, Zol92a, Zol92b]. Joining [Wei92a]. Journal [Zol93a]. Journal/Logging [Zol93a]. JServ [GW00]. Jukebox [Car97]. JumpStart [Jos01, vdW00]. Jungle [Con96, Hun92a].

Keeping [Har93e, O’G01, Ols01a]. Kerberos [Smi01a, Wit01]. Kernel [Kie99b, Ric96a]. Keys [Str02]. Kickstart [O’K00]. kill [Isa94a]. Killidle [Gan94a]. Kinder [Dea92]. Known [Sam01]. Korn [Raa01, Sch02b]. Kstat [GN01b].

LAN [Bun97, Hel98d, Hel99e, Pod01, Zha99b]. Landmining [Ric99]. Languages [Rez93f]. LANs [Won97]. Large [Che96a, HD97, Hun92c, Nak01, Pet01, Pre01e, SL95b]. Large-Scale [Hun92c, SL95b]. Latency [Hel97a]. Launcher [Zol92a]. Layer [Hel97d, Hel99d, Hel00d]. LDAP [Bri00, De 97a, Gre01, Vog00c]. League [Pre00b]. Learned [CC99]. Legacy [Gag99a]. Legitimizing [mH99]. Lessons [CC99]. Letter [San93b, WG93]. Letters [Ano97-28]. Level [Jac99a]. Levels [Dul95b, Kno02, Luc99]. Lib [New02]. Libraries [Bar99a]. Library [Bar98m, Gag98, Gan98, Muh97a]. Life [Dea93]. Like [Sch02d]. Line [Pru02, Sch00a, McC00h, NS94, Zol92c]. Liners [Fie92]. Lines [BL99]. Link [De 97a, KR94, Vog00b]. Links [Qui98]. Linux [And99, Ano01-32, BB99, Bro00, BL02, Bun97, Bus01, Coo98, Coo01, Don96c, Don96g, Don97b, Don97d, Don98e, Fel96a, Fel96c, Fer01b, Fer01a, Gag00c, Gag00b, Gol01, mH99, Her02, Kie99b, Kie99a, Lam99a, Luc00d, Mar00, Nap01, Nap02, Nie00, Reh99, Reh00, Smi01b, Sta02, Vad98, Wei01b]. Linux-Hosted [Coo98]. Lists [Hel99c, Pet01]. Little [Rez93f, Sam01, Sch00g]. Living [Don99a]. Load [Gou01, Zol92d]. Local [Abl93b, Lee94b, Mag95, Mag96]. Location [Kin94]. Locked [Bro92a]. Locking [Bur01b, Cox95, Mor96b]. Log [Agu99, Bal99, May95, Mic99, Pra01a, Tar96]. Log-Checking [Pra01a]. LogCheck [Cin01]. Logging [Isa94a, Rib02, Skl01, Wei98]. Logical [Gen94a, Reh00, San98]. Login [Cou93c, Eve95, Har93c, Hil93, Ho01, Rez95f, Ric95]. logit [May95]. Logs [Naz01, Pre01f]. LogViewer [HS00]. Long [Vel95a]. Long-Term [Vel95a]. Look [McC00f, McC01a, Sch99f, Sch00f, Wes00a]. Looking [Don98d]. Lost [Plo00]. Low [Liu96, Vad98]. Low-Cost [Liu96, Vad98]. lp [Zol93b]. LPI [Dul00b]. LPRng [Mil00b]. lptran [Cou93c]. LUDE [Gag98]. LUI [Fer01a]. LynxArray [van97a]. LynxStak [van97a].

Machine [Wad96]. Machines [She02]. Macintosh [Hal01]. Macro [Bus99a, Ric94b]. Macs [Lei95a, Liu96, Sto95]. Made [Ano97-27, Bol00, Cli97, Kom00b, Muh97b]. Magic [Haz99b, Sch00g]. Mail [Abl92c, Hil93, Hun93a, Hun93b, Pla00, Rez94d, Rez95h, Sal96, WG93]. Mailer [Den99a]. Mailman [McC01b]. Maintainable [Tan98b]. Maintaining [Ess98a, Luc99]. Maintenance [Con96, Don96b]. Make [Dia00]. Making [Bay00, Lin93a, Tee00, Vel95b]. Mammoth [Swi98]. Man [Car01a, Jac99b, Isa98]. Manageability [LMBMR98]. Management [Agu99, Agu01, Bro97, Eve95, Gag00a, Har94c, Hel99e, Hof97, Lei95b, Lym02, Pur01, RW96, Sal96, She98, Smi96, Tho99, Vog00b, Zha98, Zol93c]. Manager [Ber95, Cou93a, Gen94a, May95, Qui00, Reh00, Wol00, Woo01, Cer01a]. Managing [Bry00, Car01b, Don95a, Gag98, GN01b, Hon98b, Hu98a, Kie98, Koe98, Lei96, Pet95, Pip93, Plo97, Raj01, Rib02, Sal97b, Tis98, Tot00, Vog00c]. Many [Sch02d]. Map [Swe00a]. Marketplace [Ano01-33, Ano01-34, Ano01-35, Ano02g, Ano02h]. Masquerade [Bun97]. Masquerading [BB99]. Masshosts [Mec99a]. Matters [Hel01c]. maxtab [Isa93c]. Me [Bry97a, Don99c]. Measuring [Nak01]. Mechanisms [Mor96b]. Memory [Nak01]. Menu [Che97, Sch96, Whi96]. Menu-Driven [Sch96]. Menus [Rez95f, Tay01]. Merging [Mil00a]. Message [Ram01]. Messages [Ano92i, Ano92j, Ano92k, Ano92l, Ano93o, Ano93p, Ano93q, Ano93r, Ano93s, Ano93t, Ano94q, Ano94r, Ano94s, Ano94t, Ano94u, Ano94v, Ano95i, Ano95j, Ano95k, Ano96-29, Ano96-30, Ano96-31, Ano96-32, Ano96-33, Ano96-35, Ano96-34, Ano96-36, Ano97-29, Ano97-30, Ano97-31, Ano98-54, Ano98-55, Ano99-49, Ano99-50, Ano99-51, Ano99-52, Ano00-41, Ano00-42, Ano00-43, Ano00-44, Ano00-45, Ano00-46, Ano00-47, Ano01-36, Ano01-37, Ano01-38, Ano01-39, Ano01-40, Ano01-41, Ano02i, WA99]. Messaging [Jos02]. Method [Vel94a]. Methodology [Ber98b]. MHonArc [Gar02]. Microsoft [Wid00]. Migrating [All96, Sch00d]. Migrational [Soh01]. Millennium [Haz99a, MS99]. MIME [Sch01d]. Minded [Gra96]. Minimal [Bay01, Coo96]. Mirroring [Cer01b, Rya94]. Missing [De 97a, Sch99j]. Mixed [Don95a]. MKS [Dul00a]. Model [Dav00d, Dav00e, Hil96b]. Modem [Rez93a]. Modules [Sei00d]. Monitor [Ber95, Ess98b, Gen92, Gen93a, Har95b, Mac97]. Monitoring [Ber99, Ber94, Bic02, Con97a, Dav00a, Dav01b, Dil02, GS00, Gen94b, Gra96, Har93c, Kom00a, Ram00, She01, SN95, Wyn97, Zol92d]. Most [Lin93a]. Mounting [Mor96a]. Moving [Sch94]. MPLS [McC00f, McC00e, Zha00]. MS [SB96]. MS-DOS [SB96]. MTRG [Kra00]. Muhamedagić [Ano97-27]. Multi [Gol01, Hu98a, Nas98a, Nas98b, Pod01, Wak02]. Multi-Booting [Gol01]. Multi-File [Hu98a]. Multi-Platform [Wak02]. Multi-Tiered [Nas98a, Nas98b]. Multi-Vendor [Pod01]. Multilayered [Wet95]. Multiple [Hal02, Mor96c]. Multiport [Rez94b]. MVS [Joh95, Joh96a, Joh96b]. My [Fie92, Wel98]. MySQL [Jen01, Kla00, Sol02].

Name [McC99a, McC02]. NAT [McC00g]. Naval [FS95]. Nessus [Lau02]. ’Net [Mer99, MD94, Mer99]. Netgroup [Wyn94]. NetReg [VW00]. netstat [Dav97]. NetWare [Dul96a, Wei01b]. Network [All96, Ben98, Ber99, Bic02, Bou99, Bro97, Bus99b, Dav99b, DB02, Don96b, Dul94b, LE97, Fel97c, Gag99b, Gal01a, Goo98, Goo99, Hal01, Hal02, Har93d, Har95a, Hel97a, Hel97b, Hel97d, Hel99a, Hel00e, Hel00f, Hel00h, Hel01b, Jay98, Jos00, Kin01, Kol99, Lee94b, Lib01, Mac97, McC00g, McC00h, Qui01b, Reh97, RB01b, San99, Sch00d, Sto01, Woo01, Ber98b, Wil01a]. Networked [Sef93]. Networking [Bar97l, Har94a, Har95c, Hun92d, Kue00, Rus96, Zha00]. Networks [Bla99, EK00, Jon02, McC00a, Nap02, Sul00, Ano00-40, Tho98c, Wha98]. newping [Hor93a, Hor93b]. News [D’A96, Dav01b, Hu98b, Mag96]. Next [Nap01, Sat95a, Zha00]. Next-Generation [Nap01]. NFS [All98, Ber94, Fel94, Lin93a, McC01d, McC01c, Mor96a, Rya94, Sta02]. ngrep [McC01a]. NICs [Kra01]. NIS [Abl93b, Bak95, Pap95]. nmap [Kom00b]. nmap-web [Kom00b]. NNTP [Jos00]. No [Haz99a, Ano97-27]. Non [Bla99, Joh96c]. Non-Connected [Bla99]. Non-TCP [Joh96c]. Non-TCP/IP [Joh96c]. Notification [Bri94, Cro98]. NT [Tis98, Don97a, Don98d, Fla97, Hel96b, Hel97c, Hel97e, Hel98b, Joh97, Kos96, Nas97a, Oku98]. NT/UNIX [Nas97a]. Numbering [Har94b]. Numeric [McD95]. Numerous [Rib02]. Numlock [Rez93d].

O [Har96, McK99a, New02, Ric96a]. Object [Cou93a]. Objects [Jon98a, Lee95c, Vog02]. Observatory [FS95]. ODM [Cou93a]. Office [And99, Don98e]. older [Rez92a]. On-line [McC00h, NS94]. One [Coo01, Fie92, Har95c]. One-Liners [Fie92]. Online [Ano94-29, Ano94-30, Bea96, Bus94a, Bus94b, Har95d]. Open [Cha01, Gag02, Hag01, Luc99, Reh01b]. OpenBSD [Mur01, Wit00]. OpenServer [RW96]. OpenSSH [Les00, Sta02]. Openview [Bro97]. Operating [Hal02, Reh01b, Tus95]. Operations [Apt02, McC00h]. Opportunities [Ano96u, Ano96v, Ano96w, Ano96x, Ano96y, Ano96z, Ano96-27, Ano96-28, Ano97q, Ano97r, Ano97s, Ano97t, Ano97u, Ano97v, Ano97w, Ano97x, Ano98y, Ano98z, Ano98-27, Ano98-28, Ano98-29, Ano98-30, Ano98-31, Ano98-32, Ano98-33, Ano98-34, Ano98-35, Ano98-36, Ano99y, Ano99z, Ano99-27, Ano99-28, Ano99-29, Ano99-30, Ano99-31, Ano99-32, Ano99-33, Ano99-34, Ano99-35, Ano99-36, Ano00-27, Ano01t]. Optimized [McC00a]. Optimizing [Ber94, Hel98b, OY01]. Option [Rez92a]. Options [Joh97, Sch00a]. Oracle [Dav00c, LB96, Man01b, McK01, OY01, Pre96, Pre01f, Pre01h]. Orchestrating [Dea92]. Other [Dul94a, Nas97b, Swe99]. Out/Getting [Bee01]. Outlook [Pet01]. Outs [Don98b]. Overcoming [Pet01]. Overnight [Zol92b]. Overview [Gar98b, Oku98]. Overworked [Pra01b, Pra01c]. Own [Pre00b]. Ownership [Mor96c]. OWS [Dav00c].

Package [Ric94b, She01]. Packages

[Zwi99a]. Packet

[Dea93, Don98b, McC99e, Mec99b, Nap01]. Packets [Dav98a]. Pager [Dyk96, Gom94]. Pages [Isa98]. Paging [McD95]. Painless [Can99, Gar00]. PAM [Sei00d]. Papers [Ano92e, Ano92f, Ano92g, Ano92h, Ano93g, Ano93h, Ano93i, Ano93j, Ano93k, Ano93l, Ano94g, Ano94h, Ano94i, Ano94j, Ano94k, Ano94l, Ano95e, Ano95f, Ano95g, Ano95h, Ano96k, Ano96l, Ano96m, Ano96n, Ano96o, Ano96p, Ano96r, Ano96q, Ano96s, Ano96t, Ano97i, Ano97j, Ano97k, Ano97l, Ano97m, Ano97n, Ano97o, Ano97p, Ano98m, Ano98n, Ano98o, Ano98p, Ano98q, Ano98r, Ano98s, Ano98t, Ano98u, Ano98v, Ano98w, Ano98x, Ano99m, Ano99n, Ano99o, Ano99p, Ano99q, Ano99r, Ano99s, Ano99t, Ano99u, Ano99v, Ano99w, Ano99x, Ano00n, Ano00o, Ano00p, Ano00q, Ano00r, Ano00s, Ano00t, Ano00u, Ano00v, Ano00w, Ano00x, Ano00y, Ano00z, Ano01l, Ano01m, Ano01n, Ano01o, Ano01p, Ano01q, Ano01r, Ano01s, Ano02d, Ano02e, Ano02f]. Parsing [Sch01e]. Part [Gen94d, Gen95, Har01a, Har01b, Hel00e, Hun93a, Hun93b, McC00e, Nas98a, Nas98b, Pre00c, Pre01d, Sei00c, McC00f, McC01d, McC01c, Pre01b, Pre01c]. Partitions [Cay94b]. Party [Cay94a]. Password [Bur01b, CB96, Har92c, Har01a, Har01b, JK94, Rez95a, Ric96b, Sch99a]. Passwords [Dav98b, Haz99a, Hil00, McK98a]. Patch [Kno02, Luc99, Won97]. Patching [Gaf99]. Patrol [Boe97]. PCASSO [CC99]. PCs [LE97, Lei95a, Wei95]. 6000 [Pet94]. Cobalt [Gal02f]. Finish [Jos01]. Getting [Bee01]. IP [Joh96c, Joh97, Sch95, Sto95, Tis98]. Linux [Luc00b]. Logging [Zol93a]. NetFilter [Nap01]. NT [Gau98]. PC [Fel94]. PPP [Don95a]. RCS [Dav00c]. restore [Hu98a]. Server [May96]. SGID [Tot00]. SMTP [Fel96b]. terminfo [Tho95]. Time [Sch95]. Tk [BB01, Hal97, Tar96]. UNIX [Nas97a, Hun92d]. WAN [Nap00]. Windows [Don95b]. Workstation [DW01]. Penetration [KP99, PK99]. Perform [Dil02]. Performance [AF00, Bec99, Ber94, Ber98b, De 98a, Gen92, Gen94b, Got01, Hel96b, Hel97a, Hel98b, LS97, McK99a, New02, Nic00, RW96, RL98, RB01b, Smi96, Vad98]. Performing [Dav00a]. Perl [App00, Swe00a, Tay99, BB01, D’A96, LB96, Sch01b, She01, Tan98b, Vog00c, Vog02, dRV01, Hil95]. Perl/Tk [BB01]. Perls [Bus99c]. Permissions [Dul96a, Oli01]. Perspective [Har93f]. PGP [End97]. Phone [Sch99i, Swe99]. Phone-y [Sch99i]. PHP [Jen01, Mob98, Ped02]. Physical [Jam94]. PICA [dRV01]. Piece [Har95c]. ping [Jac99a]. pings [Sch02d]. Pipe [Sch93]. Plaintext [Haz99a]. Plan [Zwi98]. Planning [Hel96b, Zwi98, Zwi99c].

Platform [Agu00, Gil00, Wak02].

Platform-Independent [Gil00]. Playing [Ric99]. Pluggable [Sei00d]. Point [DB02]. Policy [Mar00]. Poor [Car01a, Jac99b]. Port [Don98a, Kom00b, McC00a]. Portability [Wei92b]. Porting [Har96]. PortSentry [Cin01]. Poster [Ano01-32]. Posting [Gar02]. PostScript [Cer97]. Posture [KP99]. Pot [End00]. Potential [Hel99b, RK97]. Potpourri [Hun92c]. Powerdowns [SL95b]. Powerful [Lau02, Pol01]. pp [Ano97-27]. PPP [Ano97-27, Dia00, Gar96, Har94a, Kue00, Muh97b]. Practical [App00, Jac00, May97, Mec99b]. Practices [Gal02e]. Preparing [Hel01a]. Prevent [Hel97b]. Prevention [Hil97, Sch99d]. Primer [Goo98, Pre00c, Pre01d]. Print [Ber95, Bro95, KCW98, Lei95a, SA95]. Printer [Cou92, Cou93b, Hun92c, Lym02, Pan94, Rez92b, Rez93b]. Printers [Sch01a]. Printing [Cer97, Hof97, Lee95a]. Priorities [Rez94c]. Private [Nap02, Tho98c, Wha98]. Privileges [Pip93]. Proactive [Sch99d]. Probe [Lee94a]. Problem [Bry97b]. Problems [Ber98b, Nic00, RL98]. proc [Kla98]. Process [Har94c, Har95b]. Processes [Isa92, Wei93]. Processing [Sch00a]. procmon [Har95b]. Product [Ano97-40, Bar97l, van97a, van97b]. Production [Vog00a]. Products [Ano92m, Ano92n, Ano92o, Ano92p, Ano93u, Ano93v, Ano93w, Ano93x, Ano93y, Ano93z, Ano94w, Ano94x, Ano94y, Ano94z, Ano94-27, Ano94-28, Ano95l, Ano95m, Ano95n, Ano95o, Ano96-37, Ano96-38, Ano96-39, Ano96-40, Ano96-41, Ano96-42, Ano96-43, Ano96-44, Ano96-45, Ano96-46, Ano97-32, Ano97-33, Ano97-34, Ano97-35, Ano97-36, Ano97-37, Ano97-38, Ano97-39, Ano98-56, Ano98-57, Ano98-58, Ano98-59, Ano98-60, Ano98-61, Ano98-62, Ano98-63, Ano98-64, Ano98-65, Ano98-66, Ano98-67, Ano99-53, Ano99-54, Ano99-55, Ano99-56, Ano99-57, Ano99-58, Ano99-59, Ano99-60, Ano99-61, Ano99-62, Ano99-63, Ano99-64, Ano00-48, Ano00-49, Ano00-50, Ano00-51, Ano00-52, Ano00-53, Ano00-54, Ano00-55, Ano00-56, Ano00-57, Ano00-58, Ano00-59, Ano00-60, Ano01-42, Ano01-43, Ano01-44, Ano01-45, Ano01-46, Ano01-47, Ano01-48, Ano01-49, Ano01-50, Ano01-51, Ano01-52, Ano02j, Ano02k]. Products [Ano02l]. Program [Che97, Isa95, Tar96, Whi96]. Programming [Iye97]. Programs [Fie92, Rac00, Tan98b]. Project [Zol93c]. Properties [Kin94]. Protection [Hag01]. Protocol [AF00, McC00k, McC01c, WT00]. Proxies [Won02]. Proxy [Bus99d, Fel98, Reh99, Wol00, CC01]. prq [Ber95]. Pruning [Isa93c]. Public [Jon02]. Public-Access [Jon02]. Publicly [Don99b]. Publisher [Rez95e, War92a, War92b, War92c, War92d, War93a, War93b, War93c, War93d, War93e, War93f, War94c, War94d, War94e, War94f]. Publishing [Vog02]. Pushing [Ris00]. Putting [Abl92b]. Puzzle [Ano98-50, Ano98-51, Ano98-52, Ano98-53, Ano98-68, Ano98-69].

QoS [Hel00e, Hel00f, Hel01b]. Quality [Hel00d, Hel00g, McC00e, McC01f]. Quantum [Swi98]. Qube [Gal02f]. Questions [McK99b, McK99c, McK00b, McK00c, McK00d, McK00a, McK00e, McK00f, McK00g, McK00h, McK00i, McK00j, McK00k, MR00, MR01a, MR01b, MR01c, MR01d, MR01e, MR01f, MR01g, MR01h, Ric01a, Ric01b, Ric01c, Ric01d, Ric01e, Ric02a, Ric02b, Ric02c, Ric02d, Ric02e, Sat92a, Sat92b, Sat92c, Sat93a, Sat93b, Sat93c, Sat93d, Sat93e, Sat94a, Sat94b, Sat94c, Sat94d, Sat94e, Sat94f, Sat95b, Sat95c, Sat95d, Sat95e, Sat95f, Sat96a, Sat96b, Sat96c, Sat96d, Sat96e, Sat96f, Sat96g, Sat96h, Sat96i, Sat96j, Sat97a, Sat97b, Sat97c, Sat97d, Sat97e, Sat97f, Sat97g, Sat97h, Sat97i, Sat97j, Sat98a, Sat98b, Sat98c, Sat98d, Sat98e, Sat98f, Sat98g, Sat98h, Sat98i, Sat98j, Sat98k, Sat99a, Sat99b, Sat99c, Sat99d, Sat99e, Sat99f, Sat99g, Sat99h, Sat99i]. Queue [Ber95]. Queuing [Hel00f]. Quick [DW01, Lib01]. Quickly [LA01]. Quota [KCW98, Sol02, Bot93].

R [Koe98]. R* [Fla97]. R.I.P. [McC01e]. R/3 [Koe98]. RADIUS [McC99b, Reh97]. RAID [Hel98c, LS97, McK98d, Smi01b, Vad98, van97a]. RAID-1 [Smi01b]. Rainy [Sch00h]. Rate [Hel96a]. Rays [Raj01]. RCS [DS00, Kie98, Plo98]. RDB [Pru02]. RDBMS [Smi96]. rdist [Ash92, Wat01b]. Read [Swe00b]. Readership [Hu98b]. Reading [Rea93]. Real [Cer01b, Joh95, McK98d, Reh01b]. Real-Time [Cer01b, Joh95, Reh01b]. Real-World [McK98d]. Really [Don98c]. Rebooting [Rez95g]. Recognizing [O’B96]. Recovered [Plo00]. Recovering [BL02, DS00, O’B96]. Recovery [Cer01a, Gar00, Pre01h, Ric98, WF01]. Recruitment [Ano01-53, Ano01-54, Ano01-55, Ano01-56, Ano01-57]. Recycle [Sto01]. Red [Reh01a]. Redo [Pre01f]. Reducing [Vel98a]. Redundancy [Lib01, McC00k]. Redundant [Her02, Kra01]. refer [Ric94b]. Refresh [BL00]. Registration [VW00]. Regular [Rez92c, Sch99h]. Relay [Coo98]. Reliability [Apt02, Ram00]. Reliable [Gal01a]. Remote [Cer01b, CB96, De 98b, Fer01a, Hel98a, Ho01, Hon98b, Hor93b, Lau02, MD94, Rez95f, Ano00-40]. ReOrg [Bol00]. Repair [Woo96]. Replacing [Wat01b]. Replication [DW01, McK98c, McK01]. Report [Cou93c, Hol97, Nol98, Wil93, Zin92a, Zin92b, Zin92c, Zin93a, Zin93b, Zin93c, Zin93d, Zin93e, Zin93f, Zin94a, Zin94b, Zin94c, Zin94d, Zin94e, Zin94f, Zin95a, Zin95b, Zin95c, Zin95d, Zin95e, Zin95f, Zin96a, Zin96b, Zin96c, Zin96d, Zin96e, Zin96f, Zin96g, Zin96h, Zin96i, Zin96j, Zin96k, Zin96l, Zin97a, Zin97b, Zin97c, Zin97d, Zin97e, Zin97f, Zin97g, Zin97h, Zin97i, Zin97j, Zin97k, Zin98a, Zin98b, Zin98c, Zin98d, Zin98e, Zin98f, Zin98g, Zin98h, Zin98i, Zin98j, Zin98k, Zin98l, Zin99a, Zin99b, Zin99c, Zin99d, Zin99e, Zin99f, Zin99g, Zin99h, Zin99i, Zin99j, Zin99k, Zin99l, Zin00a, Zin00b, Zin00c, Zin00d, Zin00e, Zin00f, Zin00g, Zin00h, Zin00i, Zin00j, Zin00k, Zin00l, Zin00m, Zin01a, Zin01b, Zin01c, Zin01d, Zin01e, Zin01f, Zin01g]. Report [Zin01h, Zin01i, Zin01j, Zin01k, Zin01l, Zol93d]. Reporter [Fev01]. Reporting [Agu00, Gra96, Ric94a, RL98]. Reproducing [Vel94b]. Requests [Isa94a]. Rescue [Ash92]. Resource [Vel94c]. Responsibility [BL99]. Restart [Ho01]. Restore [AS93, Ash93, Pet94, Mor95]. Restoring [Geb98]. Restricting [Wad96]. Retrieving [Hel96c]. Returning [Sch97a]. Review [BR99, van97a, van97b]. Revised [Pip93]. Revision [Plo98, Smi98]. Revisited [Gan94a, Hor93a, Jam94, Sal97a]. Revisiting [Har01a, Har01b]. Revolution [RK97]. Right [Bro93]. RIP [McC01e]. rlock [Cox95]. rma [Wad96]. Robin [Leg01]. RoboMon [van97b]. Role [Vin00]. Role-Based [Vin00]. ROM [Car97]. ROMs [Kuo98]. Root [Con97b, End98, Rac00]. Rootkit [O’B96]. Round [Leg01]. Router [Fel96c, Hel99c, Hel00f, Hel00h, Hol97, McC00k, Tis01]. Router-Based [Hel00h]. Routers [Lin99, Sus00, Tho98a, Zha98]. Routine [Sch01b]. Routing [Mar00, Ano00-40]. RPC [Orw01]. RPTRTR [Hol97]. RS [Pet94]. RS/6000 [Pet94]. RSA [Str02]. RSVP [McC01f]. Rules [Gal02b]. Rulesets [Sch99f]. Run [Dul95b, Gag02]. Running [O’G01]. s [CY98]. Safe [Bea96, EK00]. Safer [WB01]. SAINT [Ols01b]. Samba [Ada97, Knu96, Nas97a, Sto01]. Samba-Based [Sto01]. SAN [And01, Pre01b, Pre01c, Pre01g, Sch00c]. SANS [Kol99, Pre00d]. SAP [Koe98]. sar [Bry97b, Har92a, Ric97]. SAs [BL99]. Save [Won02]. Say [Don98c]. Scalable [Jos00]. Scale [Hun92c, SL95b]. Scanner [Lau02]. Scanners [BmC01]. Scanning [Don98a, Kom00b, Tis00]. Scaring [Ols00]. SCCS [Har93e]. Scenario [Goo99]. Scheduled [Rez95g]. Scheduler [Vel94c]. Scheme [KCW98, Sat95a]. Schemes [Lib01]. SCO [RW96, Ric96a]. Scotty [Bic02]. scp [Wat01b]. Screen [Isa99, Mor96b]. Screened [DeJ02]. Script [Coo01, D’A96, Gil00, Hel01a, Qui00, Sch97a, Won97]. Script-Form [Hel01a]. Scripting [Mur00, Tes96, WB01]. Scripts [Aki01, AS93, Dul96b, GT95, Hil96c, Jos01, LB96, Lee95b, Pet94, Pra01b, Pra01c, Sch99a, Sch02a, SN95, Sto92, Tan98b, Wei01a]. scsh [Sar02]. SCSI [Sch98, Swe00b]. Search [Car01a, McC00e, Pol01]. Searchable [Tes96]. Searches [GN01a]. Secrets [Sch00e]. Secure [Bal99, De 98b, Den99a, Don96d, End96, Kra99, Len94, MD94, Mal01, Mer99, Tee00, Zol93a]. Securing [Dav98b, De 97b, Den99b, Don97c, Dub00, Jon02, Qui01a, Sar01, SB99, Tis01]. Security [Abl92c, Ali99, BR99, Bah01, Bal99, Bur01b, Bus99b, Che96b, Con97b, Dav00d, Don96a, Don98d, Don99b, Dub02, FPA99, Fel96b, Fel98, GS00, Har93a, Har93d, Har98, Hel98g, Jam94, Kol99, KP99, Lam99a, Lau02, MD94, May96, May97, Nas98a, Nas98b, Oku98, Pap95, Reh97, Rez94a, Rie93, Sam01, Sef93, SD96, Tis00, Wei01a, Wet95]. Selecting [Bro00]. Selective [D’A96]. SelectNews.pl [D’A96]. Sendmail [Hun93a, Hun93b, Mag97, Pet01, Sch99f, Don95a, McC00i, Sch99b, Sch99e]. Sensitive [Hu98b, Sal98]. Sentinel [Gen93b]. Serial [Gar96, Gol98, Har96, Rez94b, Vel98a, Wei95, Pur01]. Serious [Iye97]. Serve [Lym01b, Lym01c, Lym01d]. Server [BN01, Bus99d, Bus01, Coo96, De 98a, DW01, Don96g, Don97c, Fel96a, Fel98, Gag00b, Gal02f, GW00, Hel96a, Jos00, Lei95a, Lym01b, Lym01c, Lym01d, Mag96, Pla00, Ram00, Reh01a, Sch95, SA95, SB99, Tan98a, vdW00, WM97, Zwi98]. Server-Based [Sch95]. Server-Side [WM97]. Server/Workstation [DW01]. Servers [Kom00a]. Service [Hel00d, Hel00g, McC01f, Sal01, McC00e]. Services [Bus94a, Bus94b, De 97a, FS95, Ho01, McC02, O’G01, Orw01, Wid00, Wid01, Zha00]. Sessions [Bro96, De 98b]. Setting [Cay94b, Len94, Rac00, Rez92b, Rez93a, Rez93b, Rez94c, Rez95h]. Setuid [Cox94]. Setup [Mer98b, Sus00]. SGI [Gar00]. Shared [EK00]. Sharing [All98, Fel97b, Lee94b, Pan94]. Shell [Aki01, BC99, Bay00, Dul96b, Fie92, GT95, Hil96c, Iye97, Jon98a, Raa01, Sch97a, Sch99a, Sch00a, Tan98b, Tes96, Vel94c, Wad96]. Shell-Based [Vel94c]. Shells [Sch02b]. Short [Dea93]. Side [WM97]. Signaling [McC01f]. Signals [Lee95b]. Simple [Bol00, Dul96b, End97, Fel96b, Fel98, Gra96, Hol97, Mor96b, Sto92, Wei01a]. Simple-Minded [Gra96]. Simplifying [Ane96, Tay01, Vog00a]. Simulating [Mor96c]. Single [Dav00b, Sch97a]. Single-User [Dav00b]. Site [Ane96, Don96d, PK99, Sal96, Sch99g, Swe00a, Wel98]. Site-Wide [Ane96]. Sites [Jay98, Kla00, Ped02]. SLAMming [Qui98]. SLIP [Har94a, MD94]. Slow [RL98]. SM2UR [Gra96]. Small [Mur01]. Smart [Bru94]. Smarter [Rez94d]. SMIT [Car01b]. Smurfing [Tho98b]. Sniffing [Mec99b]. SNMP [Qui01a]. SNMPv3 [Dav00d, Dav00e]. Snort [McC00j, Wes00a, Wes01]. Sockets [Joh96a].

SOCKS [Fel98, Gan95, Gan98]. SOCKSLIB.PL [Gan98]. Software [Gag98, Haz97, Hun92a, Lee94b, Moo97, Vad98]. Software-based [Haz97]. Solaris [BR99, Coo01, Dia00, Dub00, End98, Gaf99, Gal01c, Gal01d, Gal01b, Gal02b, Gal02e, Gal02c, Gal02a, Gol01, GN01b, Got01, How01, Kra01, McC00d, Oli01, OY01, Qui01b, Qui01a, Sch00d, Tee00, vdW00, WM00, Wen01, Won97, Zwi99a]. Solaris^{T M} [Gal01a, Gal02f, Gal02e, Gal02c, Gal02d, Kno02, LA01]. Solution [Car97, Hag01, MD94, McC00b, Par01, Sch00c, Str02, Wit00]. Solutions [Ano98-50, Ano98-51, Ano98-52, Ano98-53]. Solving [Bry97b]. Some [Fie92]. Sony [Swi98]. Sort [Wes01]. Source [Ano92q, Ano92r, Ano92s, Ano93-27, Ano93-28, Ano93-29, Ano93-30, Ano93-31, Ano93-32, Ano94-31, Ano94-33, Ano94-29, Ano94-30, Ano94-32, Cha01, DS00, Gag02, Luc99, Reh01b, Rez95b]. Sources [SD96]. Space [Con97a, Lei96, Mil00a]. Spam [Leo01, Sch99d, War99]. Spamivore [Leo01]. Spammers [Sch99c]. Spamming [Sch99b]. Spatchula [Won97]. Speaking [Don97a]. Speed [Bar97l, Gar96, Hel99b, Rez93a]. Speeds [McC00a]. Spidering [Vog01]. Spoofing [Don98c]. Spooler [Zol92b]. Spoolers [Bro95]. SQL [GT95]. SSH [Str02]. SSL [Don99c]. Stack [Gut96]. Standard [Abl93a, Dul94b, Len94]. Standby [Wat01a]. Starfire [Rug00]. Start [How01]. Startup [Hil96c]. stat [Isa94b]. Stateful [Nap01]. States [Dav99b]. Statistics [Dav97, Mic98]. Step [Ada97, Agu01]. Steps [Hil98, Qui01b]. Stopping [Isa92, Jon02]. Storage [Cer01a, EK00, Pur01, Sch97c, Sch00d, Sul00, Tho99, Zwi99c]. Strategies [De 97b, McC99d]. Streamlining [Ben98]. Study [Kie99b, Soh01, Wei92b]. Style [Zol92b]. Subnets [Bak95]. Subnetting [Har95c]. Subscription [SN95].

Subscription-Based [SN95]. Subsystem [Oku98]. Successful [Got01]. sudo [Pip93]. suFirewall [Cox94]. SUID [Aki01, Tot00]. SUID/SGID [Tot00]. SUIDed [Rac00]. sukill [Isa92]. Sun [Gal02b, Gal02f, Geb98, Kom00a, Raj01]. Sun/Cobalt [Gal02f]. Super [Pip93]. Super-User [Pip93]. Support [Bye99, Cou92, Dia00, Kin95, Swe99]. Supporting [Cer97, DeJ02]. Surveillance [Ric95]. Survey [Ano97-40, Bar97l, Kol99]. Surviving [Con96, SL95b]. SVR4 [Luc00b]. SVR4/Linux [Luc00b]. Switches [Zha98]. Switching [Hel99d, Zha99b]. Sybaccess [Sch96]. Sybase [Che96a, Gen94c, Gen94d, Gen95, Sch96]. sybperl [Mel96]. Symbolic [KR94, Qui98]. Symlink [Sch99j]. Sync [Ols01a]. Synchronization [Gau98, Vel98b]. Synchronizer [Sch95]. Sys [Bun96, Plo97, Ano96-47, Ano98-68, Ano98-69, Bar98m, Bec00, Isa96a, Plo98]. Sysadmin [Ano97-27]. sysctl [Luc00c]. syslog [Ank01a, Ank01b, Ank01c, Ank01d, Ank01e, Ank01f, Ank01g, Ank01i, Ank01h, Ank01j, Ank02a, Ank02b, Ank02c, Bar99h, Bar99i, Wei98, Woo94]. syslog.conf [Hil96a]. SysMon [Ess98b]. System [Bri94, Cas01a, Dav96, Dij00, Dil02, DS00, Dyk96, Gar97, Gar00, Gen92, Gom94, Gra96, Hal02, Har92d, Har94c, Hil93, Hu98a, Isa93b, Isa94b, Kin95, Kla98, Kol99, Lee95a, Lym01a, Mac97, MD94, Mag95, McC99a, Muh97a, Naz01, Pre01e, Rez94f, Ric96a, Smi98, SN95, Tar96, Tus95, VW00, Vel94a, Vog00b, Wes00a, Wyn97, Zol92d, Ber98b, Cer97]. System-Disk [Gar00]. Systems [App00, Bau96, Cay94b, Cla94, Ess98b, FPA99, Gag00a, Gal02b, Gar00, Hun92c, Iba97, Kue01, Reh01b, Sar02, Wei92b, WC97, van97a].

Table [Ano00-40]. Tag [Sch00b]. Taint [Sch00j]. Taking [Dul96b]. Tales [WF01].

Tamper [Mer98a]. Tape [Abl93c, Bar98m, Dav00b, Swi98, Vel95b]. Tape-to-Tape [Vel95b]. tar [Rea93, Str02]. Tasks [Bun96, Plo97]. Tcl [Hal97, Tar96]. Tcl/Tk [Hal97, Tar96]. TCP [Dav99b, Dav01a, Joh97, Ols00, Sch95, Sto95, Tis98]. TCP/IP [Joh96c, Joh97, Sch95, Sto95, Tis98]. tcp wrapper [Bus99b, Bus99a]. Technical [Nol98, Raj01]. Techniques [Luc00a, Mor96c, Pra01a, Pre98]. Technologies [Vel95a]. Technology [Swi98, Zha99a]. Televaulting [Hel98d]. telnet [Joh96c]. Temperature [Kom00a]. Template [Sch00h]. Temporary [Ped02, San93a]. Term [Vel95a, Bry00]. termcap [Tho95]. termcap/terminfo [Tho95]. Terminal [Bro93]. Terminals [Bro92a, Bro92b]. terminfo [Rez92b, Rez95b]. Terms [Don97a]. Terrible [Sch01d]. Testing [Bro92b]. Text [GN01a]. Their [Pre00b]. Thieves [Jon02]. Thing [Sch01d]. Things [Sch01e, Sch02d]. Third [Cay94a]. Third-Party [Cay94a]. Thought [Plo00]. Threaded [Mag96]. Tiered [Nas98a, Nas98b]. Tiger [Wes00b]. Time [Cer01b, FS95, Gau98, Har95c, Joh95, Reh01b, Rez94e, Vel98b]. Times [Dea93, MS99]. Timing [Rez93c]. Tips [Don96a, Jac00, San99]. TITAN [FPA99]. Tivoli [Cer01a]. Tomcat [Bus01]. Tool [Bro93, Eve95, Gag99b, Hel97b, Mec99a, Pol01, Pre01e, Rau99, RL98, Wil01b]. Toolkit [Bec00, Dul00a]. Tools [Bah01, Cha01, Cla01, Cou93c, Don99b, Dub01, Dul94b, Hel98g, Lei95b, Len94, Luc00a, WT00, Wes01, WC97]. Trace [Swe00b]. Tracking [Bri94, Ess96, She02]. Traditional [BL99]. Traffic [She01]. Training [Ano01-58, Ano01-59, Ano01-60, Ano02m, Ano02n]. Transfer [Bla99, Cal01]. Transition [End98]. Translator [McC00g]. Transparent [Fel94, Reh99]. Trilogy [Zol92d]. Tripwire [Kha01]. Trojan [Abl92a]. Troubleshooting [Fel97c, Kon93, Pod01, San99]. Trust [Kle97]. tty [Pet95]. Tuning [AF00, Bec99, Got01, Lin93b, Nas97a, Ric96a]. Tunnels [Don00]. Turning [Rez93d]. Two [Bak95, Cou93c]. Types [Sal01].

U.S. [FS95]. u386mon [Har92b]. ufs [Cla94]. ufsdump [DW01]. Ugly [McC00c]. Uncommon [Bar99j]. Understanding [Dul95a, Dul95b, Hel98e, Hil96a, KP99, Pre01h]. Unexpected [Isa93a]. Uni [Bur01a]. Unidentified [Lee95c]. UNIX [Wid00, van97b, Abl93a, Che96b, Dav98b, Dav00c, Dil02, Don95b, Don96a, Don97a, Don98d, Dul94b, Dul96a, Fel94, Gau98, GS00, Gol98, Har92c, Har92d, Har94c, Har01a, Har01b, HD97, Haz99b, Hel97c, Hil93, Hil98, Hof97, Hol97, Hun92a, Hun92c, Hun92d, Joh95, Joh96a, Joh96b, Jon98b, Kni92, Kos96, Kuo98, Lei95a, Len94, Liu96, Mac97, Man01b, Mar01b, May97, Pan94, Pru02, Rez94e, Rez95d, Ric96a, Rie93, Rie94, San93b, Sch93, SL95a, SB96, Sch97a, Sef93, Sei00c, SA95, Sto95, Wee99, Wei92b, WF01, Wit01]. UNIX/NT [Gau98]. UNIX/PC [Fel94]. UNIX/Windows [Don95b]. unixdoc [Mar01b]. UnixReview.com [Sch02a]. Unraveled [Nel99]. Unruly [Isa92]. Update [CB96]. Upgrading [Con99, Gal02c]. Uptime [Gra96]. Usability [Bry00]. Usage [End97, Man01a, McC01d, Mic98, Wil93]. Use [Mer99, Sus00]. Useful [Fie92, Pra01b, Pra01c]. USENET [Wei92a, Wei92b, Dav01b, Hu98b]. Usenix [Nol98]. User [Bau96, Bel98, Bun96, Bye99, Con96, Dav00b, Dav00d, Don96e, Gar98a, Har93f, Len94, Pip93, Rez94a, Sal97b, Zin92a, Zin92b, Zin92c, Zin93a, Zin93b, Zin93c, Zin93d, Zin93e, Zin93f, Zin94a, Zin94b, Zin94c, Zin94d, Zin94e, Zin94f, Zin95a, Zin95b, Zin95c, Zin95d, Zin95e, Zin95f, Zin96a, Zin96b, Zin96c, Zin96d, Zin96e, Zin96f, Zin96g, Zin96h, Zin96i, Zin96j, Zin96k, Zin96l, Zin97a, Zin97b, Zin97c, Zin97d, Zin97e, Zin97f, Zin97g, Zin97h, Zin97i, Zin97j, Zin97k, Zin98a, Zin98b, Zin98c, Zin98d, Zin98e, Zin98f, Zin98g, Zin98h, Zin98i, Zin98j, Zin98k, Zin98l, Zin99a, Zin99b, Zin99c, Zin99d, Zin99e, Zin99f, Zin99g, Zin99h, Zin99i, Zin99j, Zin99k, Zin99l, Zin00a, Zin00b, Zin00c, Zin00d, Zin00e, Zin00f, Zin00g, Zin00h, Zin00i, Zin00j, Zin00k]. User [Zin00l, Zin00m, Zin01a, Zin01b, Zin01c, Zin01d, Zin01e, Zin01f, Zin01g, Zin01h, Zin01i, Zin01j, Zin01k, Zin01l, Zol93d]. Users [Bun97, Che97, Dul94a, Hel98a, Hil95, Rez93c]. Using [BmC01, Bun97, Bus99d, Cin01, Cla94, Coo01, DW01, Dil02, Don96h, Don98e, LE97, Fel96c, Fer01a, Gag98, Gan98, GT95, Har95a, Hel98f, Her02, Hu98a, Isa99, Jen01, Joh96a, Kuo98, LB96, Lin99, Lym01d, New02, Nie00, Oli01, Ped02, Pre01b, Pre01c, Rez92c, Rez93e, Rez93f, Sar01, Sch93, Sch00b, Tar96, Wil01b, Wit01, Woo94, Woo96, Zwi99a]. Utilities [Kuo98, Mor95, Wat01b]. Utility [Cox95, Rez95d, San93b, Zol93a]. Utilization [Ber98a]. Utilizing [Str02]. UUCP [Don95a, Gom94, Har93b, Har93g, Har93f, McD95]. UUCP/PPP [Don95a]. UX [Kle97, Rez95h].

V [Cer97, Lee95a]. v6.1 [van97b].

Validating [Sch97b]. Variable [Bry00].

Vendor [Pod01]. Verification

[Bla99, Cas01b, Dav00b, Ric96b]. Verifying [MS99, Vel94a]. Version [Har93b, Har94b, Ric96b]. Versus [Jon99]. vGen [Pin97]. vi [Har95d, Sch00b]. via [Fel94, Hel96c, McD95]. Viagra [O’G01]. View [Dav00e, DeJ02, McC02, Swe99]. Virtual [Gre01, Jay98, McC00k, Nap02, Rib02, Tho98c, Wha98]. vLANs [Hel98f]. vmstat [Gen94b]. VMWare [Wil01b].

Voice [Hel00c]. vol [Ano97-27]. Volume [Gen94a, Reh00]. Volumes [San98]. VPN [Wit00]. VPNs [FM99]. VTun [Don00]. Vulnerability [BmC01, Fre02].

Waldo [Agu99]. Want [Sch00f]. Warning [Gom94]. Waste [Sch01d]. Watchdog [Isa93b, Isa96a]. Watching [Har93a]. Way [Bus01, Gag02, Isa96b, McC99e, Sch00f]. Web [How01, Bah01, Bal99, Bay00, Bro97, Bus99a, Coo96, Dav00c, De 97b, De 98a, Don96d, Don96f, Don96h, Don97c, Fel96a, Gag99a, Gar02, GW00, Hel96a, Isa98, Jay98, Kla00, Lib02, Lym02, Mac97, Mag96, McC99c, Mel96, Mic98, Mob98, Ols01a, Orw01, Ped02, Ram00, Rau99, Ric97, Sal96, Sha97, Smi98, Soh01, SB99, Tes96, Vog00a, Vog02, Wel98, Wid00, Wol00, Won02, Zwi98, Kom00b]. Web-Based [Don96f, Mag96, Bro97, Lym02, Mac97, Rau99, Sha97, Smi98, Wol00, GW00, Ric97]. Web-centric [De 97b]. Web-Enabled [Isa98, Lib02]. Web-Enabling [Gag99a]. Web-Exclusive [Bus99a]. Web-Interface [Bay00]. Web/RCS [Dav00c]. Webifying [Wee99]. Weblog [Jen01]. Weeding [Rez94f]. Week [Rez93g]. Where [Abl93c, Har92e]. Which [Rez93g, RB01a, RB01b]. Who [Don98c]. Wide [Ane96]. Window [Bro96]. WindowMaker [Tay01]. Windows [SL95a, Dia00, Fla97, Gag02, Gol01, Hel97e, Jon98b, Kos96, Oku98, Pan94, Sei01, Tis98, Wit01]. WINE [Gag02]. Wireless [DB02, Dub02, Ris00]. Wisdom [Bus99c]. Wish [Bay00]. without [CC01, Mil00a]. Words [Sch99i]. Work [Abl92b, Har92c, Isa96b, Rac00]. Workgroups [Pan94]. Working [Ols01b]. Workstation [Dia00]. Workstations [Lei95a]. World [McC00l, McK98d, Sch98]. WPM [Wol00]. Wrapper [Ols00]. Wrappers [Hal97, Rez95d]. Writing [Don97d, Rau99, Tan98b]. WTMP [Agu01]. WWW [Kin95, Sal98, Tes96]. WWW-Based [Kin95].

X [Bro96, Lin93b, Wei95]. XML [Orw01]. Xni [Gag99b]. xterm [Tho95].

Y2K [Kie99a].

Zebra [Sch01a]. Zombie [Wei93].

References

Ables:1992:BTH

[Abl92a] R. King Ables. Beware the Trojan Horse. Sys Admin: The Journal for UNIX Systems Administrators, 1(1):43–??, May/ June 1992. CODEN SYADE7. ISSN 1061-2688.

Ables:1992:PCW

[Abl92b] R. King Ables. Putting cron to work for you. Sys Admin: The Journal for UNIX Systems Administrators, 1(4):72– ??, November/December 1992. CODEN SYADE7. ISSN 1061-2688.

Ables:1992:SEM

[Abl92c] R. King Ables. Security and electronic mail. Sys Admin: The Journal for UNIX Systems Administrators, 1(2):??, July/ August 1992. CODEN SYADE7. ISSN 1061-2688.

Ables:1993:BSU

[Abl93a] R. King Ables. Backups with standard UNIX commands. Sys Admin: The Journal for UNIX Systems Administrators, 2(5):

6–??, September/October 1993. CODEN SYADE7. ISSN 1061-2688.

Ables:1993:DLI

[Abl93b] R. King Ables. Distributing local information with NIS. Sys Admin: The Journal for UNIX Systems Administrators, 2(4): 13–??, July/August 1993. CO-DEN SYADE7. ISSN 1061-2688.

Ables:1993:WDY

[Abl93c] R. King Ables. Where did you get that tape? Sys Admin: The Journal for UNIX Systems Administrators, 2(2): 41–??, March/April 1993. CO-DEN SYADE7. ISSN 1061-2688.

Adamopoulos:1997:SFS

[Ada97] Yiorgos Adamopoulos. Samba — the first step. Sys Admin: The Journal for UNIX Systems Administrators, 6(5):21–??, May 1997. CODEN SYADE7. ISSN 1061-2688.

Anderson:2000:PAP

[AF00] Adam Anderson and Dave Felps.

Protocol analysis for performance tuning. Sys Admin: The Journal for UNIX Systems Administrators, 9(12):20, 22, 24, December 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Aguiar:1999:LFE

[Agu99] Joe Aguiar. Log file event management with Waldo. Sys Admin: The Journal for UNIX Systems Administrators, 8(11): 59–60, 62–63, 65, November 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Aguiar:2000:CPE

[Agu00] Joe Aguiar. Cross platform event reporting. Sys Admin: The Journal for UNIX Systems Administrators, 9(9):33– 34, 36, September 2000. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Aguiar:2001:WAS

[Agu01] Joe Aguiar. WTMP: Another step toward centralizing event management. Sys Admin: The Journal for UNIX Systems Administrators, 10(7):49–51, July 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Akin:2001:DSS

[Aki01] Thomas Akin. Dangers of SUID shell scripts. Sys Admin: The Journal for UNIX Systems Administrators, 10(6):8, 10, 12, June 2001. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Ali:1999:FBS

[Ali99] Syed Ali. Freeware-based security. Sys Admin: The Journal for UNIX Systems Administrators, 8(1):39–42, 43–44, January 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Allen:1996:MAN

[All96] Jay Allen. Migrating to an automounted network filesystem. Sys Admin: The Journal for UNIX Systems Administrators, 5(1):19–??, January 1996. CO-DEN SYADE7. ISSN 1061-2688.

Allison:1998:ICN

[All98] Bridget Allison. Integrating CIFS and NFS file sharing. Sys Admin: The Journal for UNIX Systems Administrators, 7(3): 34, 36, 38–40, March 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Andregg:1999:ALO

[And99] Bryan C. Andregg. The all Linux office. Sys Admin: The Journal for UNIX Systems Administrators, 8(6):33–38, June 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Anderson:2001:WDI

[And01] Adam Anderson. What I did instead of buying a SAN. Sys Admin: The Journal for UNIX Systems Administrators, 10(9): 8, 10, 12, 14, September 2001. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Anello:1996:SSB

[Ane96] Benjamin J. Anello. Simplifying site-wide backups. Sys Admin: The Journal for UNIX Systems Administrators, 5(7):24–??, July 1996. CODEN SYADE7. ISSN 1061-2688.

Ankerholz:1999:EFS

[Ank99a] Amber Ankerholz. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 8(10):6, October 1999. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Ankerholz:1999:EF

[Ank99b] Amber Ankerholz. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 8(12):6, December 1999. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Ankerholz:2000:EFa

[Ank00a] Amber Ankerholz. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 9(1):6, January 2000. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Ankerholz:2000:EFb

[Ank00b] Amber Ankerholz. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 9(2):6, February 2000. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Ankerholz:2000:EFc

[Ank00c] Amber Ankerholz. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 9(3):6, March 2000. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Ankerholz:2000:EFd

[Ank00d] Amber Ankerholz. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 9(4):6, April 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Ankerholz:2000:EFe

[Ank00e] Amber Ankerholz. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 9(5):6, May 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Ankerholz:2000:EFf

[Ank00f] Amber Ankerholz. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 9(6):6, June 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Ankerholz:2000:EFg

[Ank00g] Amber Ankerholz. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 9(7):??, July 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Ankerholz:2000:EFh

[Ank00h] Amber Ankerholz. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 9(8):6, August 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Ankerholz:2000:EFi

[Ank00i] Amber Ankerholz. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 9(9):6, September 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Ankerholz:2000:EFj

[Ank00j] Amber Ankerholz. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 9(10):6, October 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Ankerholz:2000:EFk

[Ank00k] Amber Ankerholz. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 9(11):6, November 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Ankerholz:2000:EFl

[Ank00l] Amber Ankerholz. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 9(11S):6, November 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Ankerholz:2000:EFm

[Ank00m] Amber Ankerholz. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 9(12):6, December 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Ankerholz:2001:EFSa

[Ank01a] Amber Ankerholz. Editor’s forum: syslog. Sys Admin: The Journal for UNIX Systems Administrators, 10(1):6, January 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Ankerholz:2001:EFSb

[Ank01b] Amber Ankerholz. Editor’s forum: syslog. Sys Admin: The Journal for UNIX Systems Administrators, 10(2):6, February 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Ankerholz:2001:EFSc

[Ank01c] Amber Ankerholz. Editor’s forum: syslog. Sys Admin: The Journal for UNIX Systems Administrators, 10(3):6, March 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Ankerholz:2001:EFSd

[Ank01d] Amber Ankerholz. Editor’s forum: syslog. Sys Admin: The Journal for UNIX Systems Administrators, 10(4):6, April 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Ankerholz:2001:EFSe

[Ank01e] Amber Ankerholz. Editor’s forum: syslog. Sys Admin: The Journal for UNIX Systems Administrators, 10(5):6, May 2001. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Ankerholz:2001:EFSf

[Ank01f] Amber Ankerholz. Editor’s forum: syslog. Sys Admin: The Journal for UNIX Systems Administrators, 10(6):6, June 2001. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Ankerholz:2001:EFSg

[Ank01g] Amber Ankerholz. Editor’s forum: syslog. Sys Admin: The Journal for UNIX Systems Administrators, 10(7):6, July 2001. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Ankerholz:2001:ESF

[Ank01h] Amber Ankerholz. Editor’s forum: syslog. Sys Admin: The Journal for UNIX Systems Administrators, 10(8):6, August 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Ankerholz:2001:ES

[Ank01i] Amber Ankerholz. Editor’s syslog. Sys Admin: The Journal for UNIX Systems Administrators, 10(9):6, September 2001. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Ankerholz:2001:S

[Ank01j] Amber Ankerholz. syslog. Sys Admin: The Journal for UNIX Systems Administrators, 10(10): 6, October 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Ward:2001:EF

[Ank01k] Anker Ankerholz. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 10(6s):6, June 2001. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Ankerholz:2002:EFSa

[Ank02a] Amber Ankerholz. Editor’s forum: syslog. Sys Admin: The Journal for UNIX Systems Administrators, 11(3):6, March 2002. CODEN SYADE7. ISSN 1061-2688.

Ankerholz:2002:EFSb

[Ank02b] Amber Ankerholz. Editor’s forum: syslog. Sys Admin: The Journal for UNIX Systems Administrators, 11(4):6, April 2002. CODEN SYADE7. ISSN 1061-2688.

Ankerholz:2002:S

[Ank02c] Amber Ankerholz. syslog. Sys Admin: The Journal for UNIX Systems Administrators, 11(5): 6, May 2002. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1992:AIa

[Ano92a]  Anonymous.  Advertiser in-

dex. Sys Admin: The Journal for UNIX Systems Administrators, 1(1):64–??, May/June 1992. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1992:AIb

[Ano92b] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 1(2):??, July/August 1992. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1992:AIc

[Ano92c] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 1(3):72–??, September/October 1992. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1992:AId

[Ano92d]  Anonymous.  Advertiser in-

dex. Sys Admin: The Journal for UNIX Systems Administrators, 1(4):80–??, November/December 1992. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1992:CPa

[Ano92e] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 1(1): 50–??, May/June 1992. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1992:CPb

[Ano92f] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 1(2):??, July/August 1992. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1992:CPc

[Ano92g] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 1(3):64–??, September/October 1992. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1992:CPd

[Ano92h]  Anonymous.  Call for pa-

pers. Sys Admin: The Journal for UNIX Systems Administrators, 1(4):77–??, November/December 1992. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1992:NMa

[Ano92i] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 1(1): 63–??, May/June 1992. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1992:NMb

[Ano92j] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 1(2):??, July/August 1992. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1992:NMc

[Ano92k] Anonymous. New messages.

Sys Admin: The Journal for UNIX Systems Administrators, 1(3):70–??, September/October 1992. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1992:NMd

[Ano92l] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 1(4): 67–??, November/December 1992. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1992:NPa

[Ano92m] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 1(1): 61–??, May/June 1992. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1992:NPb

[Ano92n] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 1(2):??, July/August 1992. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1992:NPc

[Ano92o] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 1(3):66–??, September/October 1992. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1992:NPd

[Ano92p] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 1(4): 58–??, November/December 1992. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1992:SCAa

[Ano92q] Anonymous. Source code availability. Sys Admin: The Journal for UNIX Systems Administrators, 1(2):??, July/August 1992. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1992:SCAb

[Ano92r] Anonymous. Source code availability. Sys Admin: The Journal for UNIX Systems Administrators, 1(3):57–??, September/October 1992. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1992:SCAc

[Ano92s] Anonymous. Source code availability. Sys Admin: The Journal for UNIX Systems Administrators, 1(4):46–??, November/December 1992. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:AIb

[Ano93a]  Anonymous.  Advertiser in-

dex. Sys Admin: The Journal for UNIX Systems Administrators, 2(1):88–??, January/February 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:AId

[Ano93b] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 2(2):112–??, March/April 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:AIf

[Ano93c]  Anonymous.  Advertiser in-

dex. Sys Admin: The Journal for UNIX Systems Administrators, 2(3):96–??, May/June 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:AIh

[Ano93d] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 2(4):104–??, July/August 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:AIk

[Ano93e]  Anonymous.  Advertiser in-

dex. Sys Admin: The Journal for UNIX Systems Administrators, 2(5):112–??, September/October 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:AIn

[Ano93f]  Anonymous.  Advertiser in-

dex. Sys Admin: The Journal for UNIX Systems Administrators, 2(6):112–??, November/December 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:CPa

[Ano93g] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 2(1): 71–??, January/February 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:CPb

[Ano93h] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 2(2):101–??, March/April 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:CPc

[Ano93i] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 2(3): 74–??, May/June 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:CPd

[Ano93j] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 2(4):86–??, July/August 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:CPe

[Ano93k] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 2(5):45–??, September/October 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:CPf

[Ano93l]  Anonymous.  Call for pa-

pers. Sys Admin: The Journal for UNIX Systems Administrators, 2(6):91–??, November/December 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:AIi

[Ano93m] Anonymous. InstantInfo. Sys Admin: The Journal for UNIX Systems Administrators, 2(5):71–??, September/October 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:AIm

[Ano93n] Anonymous. InstantInfo. Sys Admin: The Journal for UNIX Systems Administrators, 2(6): 101–??, November/December 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:NMa

[Ano93o] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 2(1): 86–??, January/February 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:NMb

[Ano93p] Anonymous. New messages.

Sys Admin: The Journal for UNIX Systems Administrators, 2(2):110–??, March/April 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:NMc

[Ano93q] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 2(3): 94–??, May/June 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:NMd

[Ano93r] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 2(4):90–??, July/August 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:NMe

[Ano93s] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 2(5):110–??, September/October 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:NMf

[Ano93t] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 2(6): 110–??, November/December 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:NPa

[Ano93u] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 2(1): 82–??, January/February 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:NPb

[Ano93v] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 2(2):102–??, March/April 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:NPc

[Ano93w] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 2(3): 87–??, May/June 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:NPd

[Ano93x] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 2(4): 97–??, July/August 1993. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1993:NPe

[Ano93y] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 2(5):103–??, September/October 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:NPf

[Ano93z] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 2(6): 103–??, November/December 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:AIa

[Ano93-27] Anonymous. Source code availability. Sys Admin: The Journal for UNIX Systems Administrators, 2(1):78–??, January/February 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:AIc

[Ano93-28] Anonymous. Source code availability. Sys Admin: The Journal for UNIX Systems Administrators, 2(2):100–??, March/April 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:AIe

[Ano93-29] Anonymous. Source code availability. Sys Admin: The Journal for UNIX Systems Administrators, 2(3):85–??, May/June 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:AIg

[Ano93-30] Anonymous. Source code availability. Sys Admin: The Journal for UNIX Systems Administrators, 2(4):96–??, July/August 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:AIj

[Ano93-31] Anonymous. Source code availability. Sys Admin: The Journal for UNIX Systems Administrators, 2(5):102–??, September/October 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1993:AIl

[Ano93-32] Anonymous. Source code availability. Sys Admin: The Journal for UNIX Systems Administrators, 2(6):99–??, November/December 1993. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:AIa

[Ano94a] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 3(1):128–??, January/February 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:AIb

[Ano94b] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 3(2):120–??, March/April 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:AIc

[Ano94c]  Anonymous.  Advertiser in-

dex. Sys Admin: The Journal for UNIX Systems Administrators, 3(3):120–??, May/June 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:AId

[Ano94d] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 3(4):120–??, July/August 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:AIe

[Ano94e] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 3(5):96–97, September/October 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:AIf

[Ano94f]  Anonymous.  Advertiser in-

dex. Sys Admin: The Journal for UNIX Systems Administrators, 3(6):96–97, November/December 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:CPa

[Ano94g] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 3(1):109–??, January/February 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:CPb

[Ano94h] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 3(2):99–??, March/April 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:CPc

[Ano94i] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 3(3): 108–??, May/June 1994. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1994:CPd

[Ano94j] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 3(4):83–??, July/August 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:CPe

[Ano94k]  Anonymous.  Call for pa-

pers. Sys Admin: The Journal for UNIX Systems Administrators, 3(5):112–??, September/October 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:CPf

[Ano94l]  Anonymous.  Call for pa-

pers. Sys Admin: The Journal for UNIX Systems Administrators, 3(6):119–??, November/December 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:Ia

[Ano94m] Anonymous. InstantInfo. Sys Admin: The Journal for UNIX Systems Administrators, 3(1): 113–??, January/February 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:Ib

[Ano94n] Anonymous. InstantInfo. Sys Admin: The Journal for UNIX Systems Administrators, 3(2): 117–??, March/April 1994. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1994:Ic

[Ano94o] Anonymous. InstantInfo. Sys Admin: The Journal for UNIX Systems Administrators, 3(3): 107–??, May/June 1994. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1994:Id

[Ano94p] Anonymous. InstantInfo. Sys Admin: The Journal for UNIX Systems Administrators, 3(4): 89–??, July/August 1994. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1994:NMb

[Ano94q] Anonymous. New messages.

Sys Admin: The Journal for UNIX Systems Administrators, 3(1):124–??, January/February 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:NMc

[Ano94r] Anonymous. New messages.

Sys Admin: The Journal for UNIX Systems Administrators, 3(2):119–??, March/April 1994. CODEN SYADE7. ISSN 1061-

Anonymous:1994:NMe

[Ano94s] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 3(3): 114–??, May/June 1994. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1994:NMg

[Ano94t] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 3(4):118–??, July/August 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:NMh

[Ano94u] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 3(5):127–??, September/October 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:NMi

[Ano94v] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 3(6): 125–127, November/December 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:NPa

[Ano94w] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 3(1): 117–??, January/February 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:NPb

[Ano94x] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 3(2):112–??, March/April 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:NPc

[Ano94y] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 3(3): 102–??, May/June 1994. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1994:NPd

[Ano94z] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 3(4):113–??, July/August 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:NPf

[Ano94-27] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 3(5):122–??, September/October 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:NPg

[Ano94-28] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 3(6):122–124, 128, November/ December 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:NMf

[Ano94-29] Anonymous. Online source code.

Sys Admin: The Journal for UNIX Systems Administrators, 3(4):109–??, July/August 1994. CODEN SYADE7. ISSN 1061-

Anonymous:1994:NPe

[Ano94-30] Anonymous. Online source code. Sys Admin: The Journal for UNIX Systems Administrators, 3(5):118–??, September/October 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:NMa

[Ano94-31] Anonymous. Source code availability. Sys Admin: The Journal for UNIX Systems Administrators, 3(1):116–??, January/February 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:SCA

[Ano94-32] Anonymous. Source code availability. Sys Admin: The Journal for UNIX Systems Administrators, 3(2):108–??, March/April 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1994:NMd

[Ano94-33] Anonymous. Source code availability. Sys Admin: The Journal for UNIX Systems Administrators, 3(3):112–??, May/June 1994. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1995:AIa

[Ano95a]  Anonymous.  Advertiser in-

dex. Sys Admin: The Journal for UNIX Systems Administrators, 4(2):104–105, March/April 1995. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1995:AIb

[Ano95b] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 4(3):104–105, May/June 1995. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1995:AIc

[Ano95c]  Anonymous.  Advertiser in-

dex. Sys Admin: The Journal for UNIX Systems Administrators, 4(4):104–105, July/August 1995. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1995:AId

[Ano95d]  Anonymous.  Advertiser in-

dex. Sys Admin: The Journal for UNIX Systems Administrators, 4(6):88–89, November/December 1995. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1995:CPa

[Ano95e] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 4(2):123–??, March/April 1995. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1995:CPb

[Ano95f] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 4(3): 119–??, May/June 1995. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1995:CPc

[Ano95g] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 4(4):127–??, July/August 1995. CODEN SYADE7. ISSN 1061-

Anonymous:1995:CPd

[Ano95h]  Anonymous.  Call for pa-

pers. Sys Admin: The Journal for UNIX Systems Administrators, 4(6):86–??, November/December 1995. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1995:NMa

[Ano95i] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 4(2):136–??, March/April 1995. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1995:NMb

[Ano95j] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 4(3): 134–??, May/June 1995. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1995:NMc

[Ano95k] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 4(4):135–??, July/August 1995. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1995:NPa

[Ano95l] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 4(2):130–??, March/April 1995. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1995:NPb

[Ano95m] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 4(3):

128–??, May/June 1995. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1995:NPc

[Ano95n] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 4(4):130–??, July/August 1995. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1995:NPd

[Ano95o] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 4(6): 101–??, November/December 1995. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:AIa

[Ano96a] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 5(1):88–??, January 1996. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1996:AIb

[Ano96b] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 5(2):72–??, February 1996. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1996:AIc

[Ano96c] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 5(5):64–??, May 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:AId

[Ano96d] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 5(6):56–??, June 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:AIe

[Ano96e] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 5(7):48–??, July 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:AIf

[Ano96f]  Anonymous.  Advertiser in-

dex. Sys Admin: The Journal for UNIX Systems Administrators, 5(9):72–??, September 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:AIg

[Ano96g] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 5(10):80–??, October 1996. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1996:AIh

[Ano96h] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 5(11):64, November 1996. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1996:AIi

[Ano96i] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 5(11):64, November 1996. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1996:AIj

[Ano96j] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 5(12):72, December 1996. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1996:CPa

[Ano96k] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 5(1):94–??, January 1996. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1996:CPb

[Ano96l] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 5(2):86–??, February 1996. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1996:CPc

[Ano96m] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 5(5):74–??, May 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:CPd

[Ano96n] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 5(6):83–??, June 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:CPe

[Ano96o] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 5(7):78–??, July 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:CPf

[Ano96p] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 5(9): 78–??, September 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:CPh

[Ano96q] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 5(10): 63–??, October 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:CPg

[Ano96r] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 5(10): 93–??, October 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:CPi

[Ano96s] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 5(11): 63, November 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:CPj

[Ano96t] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 5(12): 84, December 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:COa

[Ano96u] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 5(5):66–??, May 1996. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1996:COb

[Ano96v] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 5(6):72–??, June 1996. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1996:COc

[Ano96w] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 5(7):72–??, July 1996. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1996:COd

[Ano96x] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 5(9):74–??, September 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:COe

[Ano96y] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 5(10):82–??, October 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:COf

[Ano96z] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 5(11):73–75, November 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:COg

[Ano96-27] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 5(11):73–75, November 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:COh

[Ano96-28] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 5(12):83, December 1996.

CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:NMa

[Ano96-29] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 5(1):100–??, January 1996. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1996:NMb

[Ano96-30] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 5(2): 103–??, February 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:NMc

[Ano96-31] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 5(5):80–??, May 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:NMd

[Ano96-32] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 5(6):88–??, June 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:NMe

[Ano96-33] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 5(9): 88–??, September 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:NMg

[Ano96-34] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 5(10): 80, October 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:NMf

[Ano96-35] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 5(10): 96–??, October 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:NMh

[Ano96-36] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 5(11): 80, November 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:NPa

[Ano96-37] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 5(1):101–??, January 1996. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1996:NPb

[Ano96-38] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 5(2):76–??, February 1996. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1996:NPc

[Ano96-39] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 5(5):70–??, May 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:NPd

[Ano96-40] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 5(6):84–??, June 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:NPe

[Ano96-41] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 5(7):63–??, July 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:NPf

[Ano96-42] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 5(9): 84–??, September 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:NPg

[Ano96-43] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 5(10): 76–??, October 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:NPh

[Ano96-44] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 5(10): 76–79, October 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:NPi

[Ano96-45] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 5(11): 76–79, November 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:NPj

[Ano96-46] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 5(12): 85–88, December 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1996:SAC

[Ano96-47] Anonymous. Sys Admin contents: 1992–1995. Sys Admin: The Journal for UNIX Systems Administrators, 5(2): 91–102, February 1996. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:AIa

[Ano97a] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 6(4):56–??, April 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:AIb

[Ano97b] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 6(5):72–??, May 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:AIc

[Ano97c] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 6(6):64, June 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:AIe

[Ano97d] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 6(8):48, August 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:AIf

[Ano97e] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 6(9):48, September 1997. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1997:AIg

[Ano97f] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 6(10):64, October 1997. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1997:AIh

[Ano97g] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 6(11):64, November 1997. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1997:AIi

[Ano97h] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 6(12):40, December 1997. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:1997:CPa

[Ano97i] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 6(4):62–??, April 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:CPb

[Ano97j] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 6(5):84–??, May 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:CPc

[Ano97k] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 6(6):44, June 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:CPe

[Ano97l] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 6(8):52, August 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:CPf

[Ano97m] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 6(9):62, September 1997. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1997:CPg

[Ano97n] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 6(10):72, October 1997. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1997:CPh

[Ano97o] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 6(11): 76, November 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:CPi

[Ano97p] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 6(12): 64, December 1997. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1997:COa

[Ano97q] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 6(4):66–??, April 1997. CODEN SYADE7. ISSN 1061-

Anonymous:1997:COb

[Ano97r] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 6(5):81–??, May 1997. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1997:COc

[Ano97s] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 6(6):80–81, June 1997. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1997:COe

[Ano97t] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 6(8):64, August 1997. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1997:COf

[Ano97u] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 6(9):66–67, September 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:COg

[Ano97v] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 6(10):71, October 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:COh

[Ano97w] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 6(11):71, November 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:COi

[Ano97x] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 6(12):67, December 1997. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Anonymous:1997:CA

[Ano97y] Anonymous. Code availability.

Sys Admin: The Journal for UNIX Systems Administrators, 6(12):4, December 1997. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1997:C

[Ano97z] Anonymous. Correction. Sys Admin: The Journal for UNIX Systems Administrators, 6(12): 71, December 1997. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1997:CDM

[Ano97-27] Anonymous. Correction: Dejan Muhamedagić, PPP Connections Made Easy [Sysadmin, vol. 6, no. 2, pp. 61–63, 65 (1997)]. Sys Admin: The Journal for UNIX Systems Administrators, 6(5):84–??, May 1997. CODEN SYADE7. ISSN 1061-2688. See [Muh97b].

Anonymous:1997:L

[Ano97-28] Anonymous. Letters. Sys Admin: The Journal for UNIX Systems Administrators, 6(10): 80, October 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:NMa

[Ano97-29] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 6(6):88, June 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:NMb

[Ano97-30] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 6(8):72, August 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:NMc

[Ano97-31] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 6(9):72, September 1997. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1997:NPa

[Ano97-32] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 6(4):69–??, April 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:NPb

[Ano97-33] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 6(5):85–??, May 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:NPc

[Ano97-34] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 6(6):82–87, June 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:NPe

[Ano97-35] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 6(8):68–71, August 1997. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1997:NPf

[Ano97-36] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 6(9): 68–71, September 1997. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:1997:NPg

[Ano97-37] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 6(10): 76–79, October 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:NPh

[Ano97-38] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 6(11): 77–80, November 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1997:NPi

[Ano97-39] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 6(12): 68–71, December 1997. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1997:PS

[Ano97-40] Anonymous. Product survey.

Sys Admin: The Journal for UNIX Systems Administrators, 6(5):35–??, May 1997. CODEN SYADE7. ISSN 1061-2688.

Anonymous:1998:AIa

[Ano98a] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 7(1):40, January 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:AIb

[Ano98b] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 7(2):32, February 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:AIc

[Ano98c] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 7(3):32, March 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:AId

[Ano98d] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 7(4):32, April 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:AIe

[Ano98e] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 7(5):32, May 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:AIf

[Ano98f] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 7(6):40, June 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:AIg

[Ano98g] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 7(7):56, July 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:AIh

[Ano98h] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 7(8):40, August 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:AIi

[Ano98i] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 7(9):45, September 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1998:AIj

[Ano98j] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 7(10):64, October 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1998:AIk

[Ano98k] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 7(11):32, November 1998. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:1998:AIl

[Ano98l] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 7(12):64, December 1998. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:1998:CPa

[Ano98m] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 7(1):52, January 1998. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1998:CPb

[Ano98n] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 7(2):55, February 1998. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1998:CPc

[Ano98o] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 7(3):71, March 1998. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1998:CPd

[Ano98p] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 7(4):34, April 1998. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1998:CPe

[Ano98q] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 7(5):66, May 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:CPf

[Ano98r] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 7(6):64, June 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:CPg

[Ano98s] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 7(7):83, July 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:CPh

[Ano98t] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 7(8):67, August 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:CPi

[Ano98u] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 7(9):43, September 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1998:CPj

[Ano98v] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 7(10):76, October 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1998:CPk

[Ano98w] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 7(11): 68, November 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:CPl

[Ano98x] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 7(12): 63, December 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:COa

[Ano98y] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 7(1):71, January 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1998:COb

[Ano98z] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 7(2):65, February 1998. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Anonymous:1998:COc

[Ano98-27] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 7(3):54, March 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1998:COd

[Ano98-28] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 7(4):66–67, April 1998. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Anonymous:1998:COe

[Ano98-29] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 7(5):65, May 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1998:COf

[Ano98-30] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 7(6):45, June 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1998:COg

[Ano98-31] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 7(7):76–77, July 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1998:COh

[Ano98-32] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 7(8):44–45, August 1998. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Anonymous:1998:COi

[Ano98-33] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 7(9):54–55, September 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Anonymous:1998:COj

[Ano98-34] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 7(10):54–55, October 1998. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Anonymous:1998:COk

[Ano98-35] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 7(11):73–75, November 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Anonymous:1998:COl

[Ano98-36] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 7(12):72–73, December 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Anonymous:1998:CAa

[Ano98-37] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 7(1):52, January 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:CAb

[Ano98-38] Anonymous. Code availability.

Sys Admin: The Journal for UNIX Systems Administrators, 7(2):4, February 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:CAc

[Ano98-39] Anonymous. Code availability.

Sys Admin: The Journal for UNIX Systems Administrators, 7(3):71, March 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:CAd

[Ano98-40] Anonymous. Code availability.

Sys Admin: The Journal for UNIX Systems Administrators, 7(4):4, April 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:CAe

[Ano98-41] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 7(5):4, May 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:CAf

[Ano98-42] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 7(6):4, June 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:CAg

[Ano98-43] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 7(7):4, July 1998. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1998:CAh

[Ano98-44] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 7(8):4, August 1998. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1998:CAi

[Ano98-45] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 7(9):4–6, September 1998. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:1998:CAj

[Ano98-46] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 7(10):4, October 1998. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1998:CAk

[Ano98-47] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 7(11):4, November 1998. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:1998:CAl

[Ano98-48] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 7(12):4, December 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1998:C

[Ano98-49] Anonymous. Crossword. Sys Admin: The Journal for UNIX Systems Administrators, 7(6): 67, June 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:CPSa

[Ano98-50] Anonymous. Crossword puzzle solutions. Sys Admin: The Journal for UNIX Systems Administrators, 7(7):75, July 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1998:CPSb

[Ano98-51] Anonymous. Crossword puzzle solutions. Sys Admin: The Journal for UNIX Systems Administrators, 7(8):66, August 1998. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Anonymous:1998:CPSc

[Ano98-52] Anonymous. Crossword puzzle solutions. Sys Admin: The Journal for UNIX Systems Administrators, 7(10):54, October 1998. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Anonymous:1998:CPSd

[Ano98-53] Anonymous. Crossword puzzle solutions. Sys Admin: The Journal for UNIX Systems Administrators, 7(11):73, November 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Anonymous:1998:NMa

[Ano98-54] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 7(7):87, July 1998. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1998:NMb

[Ano98-55] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 7(9):75, September 1998. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:1998:NPa

[Ano98-56] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 7(1):67–70, January 1998. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:1998:NPb

[Ano98-57] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 7(2):67–70, February 1998. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:1998:NPc

[Ano98-58] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 7(3):67–70, March 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1998:NPd

[Ano98-59] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 7(4):68–71, April 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:NPe

[Ano98-60] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 7(5):69–71, May 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:NPf

[Ano98-61] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 7(6):68–71, June 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:NPg

[Ano98-62] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 7(7):84–86, July 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1998:NPh

[Ano98-63] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 7(8):68–71, August 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1998:NPi

[Ano98-64] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 7(9): 71–74, September 1998. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:1998:NPj

[Ano98-65] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 7(10): 77–80, October 1998. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1998:NPk

[Ano98-66] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 7(11): 76–79, November 1998. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1998:NPl

[Ano98-67] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 7(12): 74–76, December 1998. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1998:SACa

[Ano98-68] Anonymous. Sys Admin crossword puzzle. Sys Admin: The Journal for UNIX Systems Administrators, 7(5):64, May 1998. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Anonymous:1998:SACb

[Ano98-69] Anonymous. Sys Admin crossword puzzle. Sys Admin: The Journal for UNIX Systems Administrators, 7(10):54, October 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Anonymous:1999:AIa

[Ano99a] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 8(1):32, January 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1999:AIb

[Ano99b] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 8(2):40, February 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1999:AIc

[Ano99c] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 8(3):40, March 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1999:AId

[Ano99d] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 8(4):40, April 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1999:AIe

[Ano99e] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 8(5):40, May 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1999:AIf

[Ano99f] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 8(6):40, June 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1999:AIg

[Ano99g] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 8(7):40, July 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1999:AIh

[Ano99h] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 8(8):64, August 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1999:AIi

[Ano99i] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 8(9):64, September 1999. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:1999:AIj

[Ano99j] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 8(10):48, October 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1999:AIk

[Ano99k] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 8(11):64, November 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1999:AIl

[Ano99l] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 8(12):72, December 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1999:CPa

[Ano99m] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 8(1):66, January 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1999:CPb

[Ano99n] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 8(2):68, February 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1999:CPc

[Ano99o] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 8(3):58, March 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1999:CPd

[Ano99p] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 8(4):56, April 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1999:CPe

[Ano99q] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 8(5):68, May 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1999:CPf

[Ano99r] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 8(6):44, June 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1999:CPg

[Ano99s] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 8(7):64, July 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1999:CPh

[Ano99t] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 8(8):75, August 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1999:CPi

[Ano99u] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 8(9):72, September 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1999:CPj

[Ano99v] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 8(10):66, October 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1999:CPk

[Ano99w] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 8(11): 76, November 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1999:CPl

[Ano99x] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 8(12): 83, December 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1999:COa

[Ano99y] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 8(1):36, January 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1999:COb

[Ano99z] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 8(2):42, February 1999. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Anonymous:1999:COc

[Ano99-27] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 8(3):67, March 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1999:COd

[Ano99-28] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 8(4):36, April 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1999:COe

[Ano99-29] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 8(5):67, May 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1999:COf

[Ano99-30] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 8(6):60, June 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1999:COg

[Ano99-31] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 8(7):34, July 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1999:COh

[Ano99-32] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 8(8):66, August 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1999:COi

[Ano99-33] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 8(9):82, September 1999. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Anonymous:1999:COj

[Ano99-34] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 8(10):65, October 1999. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Anonymous:1999:COk

[Ano99-35] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 8(11):75, November 1999. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Anonymous:1999:COl

[Ano99-36] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 8(12):84, December 1999.

CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Anonymous:1999:CAa

[Ano99-37] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 8(1):4, January 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1999:CAb

[Ano99-38] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 8(2):4, February 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1999:CAc

[Ano99-39] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 8(3):4, March 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1999:CAd

[Ano99-40] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 8(4):4, April 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1999:CAe

[Ano99-41] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 8(5):4, May 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1999:CAf

[Ano99-42] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 8(6):4, June 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1999:CAg

[Ano99-43] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 8(7):4, July 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1999:CAh

[Ano99-44] Anonymous. Code availability.

Sys Admin: The Journal for UNIX Systems Administrators, 8(8):4, August 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1999:CAi

[Ano99-45] Anonymous. Code availability.

Sys Admin: The Journal for UNIX Systems Administrators, 8(9):4, September 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1999:CAj

[Ano99-46] Anonymous. Code availability.

Sys Admin: The Journal for UNIX Systems Administrators, 8(10):4, October 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1999:CAk

[Ano99-47] Anonymous. Code availability.

Sys Admin: The Journal for UNIX Systems Administrators, 8(11):4, November 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1999:CAl

[Ano99-48] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 8(12):4, December 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1999:NMa

[Ano99-49] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 8(1):71, January 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1999:NMb

[Ano99-50] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 8(5):72, May 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1999:NMc

[Ano99-51] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 8(10): 79, October 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1999:NMe

[Ano99-52] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 8(12): 82, December 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1999:NPa

[Ano99-53] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 8(1):67–70, January 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1999:NPb

[Ano99-54] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 8(2):69–71, February 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1999:NPc

[Ano99-55] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 8(3):69–71, March 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:1999:NPd

[Ano99-56] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 8(4):68–71, April 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1999:NPe

[Ano99-57] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 8(5):69–71, May 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1999:NPf

[Ano99-58] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 8(6):68–71, June 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1999:NPg

[Ano99-59] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 8(7):68–71, July 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1999:NPh

[Ano99-60] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 8(8):76–79, August 1999. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:1999:NPi

[Ano99-61] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 8(9): 83–88, September 1999. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:1999:NPj

[Ano99-62] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 8(10): 75–78, October 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:1999:NPk

[Ano99-63] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 8(11): 77–79, November 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:1999:NPl

[Ano99-64] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 8(12): 85–87, December 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:AIa

[Ano00a] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 9(1):64, January 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:AIb

[Ano00b] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 9(2):48, February 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:AIc

[Ano00c] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 9(3):64, March 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:AId

[Ano00d] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 9(4):64, April 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2000:AIe

[Ano00e] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 9(5):64, May 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2000:AIf

[Ano00f] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 9(6):80, June 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2000:AIg

[Ano00g] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 9(7):??, July 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2000:AIh

[Ano00h] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 9(8):72, August 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2000:AIi

[Ano00i] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 9(9):64, September 2000. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:2000:AIj

[Ano00j] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 9(10):64, October 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:2000:AIk

[Ano00k] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 9(11):64, November 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:2000:AIl

[Ano00l] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 9(11S):40, November 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:2000:AIm

[Ano00m] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 9(12):64, December 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:2000:CPa

[Ano00n] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 9(1):71, January 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:CPb

[Ano00o] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 9(2):47, February 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:CPc

[Ano00p] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 9(3):66, March 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:CPd

[Ano00q] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 9(4):72, April 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:CPe

[Ano00r] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 9(5):92, May 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:CPf

[Ano00s] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 9(6):88, June 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:CPg

[Ano00t] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 9(7):??, July 2000. CODEN SYADE7.

ISSN 1061-2688. URL http://

www.samag.com/.

Anonymous:2000:CPh

[Ano00u] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 9(8):78, August 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:CPi

[Ano00v] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 9(9):90, September 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:2000:CPj

[Ano00w] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 9(10):90, October 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:2000:CPk

[Ano00x] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 9(11): 90, November 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:CPl

[Ano00y] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 9(11S): 68, November 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:CPm

[Ano00z] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 9(12): 92, December 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2000:CO

[Ano00-27] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 9(7):??, July 2000. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:2000:CAa

[Ano00-28] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 9(1):4, January 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2000:CAb

[Ano00-29] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 9(2):4, February 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2000:CAc

[Ano00-30] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 9(3):4, March 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2000:CAd

[Ano00-31] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 9(4):4, April 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:CAe

[Ano00-32] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 9(5):4, May 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:CAf

[Ano00-33] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 9(6):4, June 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:CAg

[Ano00-34] Anonymous. Code availability.

Sys Admin: The Journal for UNIX Systems Administrators, 9(7):??, July 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:CAh

[Ano00-35] Anonymous. Code availability.

Sys Admin: The Journal for UNIX Systems Administrators, 9(8):4, August 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:CAi

[Ano00-36] Anonymous. Code availability.

Sys Admin: The Journal for UNIX Systems Administrators, 9(9):4, September 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:2000:CAj

[Ano00-37] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 9(10):4, October 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:CAk

[Ano00-38] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 9(11):4, November 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:2000:CAl

[Ano00-39] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 9(12):4, December 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Table:2000:CRN

[Ano00-40] Anonymous. Converging remote networks into your routing table. Sys Admin: The Journal for UNIX Systems Administrators, 9(7):??, July 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:2000:NMa

[Ano00-41] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 9(1):79, January 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:NMb

[Ano00-42] Anonymous. New messages.

Sys Admin: The Journal for UNIX Systems Administrators, 9(6):95, June 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:NMc

[Ano00-43] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 9(7):??, July 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Anonymous:2000:NMd

[Ano00-44] Anonymous. New messages.

Sys Admin: The Journal for UNIX Systems Administrators, 9(8):87, August 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:NMe

[Ano00-45] Anonymous. New messages.

Sys Admin: The Journal for UNIX Systems Administrators, 9(9):95, September 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:2000:NMf

[Ano00-46] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 9(10): 94–95, October 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:NMg

[Ano00-47] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 9(11): 94–95, November 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:NPa

[Ano00-48] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 9(1):76–78, January 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:2000:NPb

[Ano00-49] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 9(2):75–78, February 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:2000:NPc

[Ano00-50] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 9(3):76–79, March 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:2000:NPd

[Ano00-51] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 9(4):76–79, April 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:NPe

[Ano00-52] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 9(5):93–95, May 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:NPf

[Ano00-53] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 9(6):92–94, June 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:NPg

[Ano00-54] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 9(7):??, July 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Anonymous:2000:NPh

[Ano00-55] Anonymous. New products.

Sys Admin: The Journal for UNIX Systems Administrators, 9(8):84–86, August 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:2000:NPi

[Ano00-56] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 9(9): 91–94, September 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:2000:NPj

[Ano00-57] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 9(10): 91–93, October 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2000:NPk

[Ano00-58] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 9(11): 91–93, November 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2000:NPl

[Ano00-59] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 9(11S): 69–71, November 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2000:NPm

[Ano00-60] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 9(12): 93, December 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2001:AIa

[Ano01a] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 10(1):96, January 2001. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:2001:AIb

[Ano01b] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 10(2):64, February 2001. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:2001:AIc

[Ano01c] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 10(3):64, March 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:AId

[Ano01d] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 10(4):72, April 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:AIe

[Ano01e] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 10(5):72, May 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:AIf

[Ano01f] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 10(6):72, June 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:AIg

[Ano01g] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 10(6s):32, June 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:AIh

[Ano01h] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 10(7):40, July 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2001:AIi

[Ano01i] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 10(8):40, August 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2001:AIj

[Ano01j] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 10(9):40, September 2001. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:2001:AIk

[Ano01k] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 10(10):48, October 2001. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:2001:CPa

[Ano01l] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 10(1): 106, January 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2001:CPb

[Ano01m] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 10(2): 76, February 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2001:CPc

[Ano01n] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 10(3):75, March 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:CPd

[Ano01o] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 10(4):79, April 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:CPe

[Ano01p] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 10(5):70, May 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:CPf

[Ano01q] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 10(8):56, August 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:CPg

[Ano01r] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 10(9): 66, September 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:CPh

[Ano01s] Anonymous. Call for papers.

Sys Admin: The Journal for UNIX Systems Administrators, 10(10):64, October 2001. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Anonymous:2001:CO

[Ano01t] Anonymous. Career opportunities. Sys Admin: The Journal for UNIX Systems Administrators, 10(9):??, September 2001. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Anonymous:2001:CAa

[Ano01u] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 10(1):4, January 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:CAb

[Ano01v] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 10(2):4, February 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:CAc

[Ano01w] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 10(3):4, March 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:CAd

[Ano01x] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 10(4):4, April 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:CAe

[Ano01y] Anonymous. Code availability.

Sys Admin: The Journal for UNIX Systems Administrators, 10(5):4, May 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:CAf

[Ano01z] Anonymous. Code availability.

Sys Admin: The Journal for UNIX Systems Administrators, 10(6):4, June 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:CAg

[Ano01-27] Anonymous. Code availability.

Sys Admin: The Journal for UNIX Systems Administrators, 10(6s):4, June 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:CAh

[Ano01-28] Anonymous. Code availability.

Sys Admin: The Journal for UNIX Systems Administrators, 10(7):4, July 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:CAi

[Ano01-29] Anonymous. Code availability.

Sys Admin: The Journal for UNIX Systems Administrators, 10(8):4, August 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:CAj

[Ano01-30] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 10(9):4, September 2001. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:2001:CAk

[Ano01-31] Anonymous. Code availability. Sys Admin: The Journal for UNIX Systems Administrators, 10(10):4, October 2001. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:2001:LID

[Ano01-32] Anonymous. Linux intrusion detection poster. Sys Admin: The Journal for UNIX Systems Administrators, 10(1):??, January 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Anonymous:2001:Ma

[Ano01-33] Anonymous. Marketplace. Sys Admin: The Journal for UNIX Systems Administrators, 10(4): 83, April 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2001:Mb

[Ano01-34] Anonymous. Marketplace. Sys Admin: The Journal for UNIX Systems Administrators, 10(6): 81, June 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2001:Mc

[Ano01-35] Anonymous. Marketplace. Sys Admin: The Journal for UNIX Systems Administrators, 10(9): 67, September 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:NMa

[Ano01-36] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 10(1): 110–111, January 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:NMb

[Ano01-37] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 10(2): 77, February 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:NMc

[Ano01-38] Anonymous. New messages.

Sys Admin: The Journal for UNIX Systems Administrators, 10(4):88, April 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:NMd

[Ano01-39] Anonymous. New messages.

Sys Admin: The Journal for UNIX Systems Administrators, 10(5):87, May 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:NMe

[Ano01-40] Anonymous. New messages.

Sys Admin: The Journal for UNIX Systems Administrators, 10(6):80, June 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:NMf

[Ano01-41] Anonymous. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 10(9): 72, September 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:NPa

[Ano01-42] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 10(1): 107–109, January 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:NPb

[Ano01-43] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 10(2): 73–75, February 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:NPc

[Ano01-44] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 10(3): 76–78, March 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:NPd

[Ano01-45] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 10(4): 80–82, April 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:NPe

[Ano01-46] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 10(5): 80–82, May 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:NPf

[Ano01-47] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 10(6): 77–79, June 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:NPg

[Ano01-48] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 10(6s): 53–54, June 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:NPh

[Ano01-49] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 10(7): 64–66, July 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:NPi

[Ano01-50] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 10(8): 65–67, August 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:NPj

[Ano01-51] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 10(9): 63–65, September 2001. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Anonymous:2001:NPk

[Ano01-52] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 10(10): 58–59, October 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2001:Ra

[Ano01-53] Anonymous. Recruitment. Sys Admin: The Journal for UNIX Systems Administrators, 10(2): 78–79, February 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2001:Rb

[Ano01-54] Anonymous. Recruitment. Sys Admin: The Journal for UNIX Systems Administrators, 10(3): 79–80, March 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2001:Rc

[Ano01-55] Anonymous. Recruitment. Sys Admin: The Journal for UNIX Systems Administrators, 10(4): 86–87, April 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Anonymous:2001:Rd

[Ano01-56] Anonymous. Recruitment. Sys Admin: The Journal for UNIX Systems Administrators, 10(6): 82, June 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:Re

[Ano01-57] Anonymous. Recruitment. Sys Admin: The Journal for UNIX Systems Administrators, 10(6s): 55, June 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Anonymous:2001:TEa

[Ano01-58] Anonymous. Training and education. Sys Admin: The Journal for UNIX Systems Administrators, 10(4):84–85, April 2001. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Anonymous:2001:TEb

[Ano01-59] Anonymous. Training and education. Sys Admin: The Journal for UNIX Systems Administrators, 10(6):83–87, June 2001. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Anonymous:2001:TEc

[Ano01-60] Anonymous. Training and education. Sys Admin: The Journal for UNIX Systems Administrators, 10(9):68–71, September 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Anonymous:2002:AIa

[Ano02a] Anonymous. Advertiser index.

Sys Admin: The Journal for UNIX Systems Administrators, 11(3):40, March 2002. CODEN SYADE7. ISSN 1061-2688.

Anonymous:2002:AIb

[Ano02b] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 11(4):48, April 2002. CODEN SYADE7. ISSN 1061-2688.

Anonymous:2002:AIc

[Ano02c] Anonymous. Advertiser index. Sys Admin: The Journal for UNIX Systems Administrators, 11(5):56, May 2002. CODEN SYADE7. ISSN 1061-2688.

Anonymous:2002:CPa

[Ano02d] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 11(3):51, March 2002. CODEN SYADE7. ISSN 1061-2688.

Anonymous:2002:CPb

[Ano02e] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 11(4):57, April 2002. CODEN SYADE7. ISSN 1061-2688.

Anonymous:2002:CPc

[Ano02f] Anonymous. Call for papers. Sys Admin: The Journal for UNIX Systems Administrators, 11(5):65, May 2002. CODEN SYADE7. ISSN 1061-2688.

Anonymous:2002:Ma

[Ano02g] Anonymous. Marketplace. Sys Admin: The Journal for UNIX Systems Administrators, 11(3): 68, March 2002. CODEN SYADE7. ISSN 1061-2688.

Anonymous:2002:Mb

[Ano02h]  Anonymous.  Marketplace.

Sys Admin: The Journal for UNIX Systems Administrators, 11(5):66, May 2002. CODEN SYADE7. ISSN 1061-2688.

Anonymous:2002:NM

[Ano02i] Anonymous. New messages.

Sys Admin: The Journal for UNIX Systems Administrators, 11(4):64, April 2002. CODEN SYADE7. ISSN 1061-2688.

Anonymous:2002:NPa

[Ano02j] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 11(3): 71–72, March 2002. CODEN SYADE7. ISSN 1061-2688.

Anonymous:2002:NPb

[Ano02k] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 11(4): 58–60, April 2002. CODEN SYADE7. ISSN 1061-2688.

Anonymous:2002:NPc

[Ano02l] Anonymous. New products. Sys Admin: The Journal for UNIX Systems Administrators, 11(5): 71–72, May 2002. CODEN SYADE7. ISSN 1061-2688.

Anonymous:2002:TEa

[Ano02m] Anonymous. Training and education. Sys Admin: The Journal for UNIX Systems Administrators, 11(3):69, March 2002. CO-DEN SYADE7. ISSN 1061-2688.

Anonymous:2002:TEb

[Ano02n] Anonymous. Training and education. Sys Admin: The Journal for UNIX Systems Administrators, 11(5):67–70, May 2002. CODEN SYADE7. ISSN 1061-2688.

Appleton:2000:PPS

[App00] Randy Appleton. Perl and the practical systems administrator. Sys Admin: The Journal for UNIX Systems Administrators, 9(8):65–68, August 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Apthorpe:2002:IRT

[Apt02] Bob Apthorpe. Increasing reliability through forensic operations. Sys Admin: The Journal for UNIX Systems Administrators, 11(2):24, 26–29, February 2002. CODEN SYADE7. ISSN 1061-2688.

Alder:1993:FBR

[AS93] Jon Alder and Ed Schaefer. Fast backup and restore scripts for DAT drives. Sys Admin: The Journal for UNIX Systems Administrators, 2(5):19– ??, September/October 1993. CODEN SYADE7. ISSN 1061-2688.

Schaefer:1996:DDE

[AS96] Jon Alder and Ed Schaefer. A dispatcher for database engine alarms. Sys Admin: The Journal for UNIX Systems Administrators, 5(8):??, August 1996.

CODEN SYADE7. ISSN 1061-2688.

Ashworth:1992:RR

[Ash92] Judith Ashworth. rdist to the rescue! Sys Admin: The Journal for UNIX Systems Administrators, 1(4):18–??, November/December 1992. CODEN SYADE7. ISSN 1061-2688.

Ashworth:1993:DR

[Ash93] Judith Ashworth. Dump and restore. Sys Admin: The Journal for UNIX Systems Administrators, 2(5):49–??, September/October 1993. CODEN SYADE7. ISSN 1061-2688.

Bahadur:2001:FSW

[Bah01] Gary Bahadur. Freeware security Web tools. Sys Admin: The Journal for UNIX Systems Administrators, 10(3):32, 34, 36– 37, 39–40, March 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Baker:1995:IAN

[Bak95] Mark Baker. Integrating automounter into NIS across two subnets. Sys Admin: The Journal for UNIX Systems Administrators, 4(5):40–??, September/October 1995. CODEN SYADE7. ISSN 1061-2688.

Ball:1999:WSB

[Bal99] Jimmy Ball. Web security basics with Apache: Authentication and secure log files. Sys Admin: The Journal for UNIX Systems Administrators, 8(2): 43–46, February 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Barker:1996:EFa

[Bar96a] Ralph Barker. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 5(10):6, October 1996. CODEN SYADE7. ISSN 1061-2688.

Barker:1996:EFb

[Bar96b] Ralph Barker. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 5(11):6, November 1996. CO-DEN SYADE7. ISSN 1061-2688.

Barker:1996:EFc

[Bar96c] Ralph Barker. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 5(11):6, November 1996. CO-DEN SYADE7. ISSN 1061-2688.

Barker:1996:EFd

[Bar96d] Ralph Barker. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 5(12):6, December 1996. CO-DEN SYADE7. ISSN 1061-2688.

Barker:1997:EFa

[Bar97a] Ralph Barker. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 6(1):6, January 1997. CODEN SYADE7. ISSN 1061-2688.

Barker:1997:EFb

[Bar97b] Ralph Barker. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 6(2):6, February 1997. CODEN SYADE7. ISSN 1061-2688.

Barker:1997:EFc

[Bar97c] Ralph Barker. Editor’s forum.

Sys Admin: The Journal for UNIX Systems Administrators, 6(3):6, March 1997. CODEN SYADE7. ISSN 1061-2688.

Barker:1997:EFd

[Bar97d] Ralph Barker. Editor’s forum.

Sys Admin: The Journal for UNIX Systems Administrators, 6(4):6, April 1997. CODEN SYADE7. ISSN 1061-2688.

Barker:1997:EFe

[Bar97e] Ralph Barker. Editor’s forum.

Sys Admin: The Journal for UNIX Systems Administrators, 6(5):6, May 1997. CODEN SYADE7. ISSN 1061-2688.

Barker:1997:EFf

[Bar97f] Ralph Barker. Editor’s forum.

Sys Admin: The Journal for UNIX Systems Administrators, 6(6):6, June 1997. CODEN SYADE7. ISSN 1061-2688.

Barker:1997:EFh

[Bar97g] Ralph Barker. Editor’s forum.

Sys Admin: The Journal for UNIX Systems Administrators, 6(8):6, August 1997. CODEN SYADE7. ISSN 1061-2688.

Barker:1997:EFi

[Bar97h] Ralph Barker. Editor’s forum.

Sys Admin: The Journal for UNIX Systems Administrators, 6(9):6, September 1997. CO-DEN SYADE7. ISSN 1061-2688.

Barker:1997:EFj

[Bar97i] Ralph Barker. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 6(10):6, October 1997. CODEN SYADE7. ISSN 1061-2688.

Barker:1997:EFk

[Bar97j] Ralph Barker. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 6(11):6, November 1997. CO-DEN SYADE7. ISSN 1061-2688.

Barker:1997:EFl

[Bar97k] Ralph Barker. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 6(12):6, December 1997. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Barker:1997:PSH

[Bar97l] Ralph Barker. Product survey: High-speed networking. Sys Admin: The Journal for UNIX Systems Administrators, 6(10): 21, 22, 24, 25, 27, 29–32, 34– 36, 38, October 1997. CODEN SYADE7. ISSN 1061-2688.

Barker:1998:EFa

[Bar98a] Ralph Barker. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 7(1):6, January 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Barker:1998:EFb

[Bar98b] Ralph Barker. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 7(2):6, February 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Barker:1998:EFc

[Bar98c] Ralph Barker. Editor’s forum.

Sys Admin: The Journal for UNIX Systems Administrators, 7(3):6, March 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Barker:1998:EFd

[Bar98d] Ralph Barker. Editor’s forum.

Sys Admin: The Journal for UNIX Systems Administrators, 7(4):6, April 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Barker:1998:EFe

[Bar98e] Ralph Barker. Editor’s forum.

Sys Admin: The Journal for UNIX Systems Administrators, 7(5):6, May 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Barker:1998:EFf

[Bar98f] Ralph Barker. Editor’s forum.

Sys Admin: The Journal for UNIX Systems Administrators, 7(6):6, June 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Barker:1998:EFg

[Bar98g] Ralph Barker. Editor’s forum.

Sys Admin: The Journal for UNIX Systems Administrators, 7(7):6, July 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Barker:1998:EFh

[Bar98h] Ralph Barker. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 7(8):6, August 1998. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Barker:1998:EFi

[Bar98i] Ralph Barker. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 7(9):6, September 1998. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Barker:1998:EFj

[Bar98j] Ralph Barker. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 7(10):6, October 1998. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Barker:1998:EFk

[Bar98k] Ralph Barker. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 7(11):6, November 1998. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Barker:1998:EFl

[Bar98l] Ralph Barker. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 7(12):6, December 1998. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Barker:1998:SAD

[Bar98m] Ralph Barker. Sys Admin’s desktop tape library buyer’s guide. Sys Admin: The Journal for UNIX Systems Administrators, 7(9):29, 31–41, September 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Barker:1999:AL

[Bar99a] Ralph Barker. Alternative libraries. Sys Admin: The Journal for UNIX Systems Administrators, 8(11):35–38, November 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Barker:1999:EFa

[Bar99b] Ralph Barker. Editor’s forum.

Sys Admin: The Journal for UNIX Systems Administrators, 8(1):6, January 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Barker:1999:EFb

[Bar99c] Ralph Barker. Editor’s forum.

Sys Admin: The Journal for UNIX Systems Administrators, 8(2):6, February 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Barker:1999:EFc

[Bar99d] Ralph Barker. Editor’s forum.

Sys Admin: The Journal for UNIX Systems Administrators, 8(3):6, March 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Barker:1999:EFd

[Bar99e] Ralph Barker. Editor’s forum.

Sys Admin: The Journal for UNIX Systems Administrators, 8(4):6, April 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Barker:1999:EFe

[Bar99f] Ralph Barker. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 8(5):6, May 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Barker:1999:EFf

[Bar99g] Ralph Barker. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 8(6):6, June 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Barker:1999:EFSa

[Bar99h] Ralph Barker. Editor’s forum: syslog. Sys Admin: The Journal for UNIX Systems Administrators, 8(7):6, July 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Barker:1999:EFSb

[Bar99i] Ralph Barker. Editor’s forum: syslog. Sys Admin: The Journal for UNIX Systems Administrators, 8(8):6, August 1999. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Barker:1999:EFU

[Bar99j] Ralph Barker. Editor’s forum: Uncommon common ground. Sys Admin: The Journal for UNIX Systems Administrators, 8(9):6, September 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Baum:1996:GUI

[Bau96] Steven Baum. Graphical user interface development systems. Sys Admin: The Journal for UNIX Systems Administrators, 5(4):??, April 1996. CODEN SYADE7. ISSN 1061-2688.

Bayern:2000:MWW

[Bay00] Shawn Bayern. Making a wish:

The Web-interface shell. Sys Admin: The Journal for UNIX Systems Administrators, 9(7):??, July 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Bayern:2001:MB

[Bay01] Shawn Bayern. Minimal backups. Sys Admin: The Journal for UNIX Systems Administrators, 10(4):28, 30, April 2001. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Bowden:1999:LFM

[BB99] Terrehon Bowden and Bodo Bauer. Linux firewall and masquerading: The IP chains concept in Linux 2.2. Sys Admin: The Journal for UNIX Systems Administrators, 8(6):23–24, 26– 30, 32, June 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Brian:2001:HDP

[BB01] Dan Brian and Jon Bjornstad.

Helping the disabled with Perl/ Tk. Sys Admin: The Journal for UNIX Systems Administrators, 10(12):P61–P65, December 2001. CODEN SYADE7. ISSN 1061-2688. Perl Journal supplement.

Bayern:1999:HCB

[BC99] Shawn Bayern and Brian Carp. How to crawl back inside your shell. Sys Admin: The Journal for UNIX Systems Administrators, 8(11):55–58, November 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Bear:1996:HSI

[Bea96] Cory Bear. How safe is your online backup. Sys Admin: The Journal for UNIX Systems Administrators, 5(7):26–??, July 1996. CODEN SYADE7. ISSN 1061-2688.

Beck:1999:APT

[Bec99] Joe Beck. ABCs of performance tuning. Sys Admin: The Journal for UNIX Systems Administrators, 8(7):8, 10, 12, 14, 16, 18, July 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Beck:2000:SAT

[Bec00] Joe Beck. A Sys Admin’s toolkit. Sys Admin: The Journal for UNIX Systems Administrators, 9(5):79–84, May 2000. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Beecher:2001:GG

[Bee01] David Beecher. Getting out/ getting in. Sys Admin: The Journal for UNIX Systems Administrators, 10(1):45–46, 48, 50, January 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Bellenot:1998:CUA

[Bel98] Steven Bellenot. Creating user accounts. Sys Admin: The Journal for UNIX Systems Administrators, 7(11):57–60, November 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Bennett:1998:SNA

[Ben98]  Larry Bennett.  Streamlin-

ing network administration with DHCP. Sys Admin: The Journal for UNIX Systems Administrators, 7(3):14, 16, 18, 20, 22–23, March 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Berry:1994:MON

[Ber94]  Robert Berry.  Monitoring

and optimizing NFS performance. Sys Admin: The Journal for UNIX Systems Administrators, 3(6):66–??, November/December 1994. CODEN SYADE7. ISSN 1061-2688.

Berry:1995:PPQ

[Ber95] Robert Berry. prq: A print queue monitor and manager. Sys Admin: The Journal for UNIX Systems Administrators, 4(1):??, January/February 1995. CO-DEN SYADE7. ISSN 1061-2688.

Berry:1998:ACU

[Ber98a] Joe Berry. Assessing CPU utilization. Sys Admin: The Journal for UNIX Systems Administrators, 7(5):57–60, May 1998. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Berry:1998:MAN

[Ber98b] Joseph Berry. A methodology for analyzing Network File System performance problems. Sys Admin: The Journal for UNIX Systems Administrators, 7(7):45–53, July 1998. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Berninger:1999:NM

[Ber99] John Berninger. Network monitoring. Sys Admin: The Journal for UNIX Systems Administrators, 8(2):55–58, February 1999. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Barker:1999:DD

[BF99] Chad Barker and Mark Frisbey. Designing for disaster. Sys Admin: The Journal for UNIX Systems Administrators, 8(4):47–50, April 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Bice:2002:NDM

[Bic02]  Brent Bice.  Network dia-

gramming and monitoring with Scotty. Sys Admin: The Journal for UNIX Systems Administrators, 11(1):29–32, January 2002. CODEN SYADE7. ISSN 1061-2688.

Berninger:1999:SDI

[BL99] John Berninger and Dana Lee.

SAs and DBAs: Interactions across traditional lines of responsibility. Sys Admin: The Journal for UNIX Systems Administrators, 8(7):35–39, July 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Beilman:2000:AR

[BL00] Bryon Beilman and Jamy Libsack. The art of the refresh. Sys Admin: The Journal for UNIX Systems Administrators, 9(8):69–71, August 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Buckeye:2002:RDF

[BL02] Brian Buckeye and Kevin Liston. Recovering deleted files in Linux. Sys Admin: The Journal for UNIX Systems Administrators, 11(4):42, 44–47, April 2002. CODEN SYADE7. ISSN 1061-2688.

Blader:1999:FTV

[Bla99] Robert Blader. File transfer and verification between nonconnected networks. Sys Admin: The Journal for UNIX Systems Administrators, 8(10): 61–64, October 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Bahadur:2001:UFV

[BmC01] Gary Bahadur and Yen ming Chen. Using freeware vulnerability scanners. Sys Admin: The Journal for UNIX Systems Administrators, 10(4):69– 71, 73–75, April 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Blackmon:2001:HAF

[BN01] Steve Blackmon and John Nguyen. High-availability file server with heartbeat. Sys Admin: The Journal for UNIX Systems Administrators, 10(9): 24, 26, 28, 30, 31–32, September 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Boeheim:1997:P

[Boe97] Chuck Boeheim. Patrol. Sys Admin: The Journal for UNIX Systems Administrators, 6(2): 16–??, February 1997. CODEN SYADE7. ISSN 1061-2688.

Bollenbach:2000:AFR

[Bol00] Bill Bollenbach. AIX filesystem reorg made simple. Sys Admin: The Journal for UNIX Systems Administrators, 9(11): 85–86, November 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Botsford:1993:QGE

[Bot93] Sherwood Botsford. quota: A gentle enforcer. Sys Admin: The Journal for UNIX Systems Administrators, 2(3):43–??, May/ June 1993. CODEN SYADE7. ISSN 1061-2688.

Bourke:1999:CND

[Bou99] Tony Bourke. Console network design. Sys Admin: The Journal for UNIX Systems Administrators, 8(12):77–78, December 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Bahadur:1999:SDR

[BR99] Gary Bahadur and Dan Robertson. Security diagnostic review for Solaris 2.X. Sys Admin: The Journal for UNIX Systems Administrators, 8(1):8, 10, 12, 14, January 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Brillhart:1994:DNT

[Bri94] Dave Brillhart. The downtime notification and tracking system. Sys Admin: The Journal for UNIX Systems Administrators, 3(3):6–??, May/June 1994. CODEN SYADE7. ISSN 1061-2688.

Brignoni:2000:LDB

[Bri00] Lizmari Brignoni. LDAP directory basics. Sys Admin: The Journal for UNIX Systems Administrators, 9(6):23–24, 26–27, 29, 31, June 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Brown:1992:CHL

[Bro92a] G. Clark Brown. The cure for hung up, locked up, and frozen terminals. Sys Admin: The Journal for UNIX Systems Administrators, 1(3):47– ??, September/October 1992. CODEN SYADE7. ISSN 1061-2688.

Brown:1992:TCT

[Bro92b] G. Clark Brown. Testing capabilities in terminals and emulators. Sys Admin: The Journal for UNIX Systems Administrators, 1(1):11–??, May/June 1992. CODEN SYADE7. ISSN 1061-2688.

Brown:1993:TCR

[Bro93] G. Clark Brown. Terminal choices: The right tool for the job. Sys Admin: The Journal for UNIX Systems Administrators, 2(1):34–??, January/February 1993. CODEN SYADE7. ISSN 1061-2688.

Brown:1995:CFB

[Bro95] Donald Brown. Creating filters for BSD print spoolers. Sys Admin: The Journal for UNIX Systems Administrators, 4(4): 111–??, July/August 1995. CO-DEN SYADE7. ISSN 1061-2688.

Brownrigg:1996:CXW

[Bro96] Richard L. Brownrigg. Customizing X Window sessions with fvwm. Sys Admin: The Journal for UNIX Systems Administrators, 5(4):??, April 1996. CODEN SYADE7. ISSN 1061-2688.

Brown:1997:WBN

[Bro97] Steven Brown. Web-based network management — A free addon to HP Openview. Sys Admin: The Journal for UNIX Systems Administrators, 6(9):8, 11, 12, 14, 15, September 1997. CO-DEN SYADE7. ISSN 1061-2688.

Brockmeier:2000:SGL

[Bro00] Joe “Zonker” Brockmeier. Selecting a GUI for Linux. Sys Admin: The Journal for UNIX Systems Administrators, 9(1): 59–61, January 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Brunet:1994:FSC

[Bru94] Fred Brunet. fcd: A smart change directory. Sys Admin: The Journal for UNIX Systems Administrators, 3(6):33– ??, November/December 1994. CODEN SYADE7. ISSN 1061-2688.

Bryson:1997:FM

[Bry97a] Donald Bryson. Forget me not. Sys Admin: The Journal for UNIX Systems Administrators, 6(11):43–44, 46–48, 50– 51, November 1997. CODEN SYADE7. ISSN 1061-2688.

Bryson:1997:PSS

[Bry97b] Donald Bryson. Problem solving with sar. Sys Admin: The Journal for UNIX Systems Administrators, 6(12):52, 54–55, 58– 59, December 1997. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Brym:2000:MUT

[Bry00] Mats Brymér. Managing usability with the TERM variable.

Sys Admin: The Journal for UNIX Systems Administrators, 9(4):59–60, 62–63, 65–68, April 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Bundy:1996:HGU

[Bun96] Charles Bundy. HTML as a graphical user interface for sys admin tasks. Sys Admin: The Journal for UNIX Systems Administrators, 5(2):51–56, 59, 60, 62, 63, February 1996. CODEN SYADE7. ISSN 1061-2688.

Bundy:1997:CLI

[Bun97] Charles C. Bundy, IV. Connect LAN users to the Internet using Linux and IP masquerade. Sys Admin: The Journal for UNIX Systems Administrators, 6(1):22–??, January 1997. CODEN SYADE7. ISSN 1061-2688.

Burke:2001:UC

[Bur01a] Sean M. Burke. Unidecode. Sys Admin: The Journal for UNIX Systems Administrators, 10(12): P54, P56–P60, December 2001. CODEN SYADE7. ISSN 1061-2688. Perl Journal supplement.

Burns:2001:LFD

[Bur01b] Victor Burns. Locking the front door of password security. Sys Admin: The Journal for UNIX Systems Administrators, 10(8): 47–51, August 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Bush:1994:IOSa

[Bus94a] Chris Bush. Internet online services: archie. Sys Admin: The Journal for UNIX Systems Administrators, 3(1):37–??, January/February 1994. CODEN SYADE7. ISSN 1061-2688.

Bush:1994:IOSb

[Bus94b] Chris Bush. Internet online services: gopher. Sys Admin: The Journal for UNIX Systems Administrators, 3(2): 79–??, March/April 1994. CO-DEN SYADE7. ISSN 1061-2688.

Bush:1999:WET

[Bus99a] Chris Bush. Web-exclusive:

tcp wrapper macro basics. Sys Admin: The Journal for UNIX Systems Administrators, 8(5):??, May 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/. Web version

only: not in the journal issue on paper.

Bush:1999:ENS

[Bus99b] Christopher Bush. Enhancing network security with tcp wrapper. Sys Admin: The Journal for UNIX Systems Administrators, 8(5):22, 24, 26–28, May 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Bush:1999:PW

[Bus99c] Christopher Bush. Perls of wisdom. Sys Admin: The Journal for UNIX Systems Administrators, 8(3):53–57, March 1999.

CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Bush:1999:UAI

[Bus99d]  Christopher Bush.  Using

Apache as an Internet proxy server. Sys Admin: The Journal for UNIX Systems Administrators, 8(2):16, 18, 20, 22, 24–29, February 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Bush:2001:LAS

[Bus01] Chris Bush. Linux as an application server – the Tomcat way. Sys Admin: The Journal for UNIX Systems Administrators, 10(1):31–32, 34, 36, 39–40, 42, 44, January 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Byers:1999:BUF

[Bye99] Chris Byers. Balancing user flexibility and administrative support. Sys Admin: The Journal for UNIX Systems Administrators, 8(10):29–30, 32, 34, 36, October 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Callaway:2001:FAF

[Cal01] George Callaway. A framework for automated file transfer. Sys Admin: The Journal for UNIX Systems Administrators, 10(9): ??, September 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/. Web

only.

Canavan:1999:PA

[Can99] John E. Canavan. Painless AIX.

Sys Admin: The Journal for UNIX Systems Administrators, 8(4):51–55, April 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Caruso:1997:ICR

[Car97] David T. Caruso. Implementing a CD-ROM jukebox solution. Sys Admin: The Journal for UNIX Systems Administrators, 6(4):16–??, April 1997. CODEN SYADE7. ISSN 1061-2688.

Carlson:2001:PMS

[Car01a] Br. David Carlson. Poor man’s search engine. Sys Admin: The Journal for UNIX Systems Administrators, 10(6):24, 26, 28, 30, June 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Carver:2001:MAS

[Car01b] David Carver. Managing AIX with SMIT. Sys Admin: The Journal for UNIX Systems Administrators, 10(6s):43–45, June 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Cashin:2001:FSC

[Cas01a] Ed L. Cashin. A flexible system for centralized backup. Sys Admin: The Journal for UNIX Systems Administrators, 10(9): 46, 48–50, September 2001. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Cashin:2001:IFV

[Cas01b] Ed L. Cashin. Integrit for file verification. Sys Admin: The Journal for UNIX Systems Administrators, 10(8):??, August 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/. Web only.

Caywood:1994:ITA

[Cay94a] John Caywood. Installing thirdparty applications. Sys Admin: The Journal for UNIX Systems Administrators, 3(2): 97–??, March/April 1994. CO-DEN SYADE7. ISSN 1061-2688.

Caywood:1994:SFS

[Cay94b] John Caywood. Setting up file systems and partitions. Sys Admin: The Journal for UNIX Systems Administrators, 3(4): 18–??, July/August 1994. CO-DEN SYADE7. ISSN 1061-2688.

Collier-Brown:1996:RPU

[CB96] David Collier-Brown. Remote password update. Sys Admin: The Journal for UNIX Systems Administrators, 5(6): 63–??, June 1996. CODEN SYADE7. ISSN 1061-2688.

Cowles:1999:LLP

[CC99] Esme Cowles and Leonard K. Carson. Lessons learned from PCASSO. Sys Admin: The Journal for UNIX Systems Administrators, 8(10):43–47, October 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Chuba:2001:PFB

[CC01] Robert Chuba and Anthony Caruso. Proxy FTP without the browser. Sys Admin: The Journal for UNIX Systems Administrators, 10(7):47–48, July 2001. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Cera:1997:SPP

[Cer97] Tim Cera. Supporting PostScript printing under System V. Sys Admin: The Journal for UNIX Systems Administrators, 6(3): 7–??, March 1997. CODEN SYADE7. ISSN 1061-2688.

Cervantes:2001:DRT

[Cer01a] Primitivo Cervantes. Disaster recovery with Tivoli Storage Manager. Sys Admin: The Journal for UNIX Systems Administrators, 10(9):16, 18, 20, 22– 23, September 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Cervantes:2001:RTR

[Cer01b] Primitivo Cervantes. Real-time remote data mirroring. Sys Admin: The Journal for UNIX Systems Administrators, 10(6s): 22, 24–28, June 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Chan:2001:DID

[Cha01] Jason Chan. Distributed intrusion detection with open source tools. Sys Admin: The Journal for UNIX Systems Administrators, 10(8):20, 22, 24–25, August 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Cheek:1996:ALS

[Che96a] Matthew Cheek. Automating large Sybase database backups. Sys Admin: The Journal for UNIX Systems Administrators, 5(8):??, August 1996. CODEN SYADE7. ISSN 1061-2688.

Cheek:1996:ESD

[Che96b] Matthew Cheek. Enhanced security on Digital UNIX. Sys Admin: The Journal for UNIX Systems Administrators, 5(11): 31, 32, 35, 36, 38, 40, 42, 44, 46, November 1996. CODEN SYADE7. ISSN 1061-2688.

Cheek:1997:MPU

[Che97] Matthew Cheek. A menu program for users and admins alike. Sys Admin: The Journal for UNIX Systems Administrators, 6(2):57–??, February 1997. CO-DEN SYADE7. ISSN 1061-2688.

Cinelli:2001:UPL

[Cin01] Anthony Cinelli. Using Port-Sentry and LogCheck. Sys Admin: The Journal for UNIX Systems Administrators, 10(3): 29–31, March 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Clark:1994:UFC

[Cla94] Tom Clark. Using fack to check ufs file systems. Sys Admin: The Journal for UNIX Systems Administrators, 3(5):65– ??, September/October 1994.

CODEN SYADE7. ISSN 1061-2688.

Clay:2001:CT

[Cla01] Keith Clay. E-class tools. Sys Admin: The Journal for UNIX Systems Administrators, 10(10): ??, October 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/. Web

only.

Clifton:1997:EEM

[Cli97] Chris Clifton. Exporting environments made easy. Sys Admin: The Journal for UNIX Systems Administrators, 6(9): 49–52, 54, 56, 57, September 1997. CODEN SYADE7. ISSN 1061-2688.

Conner:1996:SUM

[Con96] Carolyn Conner. Surviving the user maintenance jungle. Sys Admin: The Journal for UNIX Systems Administrators, 5(9): 38–??, September 1996. CODEN SYADE7. ISSN 1061-2688.

Conner:1997:ADS

[Con97a] Carolyn Conner. Automating database space monitoring. Sys Admin: The Journal for UNIX Systems Administrators, 6(2): 41–??, February 1997. CODEN SYADE7. ISSN 1061-2688.

Conner:1997:SCR

[Con97b] Carolyn Conner. Security: Controlling root. Sys Admin: The Journal for UNIX Systems Administrators, 6(6):19, 20, 22– 24, 27, June 1997. CODEN SYADE7. ISSN 1061-2688.

Conrad:1999:UB

[Con99]  David Conrad.  Upgrading

BIND. Sys Admin: The Journal for UNIX Systems Administrators, 8(12):16, 18, 20, 23– 24, 26, December 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Cook:1996:CMW

[Coo96] Cornelius Cook. Configuring a minimal Web server. Sys Admin: The Journal for UNIX Systems Administrators, 5(10): 23–??, October 1996. CODEN SYADE7. ISSN 1061-2688.

Cook:1998:LHF

[Coo98] Anthony L. Cook. Linux-hosted frame relay. Sys Admin: The Journal for UNIX Systems Administrators, 7(1):8, 10, 14, 16, 18–19, January 1998. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Cooper:2000:AD

[Coo00] Clark Cooper. Alias defender. Sys Admin: The Journal for UNIX Systems Administrators, 9(10):14, 16, 18, 20, October 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Cooper:2001:UOS

[Coo01] Clark Cooper. Using one script to back up Linux and Solaris. Sys Admin: The Journal for UNIX Systems Administrators, 10(4):18, 20, April 2001. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Courington:1992:APS

[Cou92] Jeff Courington. AIX printer support: Colon files. Sys Admin: The Journal for UNIX Systems Administrators, 1(3):29– ??, September/October 1992. CODEN SYADE7. ISSN 1061-2688.

Courington:1993:ODM

[Cou93a]  Jeff Courington.  The ob-

ject data manager (ODM) on AIX. Sys Admin: The Journal for UNIX Systems Administrators, 2(3):37–??, May/June 1993. CODEN SYADE7. ISSN 1061-2688.

Courington:1993:PEA

[Cou93b] Jeff Courington. Printer encryption on AIX. Sys Admin: The Journal for UNIX Systems Administrators, 2(4):47–??, July/ August 1993. CODEN SYADE7. ISSN 1061-2688.

Courington:1993:THA

[Cou93c] Jeff Courington. Two handy AIX tools: The failed login report and lptran. Sys Admin: The Journal for UNIX Systems Administrators, 2(3): 69–??, May/June 1993. CODEN SYADE7. ISSN 1061-2688.

Cox:1994:SCC

[Cox94] Chris Cox. suFirewall: Creating a controlled setuid environment. Sys Admin: The Journal for UNIX Systems Administrators, 3(6):24–??, November/December 1994. CODEN SYADE7. ISSN 1061-2688.

Cox:1995:RDF

[Cox95] Chris Cox. rlock: A distributed file locking utility. Sys Admin: The Journal for UNIX Systems Administrators, 4(4):8–??, July/ August 1995. CODEN SYADE7. ISSN 1061-2688.

Croteau:1998:BNE

[Cro98] Robert-Andre Croteau. BB-WARN: A notification extension for Big Brother. Sys Admin: The Journal for UNIX Systems Administrators, 7(9): 57–62, September 1998. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Chroniger:1998:WIF

[CY98] Chris Chroniger and David Yarashus. When 100Mb/s isn’t fast enough. Sys Admin: The Journal for UNIX Systems Administrators, 7(1):53–60, January 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

DAdamo:1996:SPP

[D’A96] Arthur J. D’Adamo. SelectNews.pl: Perl script for selective news delivery. Sys Admin: The Journal for UNIX Systems Administrators, 5(5):59–??, May 1996. CO-DEN SYADE7. ISSN 1061-2688.

Davis:1996:DSI

[Dav96] Jeffrey B. Davis. Detecting system intruders. Sys Admin: The Journal for UNIX Systems Administrators, 5(12):54, 56, 58, December 1996. CODEN SYADE7. ISSN 1061-2688.

Davids:1997:INS

[Dav97]  Noah Davids.  Interpreting

netstat statistics. Sys Admin: The Journal for UNIX Systems Administrators, 6(5):69–??, May 1997. CODEN SYADE7. ISSN 1061-2688.

Davids:1998:WPC

[Dav98a] Noah Davids. When packets collide. Sys Admin: The Journal for UNIX Systems Administrators, 7(6):55–59, June 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Davis:1998:SUP

[Dav98b] Eric Davis. Securing UNIX passwords. Sys Admin: The Journal for UNIX Systems Administrators, 7(8):8, 10, 12, 14–16, August 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Davids:1999:IF

[Dav99a] Noah Davids. IP forwarding.

Sys Admin: The Journal for UNIX Systems Administrators, 8(1):37–38, January 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Davids:1999:TCS

[Dav99b] Noah Davids. TCP connection states: A clue to network health. Sys Admin: The Journal for UNIX Systems Administrators, 8(7):41–48, July 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

David:2000:APM

[Dav00a]  William David.  Automati-

cally performing and monitoring backups. Sys Admin: The Journal for UNIX Systems Administrators, 9(4):27–28, 30, 32–33, April 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Davis:2000:ASU

[Dav00b] Alan C. Davis. Automating single-user backups with tape verification. Sys Admin: The Journal for UNIX Systems Administrators, 9(4):51–53, April 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Davis:2000:IWR

[Dav00c] Alan C. Davis. Implementing Web/RCS under Oracle OWS for UNIX. Sys Admin: The Journal for UNIX Systems Administrators, 9(2):8, 10, 12, 14, February 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Davis:2000:SUS

[Dav00d] Eric Davis. SNMPv3 — user security model. Sys Admin: The Journal for UNIX Systems Administrators, 9(5):43–44, 46, 48–50, May 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Davis:2000:SVA

[Dav00e] Eric Davis. SNMPv3 — view access control model. Sys Admin:

The Journal for UNIX Systems Administrators, 9(6):45–46, 48, 50–51, June 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Davids:2001:TEF

[Dav01a] Noah Davids. TCP — either fast or efficient. Sys Admin: The Journal for UNIX Systems Administrators, 10(5):53–57, May 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Davidsen:2001:MUN

[Dav01b] Bill Davidsen. Monitoring Usenet news. Sys Admin: The Journal for UNIX Systems Administrators, 10(7):??, July 2001. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Web only.

Dawson:2001:AC

[Daw01] Mark E. Dawson, Jr. AIX cloning. Sys Admin: The Journal for UNIX Systems Administrators, 10(3):49–51, March 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

DeGraw-Bertsch:2002:CFA

[DB02] Michael S. DeGraw-Bertsch.

Configuring a freebsd access point for your wireless network. Sys Admin: The Journal for UNIX Systems Administrators, 11(5):8, 10, 12, 14, May 2002. CODEN SYADE7. ISSN 1061-

DeLaVega:1997:MLA

[De 97a] Francisco M. De La Vega. Missing link: Administration by LDAP directory services. Sys Admin: The Journal for UNIX Systems Administrators, 6(9): 25, 27, 28, 30, 32, 34–38, 40, September 1997. CODEN SYADE7. ISSN 1061-2688.

DeLaVega:1997:SSW

[De 97b] Francisco M. De La Vega. Strategies for securing Webcentric intranets. Sys Admin: The Journal for UNIX Systems Administrators, 6(6):29, 31, 32, 34, 35, 37, 39–41, 43, June 1997. CODEN SYADE7. ISSN 1061-2688.

DeLaVega:1998:IPW

[De 98a] Francisco M. De La Vega. Improving performance with Web server clusters. Sys Admin: The Journal for UNIX Systems Administrators, 7(4):43–46, 48–50, April 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

DeLaVega:1998:SRS

[De 98b] Francisco M. De La Vega. Secure remote sessions. Sys Admin: The Journal for UNIX Systems Administrators, 7(1):36, 38–39, 41, 43–44, January 1998. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Deane:1992:OKG

[Dea92] Dorian Deane. Orchestrating a kinder, gentler disaster. Sys Admin: The Journal for UNIX Systems Administrators, 1(1):3– ??, May/June 1992. CODEN SYADE7. ISSN 1061-2688.

Deane:1993:SLH

[Dea93] Dorian Deane. The short life and hard times of an Internet packet. Sys Admin: The Journal for UNIX Systems Administrators, 2(1):42–??, January/February 1993. CODEN SYADE7. ISSN 1061-2688.

DeJong:2002:SSH

[DeJ02]  Scott DeJong.  Supporting

screened hosts with BIND 9.x view. Sys Admin: The Journal for UNIX Systems Administrators, 11(5):37–38, 40–44, May 2002. CODEN SYADE7. ISSN 1061-2688.

Dent:1999:ISM

[Den99a] Kyle Dent. IBM’s secure mailer.

Sys Admin: The Journal for UNIX Systems Administrators, 8(12):8, 10, 12, 14, December 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Dent:1999:SA

[Den99b] Kyle Dent. Securing Apache.

Sys Admin: The Journal for UNIX Systems Administrators, 8(5):35–39, May 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Diamond:2000:MSW

[Dia00] Ben Diamond. How to make a Solaris 2.5.1 workstation support PPP dial-up from Windows. Sys Admin: The Journal for UNIX Systems Administrators, 9(12):44–46, 48–50, 52, December 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Dijker:2000:CSA

[Dij00] Barbara Dijker. Careers in system administration. Sys Admin: The Journal for UNIX Systems Administrators, 9(9): ??, September 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Dilworth:2002:UEP

[Dil02] Bob Dilworth. Using email to perform UNIX system monitoring and control. Sys Admin: The Journal for UNIX Systems Administrators, 11(5): 31–32, 34–36, May 2002. CO-DEN SYADE7. ISSN 1061-2688.

Donkers:1995:MSM

[Don95a]  Arthur Donkers.  Managing

sendmail in a mixed UUCP/ PPP environment. Sys Admin: The Journal for UNIX Systems Administrators, 4(3):8–??, May/ June 1995. CODEN SYADE7. ISSN 1061-2688.

Donkers:1995:UWI

[Don95b] Arthur Donkers. A UNIX/ Windows 95 implementation. Sys Admin: The Journal for UNIX Systems Administrators, 4(5):27–??, September/October 1995. CODEN SYADE7. ISSN 1061-2688.

Donkers:1996:AST

[Don96a] Arthur Donkers. Assorted security tips for UNIX. Sys Admin: The Journal for UNIX Systems Administrators, 5(11):48, 50, 52, 54–58, November 1996. CODEN SYADE7. ISSN 1061-2688.

Donkers:1996:AYN

[Don96b] Arthur Donkers. Automating your network maintenance. Sys Admin: The Journal for UNIX Systems Administrators, 5(5): 35–??, May 1996. CODEN SYADE7. ISSN 1061-2688.

Donkers:1996:BLF

[Don96c] Arthur Donkers. Building a Linux firewall. Sys Admin: The Journal for UNIX Systems Administrators, 5(6):12–??, June 1996. CODEN SYADE7. ISSN 1061-2688.

Donkers:1996:BSW

[Don96d] Arthur Donkers. Building a secure Web site. Sys Admin: The Journal for UNIX Systems Administrators, 5(10):8–??, October 1996. CODEN SYADE7. ISSN 1061-2688.

Donkers:1996:CUI

[Don96e] Arthur Donkers. Creating user islands. Sys Admin: The Journal for UNIX Systems Administrators, 5(9):61–??, September 1996. CODEN SYADE7. ISSN 1061-2688.

Donkers:1996:EYW

[Don96f] Arthur Donkers. Enhance your Web-Based GUI with Java. Sys Admin: The Journal for UNIX Systems Administrators, 5(7): 35, 36, 39, 40, 42, 43, 45, 47, 50, July 1996. CODEN SYADE7. ISSN 1061-2688.

Donkers:1996:LIS

[Don96g] Arthur Donkers. Linux as an Internet server. Sys Admin: The Journal for UNIX Systems Administrators, 5(1):52–??, January 1996. CODEN SYADE7. ISSN 1061-2688.

Donkers:1996:UYW

[Don96h] Arthur Donkers. Using your Web browser as a GUI. Sys Admin: The Journal for UNIX Systems Administrators, 5(3):??, March 1996. CODEN SYADE7. ISSN 1061-2688.

Donkers:1997:GUN

[Don97a] Arthur Donkers. Getting UNIX and NT on speaking terms. Sys Admin: The Journal for UNIX Systems Administrators, 6(5):8–??, May 1997. CODEN SYADE7. ISSN 1061-2688.

Donkers:1997:IL

[Don97b] Arthur Donkers. ISDN and Linux. Sys Admin: The Journal for UNIX Systems Administrators, 6(10):51–53, 55, 57–60, October 1997. CODEN SYADE7. ISSN 1061-2688.

Donkers:1997:SYW

[Don97c] Arthur Donkers. Securing your Web server. Sys Admin: The Journal for UNIX Systems Administrators, 6(8):35–40, August 1997. CODEN SYADE7. ISSN 1061-2688.

Donkers:1997:WLD

[Don97d] Arthur Donkers. Writing Linux device drivers. Sys Admin: The Journal for UNIX Systems Administrators, 6(1):8–??, January 1997. CODEN SYADE7. ISSN 1061-2688.

Donkers:1998:ADP

[Don98a] Arthur Donkers. The art and detection of port scanning. Sys Admin: The Journal for UNIX Systems Administrators, 7(11): 61–64, November 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Donkers:1998:IOP

[Don98b] Arthur Donkers. The ins and outs of packet filtering. Sys Admin: The Journal for UNIX Systems Administrators, 7(5): 33–35, May 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Donkers:1998:ISY

[Don98c] Arthur Donkers. IP spoofing:

Are you really who you say you are? Sys Admin: The Journal for UNIX Systems Administrators, 7(7):69–71, July 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Donkers:1998:LNS

[Don98d] Arthur Donkers. Looking at NT security through UNIX glasses. Sys Admin: The Journal for UNIX Systems Administrators, 7(9):NT–17–NT–20, September 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Donkers:1998:ULO

[Don98e] Arthur Donkers. Using Linux in an office environment. Sys Admin: The Journal for UNIX Systems Administrators, 7(1): 45–46, 48, 50, 51, January 1998. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Donkers:1999:LE

[Don99a] Arthur Donkers. Living on the edge. Sys Admin: The Journal for UNIX Systems Administrators, 8(1):29–30, 33–35, January 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Donkers:1999:PAS

[Don99b] Arthur Donkers. Publicly available security tools. Sys Admin: The Journal for UNIX Systems Administrators, 8(4): 61–63, April 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Donkers:1999:SWD

[Don99c] Arthur Donkers. SSL: What does it buy me? Sys Admin: The Journal for UNIX Systems Administrators, 8(2):47–48, 50– 54, February 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Donkers:2000:BTV

[Don00] Arthur Donkers. Building tunnels with VTun. Sys Admin:

The Journal for UNIX Systems Administrators, 9(10):32, 34, 36, 38, 40–41, October 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

delRio:2001:PPI

[dRV01] Miguel Armas del Río and Esteban Manchado Velázquez. PICA: Perl installation and configuration agent. Sys Admin: The Journal for UNIX Systems Administrators, 10(6):??, June 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/. Web only.

Dunham:2000:RSD

[DS00] Mike Dunham and Ed Schaefer. Recovering source from a defunct RCS system. Sys Admin: The Journal for UNIX Systems Administrators, 9(4): 55–58, April 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Dubrawsky:2000:SS

[Dub00] Ido Dubrawsky. Securing Solaris. Sys Admin: The Journal for UNIX Systems Administrators, 9(11S):26, 28–35, November 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Dubrawsky:2001:FID

[Dub01] Ido Dubrawsky. Freeware intrusion detection tools. Sys Admin: The Journal for UNIX Systems Administrators, 10(8):??, August 2001. CODEN SYADE7.

ISSN 1061-2688. URL http://

                                        www.samag.com/. Web only.

Dubrawsky:2002:WS

[Dub02]  Ido Dubrawsky.  Wireless

(in)security. Sys Admin: The Journal for UNIX Systems Administrators, 11(5):16, 18, 20, 22, May 2002. CODEN SYADE7. ISSN 1061-2688.

Dulaney:1994:COU

[Dul94a] Emmett Dulaney. Communicating with other users. Sys Admin: The Journal for UNIX Systems Administrators, 3(1):31–??, January/February 1994. CODEN SYADE7. ISSN 1061-2688.

Dulaney:1994:SUN

[Dul94b] Emmett Dulaney. Standard UNIX network diagnostic tools. Sys Admin: The Journal for UNIX Systems Administrators, 3(5):53–??, September/October 1994. CODEN SYADE7. ISSN 1061-2688.

Dulaney:1994:WIG

[Dul94c] Emmett Dulaney. When inodes go bad. Sys Admin: The Journal for UNIX Systems Administrators, 3(4):6–??, July/August 1994. CODEN SYADE7. ISSN 1061-2688.

Dulaney:1995:UF

[Dul95a] Emmett Dulaney. Understanding filesystems. Sys Admin: The Journal for UNIX Systems Administrators, 4(2):40– ??, March/April 1995. CODEN SYADE7. ISSN 1061-2688.

Dulaney:1995:URL

[Dul95b] Emmett Dulaney. Understanding run levels. Sys Admin: The Journal for UNIX Systems Administrators, 4(5):103– ??, September/October 1995. CODEN SYADE7. ISSN 1061-2688.

Dulaney:1996:HPB

[Dul96a] Emmett Dulaney. Handling permissions between UNIX and NetWare. Sys Admin: The Journal for UNIX Systems Administrators, 5(1):69–??, January 1996. CODEN SYADE7. ISSN 1061-2688.

Dulaney:1996:TSS

[Dul96b] Emmett Dulaney. Taking shell scripts beyond simple batch functionality. Sys Admin: The Journal for UNIX Systems Administrators, 5(9):8–??, September 1996. CODEN SYADE7. ISSN 1061-2688.

Dulaney:2000:EMT

[Dul00a] Emmett Dulaney. Evaluating MKS Toolkit 7.0. Sys Admin: The Journal for UNIX Systems Administrators, 9(11):24, 26, 28, 30, November 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Dulaney:2000:ELC

[Dul00b] Emmett Dulaney. Examining LPI certification. Sys Admin: The Journal for UNIX Systems Administrators, 9(9):23–24, 26, 28, 30, September 2000. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Diamond:2001:QDS

[DW01] Ben Diamond and Keil Wurl. Quick and dirty server/workstation replication using ufsdump. Sys Admin: The Journal for UNIX Systems Administrators, 10(4): 32, 34, 36, 38–39, April 2001. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Dykes:1996:ASA

[Dyk96] David Dykes. Automated system administrator’s pager. Sys Admin: The Journal for UNIX Systems Administrators, 5(5): 51–??, May 1996. CODEN SYADE7. ISSN 1061-2688.

Ericson:2000:SAN

[EK00] George Ericson and Ken Kutzer. Storage area networks — achieving safe shared data access. Sys Admin: The Journal for UNIX Systems Administrators, 9(3): 45–50, March 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Endler:1996:CSC

[End96] David Endler. Creating a secure CGI environment. Sys Admin: The Journal for UNIX Systems Administrators, 5(11):21–24, 27, 29, November 1996. CODEN SYADE7. ISSN 1061-2688.

Endler:1997:PSU

[End97] David Endler. PGP: A simple usage guide. Sys Admin:

The Journal for UNIX Systems Administrators, 6(8):21, 22, 24, 25, 27–29, 31, 32, August 1997. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/archive/608art.html.

Endler:1998:DIR

[End98] David Endler. Detecting illegal root transition in Solaris. Sys Admin: The Journal for UNIX Systems Administrators, 7(8):29–32, 34–35, August 1998. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Endler:2000:CDH

[End00] Dave Endler. Creating and deploying a honey pot. Sys Admin: The Journal for UNIX Systems Administrators, 9(1):8, 10, 12, 14, 16, 18, 20–22, 24, January 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Ess:1996:CT

[Ess96]  Robert Ess.  Configuration

tracking. Sys Admin: The Journal for UNIX Systems Administrators, 5(9):20–??, September 1996. CODEN SYADE7. ISSN 1061-2688.

Ess:1998:IME

[Ess98a] Bob Ess. iptool — maintaining /etc/hosts. Sys Admin: The Journal for UNIX Systems Administrators, 7(3):43–45, March 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Ess:1998:SSM

[Ess98b] Bob Ess. SysMon — A systems monitor. Sys Admin: The Journal for UNIX Systems Administrators, 7(10):8, 10, 12, 14, 16, 18, 20, 22–24, October 1998. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Everson:1995:CLM

[Eve95] Tom Everson. closeit: A login management tool. Sys Admin: The Journal for UNIX Systems Administrators, 4(4): 39–??, July/August 1995. CO-DEN SYADE7. ISSN 1061-2688.

Feldman:1994:TUP

[Fel94] Jonathan Feldman. A transparent UNIX/PC connection via NFS. Sys Admin: The Journal for UNIX Systems Administrators, 3(4):85–??, July/August 1994. CODEN SYADE7. ISSN 1061-2688.

Feldman:1996:BLW

[Fel96a] Jonathan Feldman. Building a Linux Web server. Sys Admin: The Journal for UNIX Systems Administrators, 5(10): 41–??, October 1996. CODEN SYADE7. ISSN 1061-2688.

Feldman:1996:SSG

[Fel96b] Jonathan Feldman. Simple security: A GroupWise/SMTP connection. Sys Admin: The Journal for UNIX Systems Administrators, 5(2):8–??, February 1996. CODEN SYADE7. ISSN 1061-2688.

Feldman:1996:ULR

[Fel96c] Jonathan Feldman. Using Linux as a router. Sys Admin: The Journal for UNIX Systems Administrators, 5(1):8–??, January 1996. CODEN SYADE7. ISSN 1061-2688.

Feldman:1997:BYI

[Fel97a] Jonathan Feldman. Babysitting your Intranet and Internet. Sys Admin: The Journal for UNIX Systems Administrators, 6(3):67–??, March 1997. CO-DEN SYADE7. ISSN 1061-2688.

Feldman:1997:EIE

[Fel97b] Jonathan Feldman. Eschewing the Internet for external data sharing. Sys Admin: The Journal for UNIX Systems Administrators, 6(1):42–??, January 1997. CODEN SYADE7. ISSN 1061-2688.

Feldman:1997:INT

[Fel97c] Jonathan Feldman. Inexpensive network troubleshooting. Sys Admin: The Journal for UNIX Systems Administrators, 6(11): 53–55, 57–58, 60, November 1997. CODEN SYADE7. ISSN 1061-2688.

Feldman:1998:SSB

[Fel98] Jonathan Feldman. Simple security: Building a SOCKS proxy server. Sys Admin: The Journal for UNIX Systems Administrators, 7(1):20, 22, 24, 27, January 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Ferri:2001:RIH

[Fer01a] Richard Ferri. Remote installation of heterogeneous Linux clusters using LUI. Sys Admin: The Journal for UNIX Systems Administrators, 10(4): 50, 52, 54–57, April 2001. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Ferri:2001:LA

[Fer01b] Richard C. Ferri. Linux affinity. Sys Admin: The Journal for UNIX Systems Administrators, 10(6s):16, 18, 20, June 2001. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Fevre:2001:FR

[Fev01] Alex Le Fevre. Firewall reporter. Sys Admin: The Journal for UNIX Systems Administrators, 10(8):??, August 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/. Web

only.

Fiedler:1992:MFO

[Fie92] David Fiedler. My favorite oneliners: Some useful shell programs. Sys Admin: The Journal for UNIX Systems Administrators, 1(1):7–??, May/June 1992. CODEN SYADE7. ISSN 1061-2688.

Flannigan:1997:ARW

[Fla97] Robert Flannigan. Adding R* to Windows NT. Sys Admin: The Journal for UNIX Systems Administrators, 6(5):45–??, May 1997. CODEN SYADE7. ISSN 1061-2688.

Freeman:1999:VII

[FM99] Joe Freeman and Ronald Mc-Carty. VPNs and IPSec: An introduction. Sys Admin: The Journal for UNIX Systems Administrators, 8(6):45–49, June 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Farmer:1999:TSA

[FPA99] Dan Farmer, Brad Powell, and Matt Archibald. TITAN — A systems administration approach to security. Sys Admin: The Journal for UNIX Systems Administrators, 8(9):8, 10, 12, 14, 16, September 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Frederick:2002:CIH

[Fre02] Karen Kent Frederick. Cisco IOS HTTP authorization vulnerability. Sys Admin: The Journal for UNIX Systems Administrators, 11(3):8, 10, 12, 14, March 2002. CODEN SYADE7. ISSN 1061-2688.

Friedl:1995:TSU

[FS95] Stephen Friedl and Ed Schaefer. Time services from the U.S. Naval Observatory. Sys Admin: The Journal for UNIX Systems Administrators, 4(4): 117–??, July/August 1995. CO-DEN SYADE7. ISSN 1061-2688.

Gaffney:1999:PS

[Gaf99] Corey Gaffney. Patching Solaris. Sys Admin: The Journal for UNIX Systems Administrators, 8(4):24, 26, 28, 30, 32–35, April 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Gagnon:1998:MDS

[Gag98] Charles Gagnon. Managing a distributed software library using LUDE. Sys Admin: The Journal for UNIX Systems Administrators, 7(11):41– 46, November 1998. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Gagne:1999:WEL

[Gag99a] Marcel Gagné. Web-enabling legacy applications. Sys Admin: The Journal for UNIX Systems Administrators, 8(6): 57–59, June 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Gagne:1999:XNA

[Gag99b] Marcel Gagné. The Xni network analysis tool. Sys Admin: The Journal for UNIX Systems Administrators, 8(9):25–26, 28, 30, 32, 34, September 1999. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Gagne:2000:EDS

[Gag00a] Marcel Gagné. EnlightenDSM — distributed systems management. Sys Admin: The Journal for UNIX Systems Administrators, 9(1):41–42, 44–46, 48, January 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Gagne:2000:LES

[Gag00b] Marcel Gagné. A Linux email server. Sys Admin: The Journal for UNIX Systems Administrators, 9(2):53–58, February 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Gagn:2000:LIG

[Gag00c] Marcel Gagné. A Linux Internet gateway. Sys Admin: The Journal for UNIX Systems Administrators, 9(6):77–79, 81–83, June 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Gagne:2002:WOS

[Gag02] Marcel Gagné. WINE: The open source way to run Windows applications. Sys Admin: The Journal for UNIX Systems Administrators, 11(1):8, 10, 12, 14, January 2002. CODEN SYADE7. ISSN 1061-2688.

Galvin:2001:RNS

[Gal01a] Peter Baer Galvin. Reliable network with Solaris^{T M}. Sys Admin: The Journal for UNIX Systems Administrators, 10(11): 47, 49–50, November 2001. CO-DEN SYADE7. ISSN 1061-2688.

Galvin:2001:SCc

[Gal01b] Peter Baer Galvin. Solaris companion. Sys Admin: The Journal for UNIX Systems Administrators, 10(12):??, December 2001. CODEN SYADE7. ISSN 1061-2688. Web only.

Galvin:2001:SCa

[Gal01c] Peter Baer Galvin. Solaris corner. Sys Admin: The Journal for UNIX Systems Administrators, 10(7):??, July 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/. Web

only.

Galvin:2001:SCb

[Gal01d] Peter Baer Galvin. Solaris corner. Sys Admin: The Journal for UNIX Systems Administrators, 10(10):??, October 2001. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.

com/. Web only.

Galvin:2002:SCa

[Gal02a] Peter Baer Galvin. Solaris companion. Sys Admin: The Journal for UNIX Systems Administrators, 11(1):??, January 2002. CODEN SYADE7. ISSN 1061-2688. Web only.

Galvin:2002:SCG

[Gal02b] Peter Baer Galvin. Solaris companion: The golden rules of Sun systems administration. Sys Admin: The Journal for UNIX Systems Administrators, 11(2): 55–56, February 2002. CODEN SYADE7. ISSN 1061-2688.

Galvin:2002:SCU

[Gal02c] Peter Baer Galvin. Solaris^{T M} companion: Upgrading to Solaris 8. Sys Admin: The Journal for UNIX Systems Administrators, 11(4):39–41, April 2002. CODEN SYADE7. ISSN 1061-2688.

Galvin:2002:SCb

[Gal02d] Peter Baer Galvin. Solaris^{T M} corner. Sys Admin: The Journal for UNIX Systems Administrators, 11(4):??, April 2002. CO-DEN SYADE7. ISSN 1061-2688. Web only.

Galvin:2002:SCSb

[Gal02e] Peter Baer Galvin. Solaris^{T M} corner: Solaris administration best practices. Sys Admin: The Journal for UNIX Systems Administrators, 11(5):61–64, May 2002. CODEN SYADE7. ISSN 1061-2688.

Galvin:2002:SCSa

[Gal02f] Peter Baer Galvin. Solaris^{T M} corner: Sun/Cobalt Qube server. Sys Admin: The Journal for UNIX Systems Administrators, 11(3):??, March 2002. CODEN SYADE7. ISSN 1061-2688. Web only.

Ganis:1994:KR

[Gan94a] Matt Ganis. Killidle revisited. Sys Admin: The Journal for UNIX Systems Administrators, 3(3):59–??, May/June 1994. CODEN SYADE7. ISSN 1061-2688.

Ganis:1994:CD

[Gan94b]  Matthew Ganis.  Configur-

ing DNS. Sys Admin: The Journal for UNIX Systems Administrators, 3(1):61–??, January/February 1994. CODEN SYADE7. ISSN 1061-2688.

Ganis:1995:IS

[Gan95]  Matt Ganis.  Implementing

SOCKS. Sys Admin: The Journal for UNIX Systems Administrators, 4(3):117–??, May/June 1995. CODEN SYADE7. ISSN 1061-2688.

Ganis:1998:SPL

[Gan98] Matt Ganis. SOCKSLIB.PL — A library for using SOCKS firewalls. Sys Admin: The Journal for UNIX Systems Administrators, 7(1):29–30, 33, 34, January 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Garrett:1996:PHS

[Gar96] Thom Garrett. PPP over highspeed serial cards. Sys Admin: The Journal for UNIX Systems Administrators, 5(5):8–??, May 1996. CODEN SYADE7. ISSN 1061-2688.

Garrett:1997:SA

[Gar97] Thom Garrett. System administration: SA = f(x). Sys Admin: The Journal for UNIX Systems Administrators, 6(8):57–59, August 1997. CODEN SYADE7. ISSN 1061-2688.

Garner:1998:UCB

[Gar98a] William L. Garner. User configuration basics. Sys Admin: The Journal for UNIX Systems Administrators, 7(11):??, November 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/. This article is

available only electronically.

Garrett:1998:IBO

[Gar98b] Thom Garrett. IPv6 — A brief overview. Sys Admin: The Journal for UNIX Systems Administrators, 7(9):47–50, 52, September 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Garrett:2000:FPS

[Gar00] Dan Garrett. Fast and painless system-disk recovery on SGI systems. Sys Admin: The Journal for UNIX Systems Administrators, 9(11):75–76, 78–80, 82– 83, November 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Garcia:2002:PEW

[Gar02] Roberto João Lopes Garcia.

Posting email to the Web with MHonArc. Sys Admin: The Journal for UNIX Systems Administrators, 11(1):49, 52– 55, January 2002. CODEN SYADE7. ISSN 1061-2688.

Gaur:1998:TSU

[Gau98] Nalneesh Gaur. Time synchronization in a UNIX/NT environment. Sys Admin: The Journal for UNIX Systems Administrators, 7(4):27–30, 33, April 1998. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Gebhards:1998:RS

[Geb98] Chet Gebhards. Restoring the Sun. Sys Admin: The Journal for UNIX Systems Administrators, 7(6):8, 10, 12, 14, 16, 18, June 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Genosa:1992:SPM

[Gen92] William Genosa. The system performance monitor. Sys Admin: The Journal for UNIX Systems Administrators, 1(4): 62–??, November/December 1992. CODEN SYADE7. ISSN 1061-2688.

Genosa:1993:FAM

[Gen93a] William Genosa. The filesystem activity monitor. Sys Admin: The Journal for UNIX Systems Administrators, 2(1):59–??, January/February 1993. CODEN SYADE7. ISSN 1061-2688.

Genosa:1993:S

[Gen93b] William Genosa. The sentinel. Sys Admin: The Journal for UNIX Systems Administrators, 2(3):59–??, May/June 1993. CODEN SYADE7. ISSN 1061-2688.

Genosa:1994:ALV

[Gen94a] Bill Genosa. AIX’s logical volume manager. Sys Admin: The Journal for UNIX Systems Administrators, 3(4):95–??, July/ August 1994. CODEN SYADE7. ISSN 1061-2688.

Genosa:1994:MPI

[Gen94b] William Genosa. Monitoring performance with iostat and vmstat. Sys Admin: The Journal for UNIX Systems Administrators, 3(2):6–??, March/April 1994. CODEN SYADE7. ISSN 1061-2688.

Genosa:1994:SDAa

[Gen94c]  William Genosa.  Sybase

database administration. Sys Admin: The Journal for UNIX Systems Administrators, 3(5): 6–??, September/October 1994. CODEN SYADE7. ISSN 1061-2688.

Genosa:1994:SDAb

[Gen94d]  William Genosa.  Sybase

database administration: Part 2. Sys Admin: The Journal for UNIX Systems Administrators, 3(6):107–??, November/December 1994. CODEN SYADE7. ISSN 1061-2688.

Genosa:1995:SDA

[Gen95]  William Genosa.  Sybase

database administration: Part 3. Sys Admin: The Journal for UNIX Systems Administrators, 4(1):??, January/February 1995. CODEN SYADE7. ISSN 1061-2688.

Gilbertson:2000:PIB

[Gil00]  Jeff Gilbertson.  Platform-

independent backup script. Sys Admin: The Journal for UNIX Systems Administrators, 9(10): 57–59, October 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Golomshtok:2001:FTS

[GN01a] Alex Golomshtok and Yefim Nodelman. Fuzzy text searches with agrep and afind. Sys Admin: The Journal for UNIX Systems Administrators, 10(6): 14, 16, 18, 20, 22, June 2001. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Golomshtok:2001:MSK

[GN01b] Alexander Golomshtok and Yefim Nodelman. Managing Solaris with Kstat. Sys Admin: The Journal for UNIX Systems Administrators, 10(10):8, 10, 12, 14, 16, 18, October 2001. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Goldman:1998:USC

[Gol98] Ross Goldman. UNIX serial communications. Sys Admin: The Journal for UNIX Systems Administrators, 7(3):8, 10, 12–13, March 1998. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Gollsneider:2001:MBW

[Gol01]  Brian Gollsneider.  Multi-

booting Windows 98, Linux, and Solaris. Sys Admin: The Journal for UNIX Systems Administrators, 10(10):20, 22–23, October 2001. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Gomez:1994:UPA

[Gom94] Carlos Francisco Gomez. UUCP + pager = automated warning system. Sys Admin: The Journal for UNIX Systems Administrators, 3(2):71–??, March/April 1994. CODEN SYADE7. ISSN 1061-2688.

Goodman:1998:NBC

[Goo98] Ira Goodman. Network backup configuration: A primer. Sys Admin: The Journal for UNIX Systems Administrators, 7(11): 29–30, 33–36, November 1998. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Goodman:1999:NBC

[Goo99] Ira Goodman. Network backup configuration: A more complex scenario. Sys Admin: The Journal for UNIX Systems Administrators, 8(4):8, 10, 12, 14, 16, 18, April 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Gottry:2001:SSP

[Got01] Ken Gottry. Successful Solaris performance tuning. Sys Admin: The Journal for UNIX Systems Administrators, 10(10): 35–38, October 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Gourley:2001:ALB

[Gou01] Don Gourley. An Apache load balancing cluster. Sys Admin: The Journal for UNIX Systems Administrators, 10(1):8, 10, 12, 14, 16, January 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Gray:1996:SSS

[Gra96] Bill Gray. SM2UR: Simpleminded system monitoring and uptime reporting. Sys Admin: The Journal for UNIX Systems Administrators, 5(1):85–??, January 1996. CODEN SYADE7. ISSN 1061-2688.

Greaves:2001:VHF

[Gre01] Tristan Greaves. Virtual hosting, FTP, and LDAP. Sys Admin: The Journal for UNIX Systems Administrators, 10(5): 43–45, May 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Geiger:2000:AUS

[GS00] Robert Geiger and John Schweitzer. Automating UNIX security monitoring. Sys Admin: The Journal for UNIX Systems Administrators, 9(11):32, 34–36, November 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Gleed:1995:USS

[GT95] Kyle Gleed and Scott Tarvainen. Using SQL in shell scripts. Sys Admin: The Journal for UNIX Systems Administrators, 4(3): 73–??, May/June 1995. CODEN SYADE7. ISSN 1061-2688.

Guttman:1996:DS

[Gut96] Uri Guttman. The directory stack. Sys Admin: The Journal for UNIX Systems Administrators, 5(4):??, April 1996. CO-DEN SYADE7. ISSN 1061-2688.

Gourley:2000:BWB

[GW00] Don Gourley and Wei Wu.

Building a Web-based Java application server with Apache JServ. Sys Admin: The Journal for UNIX Systems Administrators, 9(2):59–65, February 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Hageman:2001:GOS

[Hag01] D. Hageman. GnuPG: An open solution to data protection. Sys Admin: The Journal for UNIX Systems Administrators, 10(1): 18, 20, 22, 24–26, January 2001. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Hall:1997:TT

[Hal97] Michael Hall. Creating graphical wrappers with Tcl/Tk. Sys Admin: The Journal for UNIX Systems Administrators, 6(1): 65–??, January 1997. CODEN SYADE7. ISSN 1061-2688.

Haletky:2001:IMC

[Hal01] Edward L. Haletky. Integrating Macintosh computers into your network. Sys Admin: The Journal for UNIX Systems Administrators, 10(11):41–44, November 2001. CODEN SYADE7. ISSN 1061-2688.

Haletky:2002:IBM

[Hal02] Edward L. Haletky. Integrating backup into a multiple operating system network with Amanda. Sys Admin: The Journal for UNIX Systems Administrators, 11(4):??, April 2002. CODEN SYADE7. ISSN 1061-2688. Web only.

Hare:1992:GIS

[Har92a] Chris Hare. Getting the info — sar. Sys Admin: The Journal for UNIX Systems Administrators, 1(4):50–??, November/December 1992. CODEN SYADE7. ISSN 1061-2688.

Hare:1992:GIU

[Har92b] Chris Hare. Getting the info — u386mon. Sys Admin: The Journal for UNIX Systems Administrators, 1(3):13–??, September/October 1992. CODEN SYADE7. ISSN 1061-2688.

Hare:1992:HUP

[Har92c] Chris Hare. How UNIX password controls work. Sys Admin: The Journal for UNIX Systems Administrators, 1(1): 53–??, May/June 1992. CODEN SYADE7. ISSN 1061-2688.

Hare:1992:UFS

[Har92d] Chris Hare. The UNIX file system debugger — FSDB. Sys Admin: The Journal for UNIX Systems Administrators, 1(2):??, July/August 1992. CODEN SYADE7. ISSN 1061-2688.

Hare:1992:WDC

[Har92e] Chris Hare. Where did that core file come from? Sys Admin: The Journal for UNIX Systems Administrators, 1(1):29–??, May/ June 1992. CODEN SYADE7. ISSN 1061-2688.

Hare:1993:CSI

[Har93a] Chris Hare. C2 security: Is Big Brother watching? Sys Admin: The Journal for UNIX Systems Administrators, 2(6):6– ??, November/December 1993. CODEN SYADE7. ISSN 1061-2688.

Hare:1993:CUV

[Har93b] Chris Hare. Configuring UUCP:

Version 2. Sys Admin: The Journal for UNIX Systems Administrators, 2(3):77–??, May/ June 1993. CODEN SYADE7. ISSN 1061-2688.

Hare:1993:MLA

[Har93c] Chris Hare. Monitoring login activity. Sys Admin: The Journal for UNIX Systems Administrators, 2(5):67–??, September/October 1993. CODEN SYADE7. ISSN 1061-2688.

Hare:1993:MNS

[Har93d] Chris Hare. More network security: Equivalency. Sys Admin: The Journal for UNIX Systems Administrators, 2(4): 35–??, July/August 1993. CO-DEN SYADE7. ISSN 1061-2688.

Hare:1993:SKI

[Har93e] Chris Hare. SCCS: Keeping it under control. Sys Admin: The Journal for UNIX Systems Administrators, 2(4):63–??, July/ August 1993. CODEN SYADE7. ISSN 1061-2688.

Hare:1993:UUP

[Har93f] Chris Hare. UUCP — the user’s perspective. Sys Admin:

The Journal for UNIX Systems Administrators, 2(1):64–??, January/February 1993. CODEN SYADE7. ISSN 1061-2688.

Hare:1993:UAB

[Har93g] Chris Hare. UUCP: Administering BNU. Sys Admin: The Journal for UNIX Systems Administrators, 2(2):78–??, March/April 1993. CODEN SYADE7. ISSN 1061-2688.

Hare:1994:ANP

[Har94a] Chris Hare. Asynchronous networking: PPP and SLIP. Sys Admin: The Journal for UNIX Systems Administrators, 3(1):6– ??, January/February 1994. CO-DEN SYADE7. ISSN 1061-2688.

Hare:1994:FVN

[Har94b] Chris Hare. File version numbering. Sys Admin: The Journal for UNIX Systems Administrators, 3(5):48–??, September/October 1994. CODEN SYADE7. ISSN 1061-2688.

Hare:1994:UPM

[Har94c] Chris Hare. The UNIX process management system. Sys Admin: The Journal for UNIX Systems Administrators, 3(4): 75–??, July/August 1994. CO-DEN SYADE7. ISSN 1061-2688.

Hare:1995:NCU

[Har95a] Chris Hare. Network construction: Using a firewall. Sys Admin: The Journal for UNIX Systems Administrators, 4(2):8– ??, March/April 1995. CODEN SYADE7. ISSN 1061-2688.

Hare:1995:PPM

[Har95b] Chris Hare. procmon: A process monitor. Sys Admin: The Journal for UNIX Systems Administrators, 4(5):109–??, September/October 1995. CODEN SYADE7. ISSN 1061-2688.

Hare:1995:SNO

[Har95c] Chris Hare. Subnetting: Networking one piece at a time. Sys Admin: The Journal for UNIX Systems Administrators, 4(1):??, January/February 1995. CO-DEN SYADE7. ISSN 1061-2688.

Harold:1995:COH

[Har95d] Michael Harold. Creating online help for vi. Sys Admin: The Journal for UNIX Systems Administrators, 4(6):87– ??, November/December 1995. CODEN SYADE7. ISSN 1061-2688.

Harold:1996:PSA

[Har96] Michael Harold. Porting serial I/O applications to distributed environments. Sys Admin: The Journal for UNIX Systems Administrators, 5(1):41–??, January 1996. CODEN SYADE7. ISSN 1061-2688.

Hare:1998:ISC

[Har98] Chris Hare. IT security coming of age. Sys Admin: The Journal for UNIX Systems Administrators, 7(12):57–62, December 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Hare:2001:RUPa

[Har01a] Chris Hare. Revisiting UNIX password controls – part 1. Sys Admin: The Journal for UNIX Systems Administrators, 10(10): 30, 32–34, October 2001. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Hare:2001:RUPb

[Har01b] Chris Hare. Revisiting UNIX password controls – part 2. Sys Admin: The Journal for UNIX Systems Administrators, 10(11): 35–38, November 2001. CODEN SYADE7. ISSN 1061-2688.

Hazlewood:1997:SBC

[Haz97] Victor Hazlewood. Softwarebased cluster computing. Sys Admin: The Journal for UNIX Systems Administrators, 6(3): 27–??, March 1997. CODEN SYADE7. ISSN 1061-2688.

Hazlewood:1999:HGN

[Haz99a] Victor Hazlewood. A Holy Grail for the new millennium: No plaintext passwords. Sys Admin: The Journal for UNIX Systems Administrators, 8(12): 47–48, 50–52, December 1999. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Hazlewood:1999:UAM

[Haz99b] Victor Hazlewood. UNIX accounting magic. Sys Admin: The Journal for UNIX Systems Administrators, 8(3):8, 10, 12, 14, 16, 18, March 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Hazlewood:1997:BLU

[HD97] Victor Hazlewood and Chris Daharsh. Backups for large UNIX installations. Sys Admin: The Journal for UNIX Systems Administrators, 6(11):8, 10, 12, 15, 16, 18, 19, November 1997. CO-DEN SYADE7. ISSN 1061-2688.

Held:1996:DWS

[Hel96a] Gilbert Held. Determining a Web server connection rate to the Internet. Sys Admin: The Journal for UNIX Systems Administrators, 5(10):32–??, October 1996. CODEN SYADE7. ISSN 1061-2688.

Held:1996:PPN

[Hel96b] Gilbert Held. Performance planning in an NT environment. Sys Admin: The Journal for UNIX Systems Administrators, 5(12): 22, 25–28, December 1996. CO-DEN SYADE7. ISSN 1061-2688.

Held:1996:RIE

[Hel96c] Gilbert Held. Retrieving information via email. Sys Admin: The Journal for UNIX Systems Administrators, 5(3):??, March 1996. CODEN SYADE7. ISSN 1061-2688.

Held:1997:BLD

[Hel97a] Gilbert Held. Bandwidth and latency: The dynamic duo of network performance. Sys Admin: The Journal for UNIX Systems Administrators, 6(12):8, 10, 13– 15, December 1997. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Held:1997:FTP

[Hel97b] Gilbert Held. Firewalls: The tool to prevent network. Sys Admin: The Journal for UNIX Systems Administrators, 6(6):8, 10, 12, 13, 15, 16, June 1997. CODEN SYADE7. ISSN 1061-2688.

Held:1997:INU

[Hel97c] Gilbert Held. Integrating NT into a UNIX environment. Sys Admin: The Journal for UNIX Systems Administrators, 6(4): 27–??, April 1997. CODEN SYADE7. ISSN 1061-2688.

Held:1997:NCE

[Hel97d] Gilbert Held. Network: Communications equipment by the layer. Sys Admin: The Journal for UNIX Systems Administrators, 6(10):8, 10, 12, 13, 15, 16, 18, 19, October 1997. CODEN SYADE7. ISSN 1061-2688.

Held:1997:WNB

[Hel97e] Gilbert Held. Windows NT backup. Sys Admin: The Journal for UNIX Systems Administrators, 6(11):21–24, November 1997. CODEN SYADE7. ISSN 1061-2688.

Held:1998:ARU

[Hel98a]  Gilbert Held.  Authenticat-

ing remote users. Sys Admin: The Journal for UNIX Systems Administrators, 7(8):57, 59–62, August 1998. CODEN SYADE7.

ISSN 1061-2688. URL http://

www.samag.com/.

Held:1998:ONP

[Hel98b] Gilbert Held. Optimizing NT performance. Sys Admin: The Journal for UNIX Systems Administrators, 7(9):NT–9–NT– 15, September 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Held:1998:RB

[Hel98c] Gilbert Held. RAID basics.

Sys Admin: The Journal for UNIX Systems Administrators, 7(3):55–59, March 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Held:1998:TLB

[Hel98d] Gilbert Held. Televaulting and LAN backup. Sys Admin: The Journal for UNIX Systems Administrators, 7(11):37– 40, November 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Held:1998:UC

[Hel98e] Gilbert Held. Understanding cookies. Sys Admin: The Journal for UNIX Systems Administrators, 7(4):51–54, April 1998. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Held:1998:UV

[Hel98f] Gilbert Held. Using vLANs.

Sys Admin: The Journal for UNIX Systems Administrators, 7(12):41–42, 44, 46–47, December 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Helmer:1998:STF

[Hel98g] Guy Helmer. Security tools in FreeBSD. Sys Admin: The Journal for UNIX Systems Administrators, 7(5):8, 10, 12, 14, 16, 18, May 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Held:1999:CNH

[Hel99a] Gilbert Held. Checking network health. Sys Admin: The Journal for UNIX Systems Administrators, 8(3):47–48, 50–52, March 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Held:1999:GEP

[Hel99b] Gilbert Held. Gigabit Ethernet — the potential illusion of speed. Sys Admin: The Journal for UNIX Systems Administrators, 8(6):51–55, June 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Held:1999:RAL

[Hel99c] Gilbert Held. Router access lists. Sys Admin: The Journal for UNIX Systems Administrators, 8(11):29–34, November 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Held:1999:SL

[Hel99d] Gilbert Held. Switching by the layer. Sys Admin: The Journal for UNIX Systems Administrators, 8(12):63–67, December 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Held:1999:WNL

[Hel99e] Gilbert Held. What’s new in LAN management. Sys Admin: The Journal for UNIX Systems Administrators, 8(8):8, 10, 12, 14, 16, August 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Held:2000:CE

[Hel00a] Gilbert Held. Convergence economics. Sys Admin: The Journal for UNIX Systems Administrators, 9(6):69–72, June 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Held:2000:EC

[Hel00b] Gilbert Held. The emergence of convergence. Sys Admin: The Journal for UNIX Systems Administrators, 9(4):42, 44–46, 48–49, April 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Held:2000:IVD

[Hel00c] Gilbert Held. Implementing voice and data convergence. Sys Admin: The Journal for UNIX Systems Administrators, 9(5):

51–52, 54–55, May 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Held:2000:LQS

[Hel00d] Gilbert Held. Layer 2 quality of service. Sys Admin: The Journal for UNIX Systems Administrators, 9(10):47–48, 50, 52, 54, October 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Held:2000:QNP

[Hel00e] Gilbert Held. QoS into the network: Part 2. Sys Admin: The Journal for UNIX Systems Administrators, 9(12):71–72, 74– 75, December 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Held:2000:QNR

[Hel00f] Gilbert Held. QoS into the network: Router queuing. Sys Admin: The Journal for UNIX Systems Administrators, 9(11): 57–62, November 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Held:2000:QS

[Hel00g] Gilbert Held. Quality of service. Sys Admin: The Journal for UNIX Systems Administrators, 9(9):77–81, September 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Held:2000:RBN

[Hel00h] Gilbert Held. Router-based network defense. Sys Admin:

The Journal for UNIX Systems Administrators, 9(3):61–63, 65, March 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Held:2001:PSF

[Hel01a] Gilbert Held. Preparing for script-form attack. Sys Admin: The Journal for UNIX Systems Administrators, 10(5):??, May 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www. samag.com/. Web only; not in paper issue.

Held:2001:QTN

[Hel01b] Gilbert Held. QoS through the network. Sys Admin: The Journal for UNIX Systems Administrators, 10(1):67–68, 70, 72, January 2001. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Held:2001:WEC

[Hel01c] Gilbert Held. Why energy consumption matters. Sys Admin: The Journal for UNIX Systems Administrators, 10(9):38, 41–42, 44, September 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Hellier:2001:CYC

[Hel01d] Richard Hellier. Centralize your crontabs. Sys Admin: The Journal for UNIX Systems Administrators, 10(11):39–40, November 2001. CODEN SYADE7. ISSN 1061-2688.

Herdejurgen:2002:RIC

[Her02] Seann Herdejurgen. Redundant Internet connections using Linux. Sys Admin: The Journal for UNIX Systems Administrators, 11(1):41–47, January 2002. CODEN SYADE7. ISSN 1061-2688.

Hill:1993:HLU

[Hil93] Russ Hill. How to login to any UNIX system and get your Email. Sys Admin: The Journal for UNIX Systems Administrators, 2(5):79–??, September/October 1993. CODEN SYADE7. ISSN 1061-2688.

Hill:1995:CNU

[Hil95]  Russ Hill.  Creating new

users with perl. Sys Admin: The Journal for UNIX Systems Administrators, 4(2):121– ??, March/April 1995. CODEN SYADE7. ISSN 1061-2688.

Hill:1996:USC

[Hil96a] Michael Hill. Understanding syslog.conf. Sys Admin: The Journal for UNIX Systems Administrators, 5(12):39, 40, 42, 44, 46, 48, 50–53, December 1996. CO-DEN SYADE7. ISSN 1061-2688.

Hill:1996:HDD

[Hil96b] Michael S. Hill. How to develop a distributed filesystem model. Sys Admin: The Journal for UNIX Systems Administrators, 5(8):??, August 1996. CODEN SYADE7. ISSN 1061-2688.

Hill:1996:CSS

[Hil96c] Russ Hill. Consolidating startup shell scripts. Sys Admin: The Journal for UNIX Systems Administrators, 5(4):??, April 1996. CODEN SYADE7. ISSN 1061-2688.

Hill:1997:DP

[Hil97] Russ Hill. Disaster prevention.

Sys Admin: The Journal for UNIX Systems Administrators, 6(1):52–??, January 1997. CO-DEN SYADE7. ISSN 1061-2688.

Hill:1998:SBU

[Hil98] Russ Hill. 10 steps to becoming a UNIX administrator. Sys Admin: The Journal for UNIX Systems Administrators, 7(7):55, 57–61, July 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Hill:2000:AFE

[Hil00] Russ Hill. Automating ftp with expiring passwords. Sys Admin: The Journal for UNIX Systems Administrators, 9(8):49–51, August 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Ho:2001:ARL

[Ho01] Hiu Ho. Automatically restart login services on a remote host. Sys Admin: The Journal for UNIX Systems Administrators, 10(4):63–64, April 2001. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Hoffman:1997:CPM

[Hof97] Doris Hoffman. Color printing management in a UNIX environment. Sys Admin: The Journal for UNIX Systems Administrators, 6(3):20–??, March 1997. CODEN SYADE7. ISSN 1061-2688.

Holoch:1997:RSU

[Hol97] Rich Holoch. RPTRTR — A simple UNIX report router. Sys Admin: The Journal for UNIX Systems Administrators, 6(5):63–??, May 1997. CODEN SYADE7. ISSN 1061-2688.

Hontanon:1998:EC

[Hon98a] Ramón J. Hontañón. Encryption 101 — the choices. Sys Admin: The Journal for UNIX Systems Administrators, 7(8): 37–39, 40–44, August 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Hontanon:1998:MRA

[Hon98b] Ramon J. Hontanon. Managing remote access to your intranet. Sys Admin: The Journal for UNIX Systems Administrators, 7(4):8, 10, 12, 14, 16, April 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Horne:1993:NR

[Hor93a] Eric Horne. newping revisited. Sys Admin: The Journal for UNIX Systems Administrators, 2(5):85–??, September/October 1993. CODEN SYADE7. ISSN 1061-2688.

Horne:1993:NRH

[Hor93b] Eric T. Horne. newping: Remote host downtime detection. Sys Admin: The Journal for UNIX Systems Administrators, 2(4):5–??, July/August 1993. CODEN SYADE7. ISSN 1061-2688.

Howard:2001:ISW

[How01] John S. Howard. Introduction to Solaris Web Start Flash. Sys Admin: The Journal for UNIX Systems Administrators, 10(10): 42, 44–46, October 2001. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Husmann:2000:L

[HS00] Darrin Husmann and Andy Sackrieter. LogViewer. Sys Admin: The Journal for UNIX Systems Administrators, 9(3): 33–34, 36, 38, 40–43, March 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Hu:1998:MMF

[Hu98a] Yufan Hu. Managing multifile system backup using dump/ restore. Sys Admin: The Journal for UNIX Systems Administrators, 7(11):15–16, 18, 20, 22, 24, 26, November 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Hu:1998:RSE

[Hu98b] Yufan Hu. Readership sensitive expiration control for Usenet news. Sys Admin: The Journal for UNIX Systems Administrators, 7(8):19–20, 23–24, 26–27, August 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Hu:2000:UD

[Hu00] Yufan Hu. Up and down. Sys Admin: The Journal for UNIX Systems Administrators, 9(8): 55–56, 58–59, August 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Hughes:2001:BH

[Hug01] David Hughes. Have I been hacked? Sys Admin: The Journal for UNIX Systems Administrators, 10(8):26, 28, 30, August 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Hunter:1992:AGU

[Hun92a] Bruce Hunter. An administrator’s guide to the UNIX software jungle. Sys Admin: The Journal for UNIX Systems Administrators, 1(4):68–??, November/December 1992. CODEN SYADE7. ISSN 1061-2688.

Hunter:1992:IA

[Hun92b] Bruce H. Hunter. Installing AIX. Sys Admin: The Journal for UNIX Systems Administrators, 1(1):48–??, May/June 1992. CODEN SYADE7. ISSN 1061-2688.

Hunter:1992:PAP

[Hun92c] Bruce H. Hunter. Printer administration potpourri for largescale UNIX systems. Sys Admin:

The Journal for UNIX Systems Administrators, 1(2):??, July/ August 1992. CODEN SYADE7. ISSN 1061-2688.

Hunter:1992:UNU

[Hun92d] Bruce H. Hunter. UNIX networking in the 90s — DOS/ UNIX connectivity. Sys Admin: The Journal for UNIX Systems Administrators, 1(3):35– ??, September/October 1992. CODEN SYADE7. ISSN 1061-2688.

Hunter:1993:MSAa

[Hun93a] Bruce H. Hunter. Mail and sendmail administration: Part 1. Sys Admin: The Journal for UNIX Systems Administrators, 2(1): 53–??, January/February 1993. CODEN SYADE7. ISSN 1061-2688.

Hunter:1993:MSAb

[Hun93b] Bruce H. Hunter. Mail and sendmail administration: Part 2. Sys Admin: The Journal for UNIX Systems Administrators, 2(2):69–??, March/April 1993. CODEN SYADE7. ISSN 1061-2688.

Ibanez:1997:CFI

[Iba97] Mario A. Ibanez. A comparison of firewalls and intrusion detection systems. Sys Admin: The Journal for UNIX Systems Administrators, 6(12): 37–39, 41, December 1997. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Isaacson:1992:SSU

[Isa92] Steven G. Isaacson. sukill: Stopping unruly processes. Sys Admin: The Journal for UNIX Systems Administrators, 1(4): 31–??, November/December 1992. CODEN SYADE7. ISSN 1061-2688.

Isaacson:1993:CCU

[Isa93a] Steven G. Isaacson. checkcron: Checking for the unexpected. Sys Admin: The Journal for UNIX Systems Administrators, 2(6):71–??, November/December 1993. CODEN SYADE7. ISSN 1061-2688.

Isaacson:1993:FFS

[Isa93b] Steven G. Isaacson. fchange: A file system watchdog. Sys Admin: The Journal for UNIX Systems Administrators, 2(3):5– ??, May/June 1993. CODEN SYADE7. ISSN 1061-2688.

Isaacson:1993:MAF

[Isa93c] Steven G. Isaacson. maxtab: Automatic file pruning. Sys Admin: The Journal for UNIX Systems Administrators, 2(2): 33–??, March/April 1993. CO-DEN SYADE7. ISSN 1061-2688.

Isaacson:1994:LRK

[Isa94a] Steven G. Isaacson. Logging rm and kill requests. Sys Admin: The Journal for UNIX Systems Administrators, 3(6):14– ??, November/December 1994. CODEN SYADE7. ISSN 1061-2688.

Isaacson:1994:SNA

[Isa94b] Steven G. Isaacson. stat: New access to your file system. Sys Admin: The Journal for UNIX Systems Administrators, 3(3): 24–??, May/June 1994. CODEN SYADE7. ISSN 1061-2688.

Isaacson:1995:CYC

[Isa95] Steven G. Isaacson. Customizing your calendar program. Sys Admin: The Journal for UNIX Systems Administrators, 4(2): 91–??, March/April 1995. CO-DEN SYADE7. ISSN 1061-2688.

Isaacson:1996:FSA

[Isa96a] Steven G. Isaacson. fchange:

The Sys Admin watchdog. Sys Admin: The Journal for UNIX Systems Administrators, 5(10): 66–??, October 1996. CODEN SYADE7. ISSN 1061-2688.

Isaacson:1996:IEW

[Isa96b] Steven G. Isaacson. isascii:

The easy way to work around binary files. Sys Admin: The Journal for UNIX Systems Administrators, 5(2):18–??, February 1996. CODEN SYADE7. ISSN 1061-2688.

Isaacson:1998:WEM

[Isa98]  Steven G. Isaacson.  Web-

enabled man pages. Sys Admin: The Journal for UNIX Systems Administrators, 7(5):37–40, May 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Isaacson:1999:US

[Isa99] Steven G. Isaacson. Using screen. Sys Admin: The Journal for UNIX Systems Administrators, 8(10):23–24, 26, October 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Iyer:1997:SSP

[Iye97] Venkat Iyer. Serious shell programming. Sys Admin: The Journal for UNIX Systems Administrators, 6(4):37–??, April 1997. CODEN SYADE7. ISSN 1061-2688.

Jachim:1999:ALP

[Jac99a] Ron Jachim. Application level ping: alp. Sys Admin: The Journal for UNIX Systems Administrators, 8(1):45–48, January 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Jachim:1999:PMH

[Jac99b] Ron Jachim. Poor man’s high availability. Sys Admin: The Journal for UNIX Systems Administrators, 8(8):37–38, 40–43, August 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Jachim:2000:PID

[Jac00]  Ron Jachim.  Practical IP

and DNS tips. Sys Admin: The Journal for UNIX Systems Administrators, 9(1):63, 65–66, January 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Jamison:1994:CSR

[Jam94] Neal S. Jamison. Computer security revisited: Physical security. Sys Admin: The Journal for UNIX Systems Administrators, 3(6):8–??, November/December 1994. CODEN SYADE7. ISSN 1061-2688.

Jayaswal:1998:VNI

[Jay98] Kailash Jayaswal. Virtual network interfaces and dependent Web sites. Sys Admin: The Journal for UNIX Systems Administrators, 7(4):59–60, April 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Jensen:2001:UPM

[Jen01] Anders Østergård Jensen. Using PHP and MySQL as a Weblog. Sys Admin: The Journal for UNIX Systems Administrators, 10(5):8, 10, 12, 14, May 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Jacobs:1994:HPA

[JK94] Brad Jacobs and Thomas Kraus.

Heterogeneous password assignment. Sys Admin: The Journal for UNIX Systems Administrators, 3(2):50–??, March/April 1994. CODEN SYADE7. ISSN 1061-2688.

Johnson:1995:RBU

[Joh95] Colt Johnson. Real-time between DOS, UNIX, and MVS. Sys Admin: The Journal for UNIX Systems Administrators, 4(5):8–??, September/October 1995. CODEN SYADE7. ISSN 1061-2688.

Johnson:1996:ICB

[Joh96a] Colt Johnson. Interprocess communication between UNIX and MVS applications using sockets. Sys Admin: The Journal for UNIX Systems Administrators, 5(8):??, August 1996. CODEN SYADE7. ISSN 1061-2688.

Johnson:1996:IUM

[Joh96b] Colt S. Johnson. Interlinking UNIX and MVS. Sys Admin: The Journal for UNIX Systems Administrators, 5(3):??, March 1996. CODEN SYADE7. ISSN 1061-2688.

Johnson:1996:ANI

[Joh96c] Stefan C. Johnson. Attaching non-TCP/IP devices with telnet. Sys Admin: The Journal for UNIX Systems Administrators, 5(6):51–??, June 1996. CODEN SYADE7. ISSN 1061-2688.

Johnson:1997:TIC

[Joh97] Colton S. Johnson. TCP/IP connectivity options under NT 4.0. Sys Admin: The Journal for UNIX Systems Administrators, 6(8):8, 11, 12, 15, 16, 19, August 1997. CODEN SYADE7. ISSN 1061-2688.

Jones:1998:SO

[Jon98a] Christopher A. Jones. Shell objects. Sys Admin: The Journal for UNIX Systems Administrators, 7(3):46, 48–53, March 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Jones:1998:BWC

[Jon98b] Tim Jones. Backing up Windows clients on UNIX. Sys Admin: The Journal for UNIX Systems Administrators, 7(9): NT–1–NT–7, September 1998. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Jones:1999:IBV

[Jon99] Tim Jones. Image-based versus file-by-file backup. Sys Admin: The Journal for UNIX Systems Administrators, 8(4): 20, 22, April 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Jones:2002:SPA

[Jon02] Walt Jones. Securing publicaccess networks: Stopping the IP thieves. Sys Admin: The Journal for UNIX Systems Administrators, 11(3):35–36, 38– 39, 41–42, March 2002. CODEN SYADE7. ISSN 1061-2688.

Josephes:2000:DSN

[Jos00] Chris Josephes. Designing a scalable NNTP server network. Sys Admin: The Journal for UNIX Systems Administrators, 9(12):81–88, December 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Josephes:2001:FJB

[Jos01]  Chris Josephes.  A frame-

work for JumpStart begin/finish scripts. Sys Admin: The Journal for UNIX Systems Administrators, 10(10):39–41, October 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Josephes:2002:IMJ

[Jos02] Chris Josephes. Instant messaging with Jabber. Sys Admin: The Journal for UNIX Systems Administrators, 11(1): 23–26, January 2002. CODEN SYADE7. ISSN 1061-2688.

Kan:1998:PQS

[KCW98] Yao-Chiang Kan, Jackie Carlson, and Jen-Sheng Wang. A print quota scheme. Sys Admin: The Journal for UNIX Systems Administrators, 7(6): 37–38, 41–44, June 1998. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Khan:2001:TEI

[Kha01] Elena Khan. Tripwire in the enterprise: Integrating Tripwire into Big Brother. Sys Admin: The Journal for UNIX Systems Administrators, 10(8):8, 10, 12, 14, 16, 18–19, August 2001. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Kiesling:1998:MDR

[Kie98] Robert Kiesling. Managing documentation with RCS. Sys Admin: The Journal for UNIX Systems Administrators, 7(12): 31–32, 34–36, December 1998. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Kiesling:1999:LYB

[Kie99a] Robert Kiesling. Linux and the Y2K bug. Sys Admin: The Journal for UNIX Systems Administrators, 8(1):16, 18–20, January 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Kiesling:1999:LKC

[Kie99b] Robert Kiesling. The Linux kernel: A case study for CVS 39. Sys Admin: The Journal for UNIX Systems Administrators, 8(6):39, 41–43, June 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

King:1994:DFL

[Kin94] Clive King. A distributed file location and properties checker. Sys Admin: The Journal for UNIX Systems Administrators, 3(2):32–??, March/April 1994. CODEN SYADE7. ISSN 1061-2688.

Kinsella:1995:WCS

[Kin95] John Kinsella. A WWW-based computer support system. Sys Admin: The Journal for UNIX Systems Administrators, 4(3): 39–??, May/June 1995. CODEN SYADE7. ISSN 1061-2688.

King:2001:CNI

[Kin01] Benjamin King. A complete network information center.

Sys Admin: The Journal for UNIX Systems Administrators, 10(6):??, June 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/. Web

only.

Klauser:1998:PFS

[Kla98] Werner Klauser. The /proc file system. Sys Admin: The Journal for UNIX Systems Administrators, 7(6):32, 34–35, June 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Klauser:2000:MCW

[Kla00] Werner Klauser. MySQL for Ecommerce Web sites. Sys Admin: The Journal for UNIX Systems Administrators, 9(10): 61–63, 65, October 2000. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Klein:1997:CYT

[Kle97] Ben N. Klein. Can you trust HP-UX? Sys Admin: The Journal for UNIX Systems Administrators, 6(6):52, 54, 55, 57, 58, 60, 61, 63, 65, 66, June 1997. CO-DEN SYADE7. ISSN 1061-2688.

Knight:1992:CDU

[Kni92]  David M. Knight.  Coher-

ent 4.0: A desktop UNIX clone. Sys Admin: The Journal for UNIX Systems Administrators, 1(4):37–??, November/December 1992. CODEN SYADE7. ISSN 1061-2688.

Knox:2002:SPL

[Kno02] Thomas Knox. Solaris^{T M} patch levels. Sys Admin: The Journal for UNIX Systems Administrators, 11(2):51–53, February 2002. CODEN SYADE7. ISSN 1061-2688.

Knudstrup:1996:S

[Knu96] Eric Knudstrup. Samba. Sys Admin: The Journal for UNIX Systems Administrators, 5(6): 8–??, June 1996. CODEN SYADE7. ISSN 1061-2688.

Koensgen:1998:MAS

[Koe98] Frank Koensgen. Managing availability of SAP R/3. Sys Admin: The Journal for UNIX Systems Administrators, 7(12): 25, 26, 28, December 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Kolstad:1999:SSN

[Kol99] Rob Kolstad. 1998 SANS system, network, and security administration survey. Sys Admin: The Journal for UNIX Systems Administrators, 8(2): 29–30, 32, February 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Komarnitsky:2000:GBT

[Kom00a] Alek Komarnitsky. Gettemp — built-in temperature monitoring for Sun Enterprise servers. Sys Admin: The Journal for UNIX Systems Administrators, 9(7):??, July 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Komarnitsky:2000:NWP

[Kom00b] Alek Komarnitsky. nmap-web: Port scanning made easy. Sys Admin: The Journal for UNIX Systems Administrators, 9(10): 22, 24, 26, 28–30, October 2000. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Konarski:1993:ETC

[Kon93] Beirne Konarski. Easy troubleshooting with crash. Sys Admin: The Journal for UNIX Systems Administrators, 2(2): 61–??, March/April 1993. CO-DEN SYADE7. ISSN 1061-2688.

Koscielny:1996:WNF

[Kos96] Al Koscielny. A Windows NT fileserver with UNIX clients. Sys Admin: The Journal for UNIX Systems Administrators, 5(6): 43–??, June 1996. CODEN SYADE7. ISSN 1061-2688.

Kurtz:1999:AAP

[KP99] George Kurtz and Chris Prosise. The art of attack and penetration: Understanding your security posture. Sys Admin: The Journal for UNIX Systems Administrators, 8(3):31, 33–38, March 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

King:1994:ASL

[KR94] Clive King and Adrian Rixon. Automating symbolic link generation. Sys Admin: The Journal for UNIX Systems Administrators, 3(5):43–??, September/October 1994. CODEN SYADE7. ISSN 1061-2688.

Kramp:1999:SD

[Kra99] William Kramp. Secure by design. Sys Admin: The Journal for UNIX Systems Administrators, 8(5):41–44, May 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Kramp:2000:BM

[Kra00] Bill Kramp. Beyond MTRG.

Sys Admin: The Journal for UNIX Systems Administrators, 9(3):18, 20, 22, 24, 27, 29–30, 32, March 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Kranz:2001:RNS

[Kra01] Thomas Kranz. Redundant NICs on Solaris. Sys Admin: The Journal for UNIX Systems Administrators, 10(6): 56, 58, June 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Kuethe:2000:DNP

[Kue00] Chris Kuethe. Dial-up networking with PPP. Sys Admin: The Journal for UNIX Systems Administrators, 9(7):??, July 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Kuethe:2001:HID

[Kue01] Chris Kuethe. Homebrew intrusion detection systems. Sys Admin: The Journal for UNIX Systems Administrators, 10(8):

??, August 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/. Web

only.

Kuo:1998:UUU

[Kuo98]  Benjamin F. Kuo.  Using

UNIX utilities to create CD-ROMs. Sys Admin: The Journal for UNIX Systems Administrators, 7(12):53–56, December 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

LaPorte:2001:GQS

[LA01] David LaPorte and Kevin Amorin. Getting on the 6bone quickly with Solaris^{T M} 8. Sys Admin: The Journal for UNIX Systems Administrators, 10(11): 28, 30, 32–34, November 2001. CODEN SYADE7. ISSN 1061-2688.

Lambert:1999:ISL

[Lam99a] Patrick Lambert. Implementing security on Linux. Sys Admin: The Journal for UNIX Systems Administrators, 8(10): 67–70, October 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Lammerse:1999:CYC

[Lam99b] M. L. A. Lammerse. Can you C2? Sys Admin: The Journal for UNIX Systems Administrators, 8(9):18, 20, 22–24, September 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Laudicina:1999:E

[Lau99] Alan Laudicina. Exploits. Sys Admin: The Journal for UNIX Systems Administrators, 8(7): 55–59, July 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Laudicina:2002:NPF

[Lau02] Alan P. Laudicina. Nessus — a powerful, free remote security scanner. Sys Admin: The Journal for UNIX Systems Administrators, 11(5):24, 26, 28–30, May 2002. CODEN SYADE7. ISSN 1061-2688.

Lawrence:1996:AOC

[LB96] Danny Lawrence and Scott Beimer. Access Oracle from CGI scripts using Perl. Sys Admin: The Journal for UNIX Systems Administrators, 5(9): 46–??, September 1996. CODEN SYADE7. ISSN 1061-2688.

Eisenberg:1997:PUD

[LE97] David Liebfritz and Michael Eisenberg. Administering network information for PCs using a database. Sys Admin: The Journal for UNIX Systems Administrators, 6(1):74–??, January 1997. CODEN SYADE7. ISSN 1061-2688.

Lees:1994:HHP

[Lee94a] John Lees. A host health probe.

Sys Admin: The Journal for UNIX Systems Administrators, 3(2):17–??, March/April 1994. CODEN SYADE7. ISSN 1061-

Lees:1994:SLS

[Lee94b] John Lees. Sharing local software on a network. Sys Admin: The Journal for UNIX Systems Administrators, 3(4): 48–??, July/August 1994. CO-DEN SYADE7. ISSN 1061-2688.

Lees:1995:PBS

[Lee95a] John Lees. Printing: BSD or system V? Sys Admin: The Journal for UNIX Systems Administrators, 4(2):97– ??, March/April 1995. CODEN SYADE7. ISSN 1061-2688.

Lees:1995:SS

[Lee95b] John Lees. Signals and scripts. Sys Admin: The Journal for UNIX Systems Administrators, 4(1):??, January/February 1995. CODEN SYADE7. ISSN 1061-2688.

Lees:1995:UFO

[Lee95c] Lisa Lees. Unidentified file objects. Sys Admin: The Journal for UNIX Systems Administrators, 4(4):24–??, July/August 1995. CODEN SYADE7. ISSN 1061-2688.

Lees:1996:EYH

[Lee96] Lisa Lees. Ethernet to your home. Sys Admin: The Journal for UNIX Systems Administrators, 5(10):53–??, October 1996. CODEN SYADE7. ISSN 1061-2688.

Leghart:2001:DRR

[Leg01] Jonathan D. Leghart. Dynamic round robin. Sys Admin:

The Journal for UNIX Systems Administrators, 10(11):24, 26– 27, November 2001. CODEN SYADE7. ISSN 1061-2688.

Leibfritz:1995:UPS

[Lei95a] David Leibfritz. A UNIX print server for UNIX workstations, PCs, and macs. Sys Admin: The Journal for UNIX Systems Administrators, 4(5):86– ??, September/October 1995. CODEN SYADE7. ISSN 1061-2688.

Leisner:1995:FMT

[Lei95b] Marty Leisner. File management tools. Sys Admin: The Journal for UNIX Systems Administrators, 4(2):59–??, March/April 1995. CODEN SYADE7. ISSN 1061-2688.

Leisner:1996:MDS

[Lei96] Marty Leisner. Managing disk space. Sys Admin: The Journal for UNIX Systems Administrators, 5(2):31–??, February 1996. CODEN SYADE7. ISSN 1061-2688.

Lenk:1994:SSU

[Len94] Fred Lenk. Setting up a secure user environment with standard UNIX tools. Sys Admin: The Journal for UNIX Systems Administrators, 3(3):85–??, May/ June 1994. CODEN SYADE7. ISSN 1061-2688.

Leonard:2001:DSS

[Leo01] Charles V. Leonard. Devouring spam with Spamivore. Sys Admin: The Journal for UNIX Systems Administrators, 10(2): 39–40, 42–44, 46–47, February 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Lesko:2000:ICO

[Les00] Matt Lesko. Installing and configuring OpenSSH. Sys Admin: The Journal for UNIX Systems Administrators, 9(10): 8, 10, 12, October 2000. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Lesko:2001:E

[Les01] Matt Lesko. Ethereal. Sys Admin: The Journal for UNIX Systems Administrators, 10(11): 8, 10, 12, 14, November 2001. CODEN SYADE7. ISSN 1061-2688.

Liberti:2001:QNR

[Lib01] Leo Liberti. Quick network redundancy schemes. Sys Admin: The Journal for UNIX Systems Administrators, 10(4):8, 10, 12, 14, 16, April 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Liberti:2002:WEF

[Lib02]  Leo Liberti.  Web-enabled

filesystem-based databases. Sys Admin: The Journal for UNIX Systems Administrators, 11(3): 16, 18–19, March 2002. CODEN SYADE7. ISSN 1061-2688.

Linthicum:1993:MMN

[Lin93a] David S. Linthicum. Making the most of NFS. Sys Admin: The Journal for UNIX Systems Administrators, 2(4):18–??, July/ August 1993. CODEN SYADE7. ISSN 1061-2688.

Linthicum:1993:TX

[Lin93b] David S. Linthicum. Tuning for X. Sys Admin: The Journal for UNIX Systems Administrators, 2(3):13–??, May/June 1993. CODEN SYADE7. ISSN 1061-2688.

Lin:1999:UEC

[Lin99] Clay Lin. Using expect to configure Cisco routers. Sys Admin: The Journal for UNIX Systems Administrators, 8(3):23–24, 27, 29–30, March 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Liu:1996:LEC

[Liu96] Huiqi Liu. Low-cost email connection between UNIX and Macs. Sys Admin: The Journal for UNIX Systems Administrators, 5(6):37–??, June 1996. CO-DEN SYADE7. ISSN 1061-2688.

Lirov:1998:CHD

[LMBMR98] Yuval Lirov, Dave Macolino, Martha Ben-Michael, and Andrew Rieger. Centaur: Hybrid desktops with manageability. Sys Admin: The Journal for UNIX Systems Administrators, 7(8):47, 49, 51–56, August 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Lehr:1997:RPI

[LS97] Ed Lehr and Chris Schultz. RAID performance issues. Sys Admin: The Journal for UNIX Systems Administrators, 6(12): 16, 18–20, 22, December 1997. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Lucas:1999:MPL

[Luc99] Michael Lucas. Maintaining patch levels with open source BSDs. Sys Admin: The Journal for UNIX Systems Administrators, 8(9):35–36, 38, 40, 42, September 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Lucas:2000:FFT

[Luc00a] Michael Lucas. FreeBSD firewall tools & techniques. Sys Admin: The Journal for UNIX Systems Administrators, 9(6):32, 34, 36– 37, 39, 41–42, 44, June 2000. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Lucas:2000:FSL

[Luc00b] Michael Lucas. FreeBSD for the SVR4/Linux administrator. Sys Admin: The Journal for UNIX Systems Administrators, 9(3):8, 10, 12, 14, 16, March 2000. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Lucas:2000:FSI

[Luc00c]  Michael Lucas.  FreeBSD’s

sysctl interface. Sys Admin: The Journal for UNIX Systems Administrators, 9(9): 16, 18, 22, September 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Lucas:2000:LUF

[Luc00d] Michael Lucas. Linux under FreeBSD. Sys Admin: The Journal for UNIX Systems Administrators, 9(1):26, 29–30, 32, January 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Lymn:2001:CDF

[Lym01a] Brett Lymn. Coda – the disconnectable file system. Sys Admin: The Journal for UNIX Systems Administrators, 10(7):??, July 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/. Web only.

Lymn:2001:WDWa

[Lym01b] Brett Lymn. What to do when the server doesn’t serve. Sys Admin: The Journal for UNIX Systems Administrators, 10(2): 24, 26, 28, February 2001. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Lymn:2001:WDWb

[Lym01c] Brett Lymn. What to do when the server doesn’t serve – duplicating data. Sys Admin: The Journal for UNIX Systems Administrators, 10(3):22, 24, 26, March 2001. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Lymn:2001:WDWc

[Lym01d] Brett Lymn. What to do when the server doesn’t serve – using CODA. Sys Admin: The Journal for UNIX Systems Administrators, 10(4):22, 24, 26, April 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Lymn:2002:WBP

[Lym02] Brett Lymn. Web-based printer management. Sys Admin: The Journal for UNIX Systems Administrators, 11(2):20, 22– 23, February 2002. CODEN SYADE7. ISSN 1061-2688.

MacGuire:1997:BBW

[Mac97] Sean MacGuire. Big brother: A Web-based UNIX system and network monitor. Sys Admin: The Journal for UNIX Systems Administrators, 6(3): 43–??, March 1997. CODEN SYADE7. ISSN 1061-2688.

Magnusson:1995:IGL

[Mag95] Lars Magnusson. An Internet gate for a local email system. Sys Admin: The Journal for UNIX Systems Administrators, 4(1):??, January/February 1995. CO-DEN SYADE7. ISSN 1061-2688.

Magnusson:1996:LNE

[Mag96] Lars Magnusson. Local news – exploring Web-based threaded discussion server functionality. Sys Admin: The Journal for UNIX Systems Administrators, 5(10):47–??, October 1996. CO-DEN SYADE7. ISSN 1061-2688.

Magnusson:1997:CS

[Mag97] Lars Magnusson. Customizing sendmail. Sys Admin: The Journal for UNIX Systems Administrators, 6(3):61–??, March 1997. CODEN SYADE7. ISSN 1061-2688.

Maletic:2001:SAF

[Mal01] Tim Maletic. Secure, automated file distribution. Sys Admin: The Journal for UNIX Systems Administrators, 10(6): 31–32, 34, 36, June 2001. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Mancardi:2001:AUA

[Man01a] Francisco Mancardi. Application usage accounting. Sys Admin: The Journal for UNIX Systems Administrators, 10(5): 58, 60–61, May 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Mansfield:2001:OBU

[Man01b] Bill Mansfield. Oracle backup for UNIX administrators. Sys Admin: The Journal for UNIX Systems Administrators, 10(5): 24, 26, 28–29, May 2001. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Marsh:2000:PRL

[Mar00] Matthew Marsh. Policy routing in Linux. Sys Admin: The Journal for UNIX Systems Administrators, 9(1):49–50, 52–58, January 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Marsh:2001:AAD

[Mar01a] Jeff Marsh. AIX alternate disk installation. Sys Admin: The Journal for UNIX Systems Administrators, 10(6s):33–37, June 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Marxer:2001:AUD

[Mar01b] Roman Marxer. Automatic UNIX documentation with unixdoc. Sys Admin: The Journal for UNIX Systems Administrators, 10(6):??, June 2001. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.

com/. Web only.

Mayfield:1995:LLF

[May95] Tim Mayfield. logit: A log file manager. Sys Admin: The Journal for UNIX Systems Administrators, 4(6):8–??, November/December 1995. CODEN SYADE7. ISSN 1061-2688.

Maynard:1996:ICS

[May96] Jack Maynard. An introduction to client/server security. Sys Admin: The Journal for UNIX Systems Administrators, 5(11): 59–62, November 1996. CODEN SYADE7. ISSN 1061-2688.

Maynard:1997:USA

[May97] Jack Maynard. UNIX security auditing: A practical guide. Sys Admin: The Journal for UNIX Systems Administrators, 6(6):67–72, June 1997. CODEN SYADE7. ISSN 1061-2688.

McCormick:1998:ACC

[McC98] Bob McCormick. Automating configs with Cfengine. Sys Admin: The Journal for UNIX Systems Administrators, 7(7): 41–44, July 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

McCarty:1999:DNS

[McC99a] Ron McCarty. Domain name system design considerations. Sys Admin: The Journal for UNIX Systems Administrators, 8(11):45–48, November 1999. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

McCarty:1999:R

[McC99b] Ron McCarty. RADIUS. Sys Admin: The Journal for UNIX Systems Administrators, 8(2): 33–39, 41, February 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

McCarty:1999:WCW

[McC99c] Ron McCarty. Web caching:

Why the continued interest? Sys Admin: The Journal for UNIX Systems Administrators, 8(12): 59–62, December 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

McCarty:1999:IDS

[McC99d] Ronald McCarty. Intrusion detection strategies and design considerations. Sys Admin: The Journal for UNIX Systems Administrators, 8(9):57, 59–63, September 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

McCarty:1999:PFC

[McC99e] Ronald McCarty. Packet filtering and Cisco’s way. Sys Admin: The Journal for UNIX Systems Administrators, 8(5):8, 10, 12, 14, 16, 18, 20, May 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

McCarty:2000:FPS

[McC00a] Ron McCarty. Fast port speeds do not build optimized networks. Sys Admin: The Journal for UNIX Systems Administrators, 9(1):37–38, 40, January 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

McCarty:2000:FSH

[McC00b] Ron McCarty. FCheck: A solution to host-based intrusion detection. Sys Admin: The Journal for UNIX Systems Administrators, 9(12):61–63, December 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

McCarty:2000:IGB

[McC00c] Ron McCarty. ICMP: The good, the bad, and the ugly. Sys Admin: The Journal for UNIX Systems Administrators, 9(4):69–71, April 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

McCarty:2000:IFS

[McC00d] Ron McCarty. IP filter on Solaris. Sys Admin: The Journal for UNIX Systems Administrators, 9(11S):53–56, November 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

McCarty:2000:MPS

[McC00e] Ron McCarty. MPLS Part I:

The search for IP Quality of Service. Sys Admin: The Journal for UNIX Systems Administrators, 9(5):77–78, May 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

McCarty:2000:MPI

[McC00f] Ron McCarty. MPLS Part II:

A closer look. Sys Admin: The Journal for UNIX Systems Administrators, 9(6):73–75, June 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

McCarty:2000:NNA

[McC00g] Ron McCarty. NAT: Network address translator. Sys Admin: The Journal for UNIX Systems Administrators, 9(3):55–57, 60, March 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

McCarty:2000:NOC

[McC00h] Ron McCarty. Network operations center on-line. Sys Admin: The Journal for UNIX Systems Administrators, 9(8):73–74, 76– 77, August 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

McCarty:2000:SA

[McC00i] Ron McCarty. Sendmail alternatives. Sys Admin: The Journal for UNIX Systems Administrators, 9(9):83–85, September 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

McCarty:2000:S

[McC00j] Ron McCarty. Snort. Sys Admin: The Journal for UNIX Systems Administrators, 9(2): 49–52, February 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

McCarty:2000:VRR

[McC00k] Ron McCarty. Virtual router redundancy protocol. Sys Admin: The Journal for UNIX Systems Administrators, 9(11): 38, 40, 43, November 2000. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

McCarty:2000:WAA

[McC00l] Ron McCarty. The world according to ARP. Sys Admin: The Journal for UNIX Systems Administrators, 9(10): 83–85, October 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

McCarty:2001:LN

[McC01a] Ron McCarty. A look at ngrep. Sys Admin: The Journal for UNIX Systems Administrators, 10(5):75–76, May 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

McCarty:2001:M

[McC01b] Ron McCarty. Mailman. Sys Admin: The Journal for UNIX Systems Administrators, 10(6): 59–61, June 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

McCarty:2001:NPP

[McC01c] Ron McCarty. NFS: Part I, the protocol. Sys Admin: The Journal for UNIX Systems Administrators, 10(10):47, 49–50, October 2001. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

McCarty:2001:NPI

[McC01d] Ron McCarty. NFS Part II, usage. Sys Admin: The Journal for UNIX Systems Administrators, 10(12):??, December 2001. CODEN SYADE7. ISSN 1061-2688. Web only.

McCarty:2001:RPR

[McC01e] Ron McCarty. R.I.P. RIP? Sys Admin: The Journal for UNIX Systems Administrators, 10(3): 61–62, March 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

McCarty:2001:RSQ

[McC01f] Ron McCarty. RSVP: Signaling quality of service. Sys Admin: The Journal for UNIX Systems Administrators, 10(1): 73–74, January 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

McCarty:2002:NSA

[McC02] Ron McCarty. Name services: Another view. Sys Admin: The Journal for UNIX Systems Administrators, 11(4):49–52, April 2002. CODEN SYADE7. ISSN 1061-2688.

McDonagh:1995:ANP

[McD95] Mark McDonagh. Automated numeric paging via uucp. Sys Admin: The Journal for UNIX Systems Administrators, 4(4): 95–??, July/August 1995. CO-DEN SYADE7. ISSN 1061-2688.

McKinley:1998:AP

[McK98a] Paul McKinley. Aging passwords. Sys Admin: The Journal for UNIX Systems Administrators, 7(10):35–37, October 1998. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

McKinstry:1998:ABD

[McK98b] Jim McKinstry. Alternate boot devices. Sys Admin: The Journal for UNIX Systems Administrators, 7(11):47, 49–50, 52– 55, November 1998. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

McKinstry:1998:FR

[McK98c] Jim McKinstry. File replication. Sys Admin: The Journal for UNIX Systems Administrators, 7(2):33–37, February 1998.

CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

McKinstry:1998:RWR

[McK98d] Jim McKinstry. Real-world RAID implementation. Sys Admin: The Journal for UNIX Systems Administrators, 7(7): 63–67, July 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

McKinstry:1999:CDP

[McK99a] Jim McKinstry. I/O considerations for database performance. Sys Admin: The Journal for UNIX Systems Administrators, 8(7):49–54, July 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

McKinstry:1999:QAa

[McK99b] Jim McKinstry. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 8(11):67–69, November 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

McKinstry:1999:QAb

[McK99c] Jim McKinstry. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 8(12):69–71, December 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

McKinstry:2000:QAd

[McK00a] Jim McKinstry. Additional questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 9(4):??, April 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/. Online only.

McKinstry:2000:QAa

[McK00b] Jim McKinstry. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 9(1):73–75, January 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

McKinstry:2000:QAb

[McK00c] Jim McKinstry. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 9(2):67–69, February 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

McKinstry:2000:QAc

[McK00d] Jim McKinstry. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 9(3):67–69, March 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

McKinstry:2000:QAe

[McK00e] Jim McKinstry. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 9(5):89–91, May 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

McKinstry:2000:QAf

[McK00f] Jim McKinstry. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 9(6):89–91, June 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

McKinstry:2000:QAg

[McK00g] Jim McKinstry. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 9(7):??, July 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

McKinstry:2000:QAh

[McK00h] Jim McKinstry. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 9(8):79–80, August 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

McKinstry:2000:QAi

[McK00i] Jim McKinstry. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 9(9):87–89, September 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

McKinstry:2000:QAj

[McK00j] Jim McKinstry. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 9(10):86, 88–89, October 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

McKinstry:2000:QAk

[McK00k] Jim McKinstry. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 9(11S):64, 66– 67, November 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

McKinstry:2001:HBR

[McK01] Jim McKinstry. Host-based replication and Oracle. Sys Admin: The Journal for UNIX Systems Administrators, 10(5): 63–64, 66–69, May 2001. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

MacKinnon:1994:RSS

[MD94] Rob MacKinnon and Mark Dapoz. Remote system security: A secure net and SLIP solution. Sys Admin: The Journal for UNIX Systems Administrators, 3(4):26–??, July/August 1994. CODEN SYADE7. ISSN 1061-2688.

Mechalas:1999:MT

[Mec99a] John Mechalas. The masshosts tool. Sys Admin: The Journal for UNIX Systems Administrators, 8(3):39, 41–46, March 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Mechalas:1999:PPS

[Mec99b] John Mechalas. Practical packet sniffing. Sys Admin: The Journal for UNIX Systems Administrators, 8(9):65–71, September 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Mellman:1996:CWH

[Mel96] Mark Mellman. Creating a Web help desk with sybperl. Sys Admin: The Journal for UNIX Systems Administrators, 5(12): 59–64, December 1996. CODEN SYADE7. ISSN 1061-2688.

Meritt:1998:FTC

[Mer98a] James W. Meritt. File tamper checking. Sys Admin: The Journal for UNIX Systems Administrators, 7(8), August 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/current/web_

feature.html. This article is only available electronically.

Meritt:1998:SC

[Mer98b] James W. Meritt. Setup check.

Sys Admin: The Journal for UNIX Systems Administrators, 7(5):51–56, May 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Meritt:1999:UNS

[Mer99] James W. Meritt. Use the ’net to secure your net. Sys Admin: The Journal for UNIX Systems Administrators, 8(9): 73–74, September 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Messenger:1994:CF

[Mes94] Arthur Messenger. Configuring FTP. Sys Admin: The Journal for UNIX Systems Administrators, 3(1):45–??, January/February 1994. CODEN SYADE7. ISSN 1061-2688.

Meyer:2001:BFF

[Mey01] Andreas Meyer. Build a floppy firewall. Sys Admin: The Journal for UNIX Systems Administrators, 10(1):76, 78–79, January 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Hall:1999:LL

[mH99] Jon “maddog” Hall. Legitimizing Linux. Sys Admin: The Journal for UNIX Systems Administrators, 8(6):21–22, June 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Mickish:1998:GUS

[Mic98] Andrew Mickish. Graphing usage statistics on the Web. Sys Admin: The Journal for UNIX Systems Administrators, 7(5): 21–22, 25, 27–28, May 1998. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Mickish:1999:CAL

[Mic99] Andrew Mickish. A comprehensive approach to log files. Sys Admin: The Journal for UNIX Systems Administrators, 8(8):49–55, August 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Miles:2000:MFD

[Mil00a] Kerry Miles. Merging files without disk space. Sys Admin: The Journal for UNIX Systems Administrators, 9(8):61–63, August 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Miles:2000:L

[Mil00b] Wyman Eric Miles. LPRng. Sys Admin: The Journal for UNIX Systems Administrators, 9(6):8, 10, 12, 14, 16, 18, 22, June 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Mobily:1998:PDW

[Mob98] Tony Mobily. PHP, databases, and the Web. Sys Admin: The Journal for UNIX Systems Administrators, 7(7):24, 26, 28, 30, 32, 34, 36, 38, July 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Moore:1997:ASI

[Moo97] Rick Moore. Automating software installation. Sys Admin: The Journal for UNIX Systems Administrators, 6(9):16, 18, 19, 21, 22, September 1997. CO-DEN SYADE7. ISSN 1061-2688.

Moraca:1995:ABR

[Mor95] Bill Moraca. AIX’s backup and restore utilities. Sys Admin: The Journal for UNIX Systems Administrators, 4(4): 57–??, July/August 1995. CO-DEN SYADE7. ISSN 1061-2688.

Morris:1996:HDN

[Mor96a] Doug Morris. Hidden dangers of NFS mounting foreign filesystems. Sys Admin: The Journal for UNIX Systems Administrators, 5(8):??, August 1996. CO-DEN SYADE7. ISSN 1061-2688.

Morris:1996:SSA

[Mor96b] Doug Morris. Simple screen auto-locking mechanisms. Sys Admin: The Journal for UNIX Systems Administrators, 5(4):??, April 1996. CODEN SYADE7. ISSN 1061-2688.

Morris:1996:TSM

[Mor96c] Doug Morris. Techniques for simulating multiple group ownership. Sys Admin: The Journal for UNIX Systems Administrators, 5(10):71–??, October 1996. CODEN SYADE7. ISSN 1061-2688.

McKinstry:2000:QAl

[MR00] Jim McKinstry and Amy Rich. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 9(12): 66, 68–70, December 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

McKinstry:2001:QAa

[MR01a] Jim McKinstry and Amy Rich. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 10(1): 97–98, 100–101, January 2001. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

McKinstry:2001:QAb

[MR01b] Jim McKinstry and Amy Rich.

Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 10(2): 68, 70–71, February 2001. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

McKinstry:2001:QAc

[MR01c] Jim McKinstry and Amy Rich.

Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 10(3): 73–74, March 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

McKinstry:2001:QAd

[MR01d] Jim McKinstry and Amy Rich.

Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 10(4): 77–78, April 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

McKinstry:2001:QAe

[MR01e] Jim McKinstry and Amy Rich.

Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 10(5): 77–79, May 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

McKinstry:2001:QAf

[MR01f] Jim McKinstry and Amy Rich.

Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 10(6): 68, 70–71, June 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

McKinstry:2001:QAg

[MR01g] Jim McKinstry and Amy Rich. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 10(6s): 47–48, June 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

McKinstry:2001:QAh

[MR01h] Jim McKinstry and Amy Rich. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 10(7): 56, 58–60, July 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Morin:1999:VDT

[MS99] Lenard Morin and Ed Schaefer. Verifying dates and times for the millennium. Sys Admin: The Journal for UNIX Systems Administrators, 8(12):73–75, December 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Muhamedagic:1997:FLS

[Muh97a] Dejan Muhamedagić. A filter library system. Sys Admin: The Journal for UNIX Systems Administrators, 6(12):43–46, 48, 50–51, December 1997. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Muhamedagic:1997:PCM

[Muh97b] Dejan Muhamedagić. PPP connections made easy. Sys Admin: The Journal for UNIX Systems Administrators, 6(2): 61–63, 65, February 1997. CO-DEN SYADE7. ISSN 1061-2688. See correction [Ano97-27].

Murphy:2000:SH

[Mur00] Mike Murphy. Scripting and HTML. Sys Admin: The Journal for UNIX Systems Administrators, 9(7):??, July 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Murray:2001:IOS

[Mur01]  Mike Murray.  Installing

OpenBSD on small disks. Sys Admin: The Journal for UNIX Systems Administrators, 10(1): 28, 30, January 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Murray:2002:HF

[Mur02] Mike Murray. Halted firewalls.

Sys Admin: The Journal for UNIX Systems Administrators, 11(1):27–28, January 2002. CO-DEN SYADE7. ISSN 1061-2688.

Nakhimovsky:2001:MIM

[Nak01] Greg Nakhimovsky. Measuring and improving memory efficiency of large applications. Sys Admin: The Journal for UNIX Systems Administrators, 10(1): 51–52, 54, 56–57, January 2001. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Napier:2000:IFW

[Nap00] Duncan Napier. Introducing FreeS/WAN and IPsec. Sys Admin: The Journal for UNIX Systems Administrators, 9(11): 63, 65–69, November 2000. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Napier:2001:INL

[Nap01]  Duncan Napier.  IPTables/

NetFilter – Linux’s nextgeneration stateful packet filter. Sys Admin: The Journal for UNIX Systems Administrators, 10(12):8, 10, 12, 14, 16, December 2001. CODEN SYADE7. ISSN 1061-2688.

Napier:2002:ALI

[Nap02] Duncan Napier. Administering Linux IPSec virtual private networks. Sys Admin: The Journal for UNIX Systems Administrators, 11(3):20, 22, 25–27, 29, March 2002. CODEN SYADE7. ISSN 1061-2688.

Nassal:1997:STN

[Nas97a] Mark Nassal. Samba — tuning the NT/UNIX dance. Sys Admin: The Journal for UNIX Systems Administrators, 6(5): 27–??, May 1997. CODEN SYADE7. ISSN 1061-2688.

Nassal:1997:OBD

[Nas97b] Mark L. Nassal. The other backup: dump. Sys Admin: The Journal for UNIX Systems Administrators, 6(11):27–28, 30, 32, 34, 36–40, 42, November 1997. CODEN SYADE7. ISSN 1061-2688.

Nassal:1998:SMTa

[Nas98a] Mark Nassal. Security: A multitiered approach — part 1. Sys Admin: The Journal for UNIX Systems Administrators, 7(2): 39–40, 42–44, 46, February 1998. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Nassal:1998:SMTb

[Nas98b] Mark Nassal. Security: A multi-tiered approach — part 2. Sys Admin: The Journal for UNIX Systems Administrators, 7(3):24, 26, 28, 30–31, 33, March 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Nazario:2001:CYS

[Naz01] Jose Nazario. Checking your system logs with awk. Sys Admin: The Journal for UNIX Systems Administrators, 10(9): ??, September 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/. Web

only.

Nelson:1999:FCU

[Nel99] Steve Nelson. Fibre channel unraveled. Sys Admin: The Journal for UNIX Systems Administrators, 8(8):57–63, 65, August 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Newman:2002:ULC

[New02] Henry Newman. Using Lib C and I/O and performance. Sys Admin: The Journal for UNIX Systems Administrators, 11(4): 33–35, April 2002. CODEN SYADE7. ISSN 1061-2688.

Nice:2000:IPP

[Nic00] Steve Nice. Isolating performance problems. Sys Admin: The Journal for UNIX Systems Administrators, 9(2):33–34, 38– 41, February 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Nielsen:2000:ULD

[Nie00] Mark Nielsen. Using Linux to duplicate computers. Sys Admin: The Journal for UNIX Systems Administrators, 9(4): 16, 18, 20, 22, 25, April 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Nolan:1998:RUT

[Nol98] Ted Nolan. Report on 1998 Usenix Technical Conference. Sys Admin: The Journal for UNIX Systems Administrators, 7(9), September 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/. Pub-

lished only on the World-Wide Web.

Nichols:1994:AIO

[NS94] Carl Nichols and Ed Schaefer. Archiving Informix on-line from cron. Sys Admin: The Journal for UNIX Systems Administrators, 3(3):79–??, May/June 1994. CODEN SYADE7. ISSN 1061-2688.

OBrien:1996:RRR

[O’B96] David O’Brien. Recognizing and recovering from rootkit attacks. Sys Admin: The Journal for UNIX Systems Administrators, 5(11):8, 10, 13, 14, 16, 19, 20, November 1996. CODEN SYADE7. ISSN 1061-2688.

OGorman:2001:VKS

[O’G01]  James O’Gorman.  Viagra:

Keeping services running on BSD. Sys Admin: The Journal for UNIX Systems Administrators, 10(2):49–51, February 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

OKane:2000:K

[O’K00] James O’Kane. Kickstart. Sys Admin: The Journal for UNIX Systems Administrators, 9(1): 33–34, 36, January 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Okuntseff:1998:OWN

[Oku98] Nik Okuntseff. Overview of Windows NT security subsystem. Sys Admin: The Journal for UNIX Systems Administrators, 7(7):8, 10, 12, 14, 16, July 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Oliver:2001:AMF

[Oli01] Ross Oliver. Achieving more flexible file permissions using Solaris ACLs. Sys Admin: The Journal for UNIX Systems Administrators, 10(5):46, 48, 50–51, May 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Olson:2000:SCT

[Ols00] Adam Olson. Scaring crackers away with TCP wrapper. Sys Admin: The Journal for UNIX Systems Administrators, 9(10): 67–71, October 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Olson:2001:KYW

[Ols01a] Adam Olson. Keeping your Web content in sync. Sys Admin: The Journal for UNIX Systems Administrators, 10(2): 37–38, February 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Olson:2001:WS

[Ols01b] Adam Olson. Working with SAINT. Sys Admin: The Journal for UNIX Systems Administrators, 10(3):14, 16, 18, 20, March 2001. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Orwant:2001:CXR

[Orw01] Jon Orwant. Creating XML– RPC Web services. Sys Admin: The Journal for UNIX Systems Administrators, 10(12): P50, P52–P53, December 2001. CODEN SYADE7. ISSN 1061-2688. Perl Journal supplement.

Orellano:2001:OSO

[OY01] Leonardo Orellano and Timothy Young. Optimizing Solaris for Oracle 8.x. Sys Admin: The Journal for UNIX Systems Administrators, 10(5): 16, 18, 20, 22, May 2001. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Panattoni:1994:PSB

[Pan94] Dale Panattoni. Printer sharing between UNIX and Windows for Workgroups. Sys Admin: The Journal for UNIX Systems Administrators, 3(6):41– ??, November/December 1994. CODEN SYADE7. ISSN 1061-2688.

Papp:1995:NS

[Pap95] Andy Papp. NIS+ security. Sys Admin: The Journal for UNIX Systems Administrators, 4(1):??, January/February 1995. CO-DEN SYADE7. ISSN 1061-2688.

Pardo:2001:CFA

[Par01] Dani Pardo. A centralized and flexible antivirus solution. Sys Admin: The Journal for UNIX Systems Administrators, 10(2): 31–34, February 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Pedro:2002:DTF

[Ped02] Bruno Pedro. Deleting temporary files created by Web sites using PHP. Sys Admin: The Journal for UNIX Systems Administrators, 11(2):30, 32– 33, February 2002. CODEN SYADE7. ISSN 1061-2688.

Peterson:1994:RBR

[Pet94] Stephen Peterson. RS/6000 backup and restore scripts. Sys Admin: The Journal for UNIX Systems Administrators, 3(2): 107–??, March/April 1994. CO-DEN SYADE7. ISSN 1061-2688.

Peterson:1995:MTA

[Pet95] Stephen Peterson. Managing tty attributes in AIX. Sys Admin: The Journal for UNIX Systems Administrators, 4(2): 53–??, March/April 1995. CO-DEN SYADE7. ISSN 1061-2688.

Petersen:2001:OLO

[Pet01]  Davin Petersen.  Overcom-

ing large outlook distribution lists: An exercise in sendmail aliases. Sys Admin: The Journal for UNIX Systems Administrators, 10(2):55–57, February 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Pinson:1997:V

[Pin97] Jim Pinson. vGen. Sys Admin: The Journal for UNIX Systems Administrators, 6(12):23–24, 26, 28, 30, 33–36, December 1997. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Pipkin:1993:RSM

[Pip93] Don Pipkin. A revised sudo: Managing super-user privileges. Sys Admin: The Journal for UNIX Systems Administrators, 2(6):51–??, November/December 1993. CODEN SYADE7. ISSN 1061-2688.

Prosise:1999:AAP

[PK99] Chris Prosise and George Kurtz.

The art of attack and penetration: Defending your site. Sys Admin: The Journal for UNIX Systems Administrators, 8(4): 37–39, 41–42, 44, April 1999. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Place:2000:FMS

[Pla00] John C. Place. FreeBSD mail server with GUI. Sys Admin: The Journal for UNIX Systems Administrators, 9(9): 55–58, September 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Plonka:1997:MSA

[Plo97] Dave Plonka. Managing sys admin tasks with GNATS. Sys Admin: The Journal for UNIX Systems Administrators, 6(2): 51–??, February 1997. CODEN SYADE7. ISSN 1061-2688.

Plonka:1998:SAF

[Plo98] Dave Plonka. Sys Admin file revision control with RCS. Sys Admin: The Journal for UNIX Systems Administrators, 7(12): 8, 10, 12, 14, 16, 18, 20–22, December 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Plonka:2000:HRD

[Plo00] Dave Plonka. How I recovered data I thought I’d lost.

Sys Admin: The Journal for UNIX Systems Administrators, 9(8):26, 28–30, August 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Podnar:2001:MVL

[Pod01] Tom Podnar. Multi-vendor LAN troubleshooting. Sys Admin: The Journal for UNIX Systems Administrators, 10(7):18, 20, 22–23, July 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Pol:2001:PST

[Pol01] Koos Pol. A powerful search tool for ASCII files. Sys Admin: The Journal for UNIX Systems Administrators, 10(6):49–51, June 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Prager:2001:FLC

[Pra01a] Mark Prager. Firewall logchecking techniques. Sys Admin: The Journal for UNIX Systems Administrators, 10(8): 33–34, 36–37, August 2001. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Prager:2001:USOa

[Pra01b] Mark Prager. Useful scripts for overworked administrators. Sys Admin: The Journal for UNIX Systems Administrators, 10(6): 44, 46–48, June 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Prager:2001:USOb

[Pra01c] Mark Prager. Useful scripts for overworked administrators. Sys Admin: The Journal for UNIX Systems Administrators, 10(7):??, July 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Preston:1996:ODB

[Pre96] Curtis Preston. Oracle database backup. Sys Admin: The Journal for UNIX Systems Administrators, 5(7):8–??, July 1996. CODEN SYADE7. ISSN 1061-2688.

Preston:1998:BT

[Pre98] W. Curtis Preston. Backup techniques. Sys Admin: The Journal for UNIX Systems Administrators, 7(2):47–48, 50, 52– 54, February 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Preston:2000:EIG

[Pre00a] W. Curtis Preston. As easy as it gets. Sys Admin: The Journal for UNIX Systems Administrators, 9(7):??, July 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Preston:2000:BLT

[Pre00b] W. Curtis Preston. Backups — in a league of their own. Sys Admin: The Journal for UNIX Systems Administrators, 9(5):57–61, May 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Preston:2000:FCP

[Pre00c] W. Curtis Preston. A fibre channel primer: Part 1. Sys Admin: The Journal for UNIX Systems Administrators, 9(11): 71–74, November 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Preston:2000:WS

[Pre00d]  W. Curtis Preston.  Why

SANs? Sys Admin: The Journal for UNIX Systems Administrators, 9(9):67–72, September 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Preston:2001:BB

[Pre01a] W. Curtis Preston. Backup on a budget. Sys Admin: The Journal for UNIX Systems Administrators, 10(7):61–63, July 2001. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Preston:2001:BUSa

[Pre01b] W. Curtis Preston. Building and using a SAN: Part I. Sys Admin: The Journal for UNIX Systems Administrators, 10(5):38, 40–41, May 2001. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Preston:2001:BUSb

[Pre01c] W. Curtis Preston. Building and using a SAN (Part II). Sys Admin: The Journal for UNIX Systems Administrators, 10(6): 41–43, June 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Preston:2001:FCP

[Pre01d] W. Curtis Preston. A Fibre Channel primer: Part 2. Sys Admin: The Journal for UNIX Systems Administrators, 10(1): 87–90, January 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Preston:2001:LFS

[Pre01e] W. Curtis Preston. Large file system backup tool. Sys Admin: The Journal for UNIX Systems Administrators, 10(8):52, 54–55, August 2001. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Preston:2001:OHB

[Pre01f] W. Curtis Preston. Oracle, hot backups, and redo logs. Sys Admin: The Journal for UNIX Systems Administrators, 10(11): 53–55, November 2001. CODEN SYADE7. ISSN 1061-2688.

Preston:2001:SBB

[Pre01g]  W. Curtis Preston.  SAN

building blocks. Sys Admin: The Journal for UNIX Systems Administrators, 10(3):69– 72, March 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Preston:2001:UOB

[Pre01h] W. Curtis Preston. Understanding Oracle backup and recovery. Sys Admin: The Journal for UNIX Systems Administrators, 10(9):54, 56–58, September 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Pruett:2002:RUC

[Pru02] Mark Pruett. RDB – A UNIX command-line database. Sys Admin: The Journal for UNIX Systems Administrators, 11(1): 33–36, January 2002. CODEN SYADE7. ISSN 1061-2688.

Purdy:2001:SSA

[Pur01] Miles Purdy. Serial Storage Architecture management. Sys Admin: The Journal for UNIX Systems Administrators, 10(9): ??, September 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/. Web

only.

Quillen:1998:SSL

[Qui98] Ed Quillen. SLAMming symbolic links. Sys Admin: The Journal for UNIX Systems Administrators, 7(7):18, 20, 22, July 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Quillen:2000:SM

[Qui00] Ed Quillen. Script manager. Sys Admin: The Journal for UNIX Systems Administrators, 9(3):51–52, 54, March 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Quinton:2001:SSS

[Qui01a] Reg Quinton. Securing SNMP on Solaris. Sys Admin: The Journal for UNIX Systems Administrators, 10(7):??, July 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/. Web only.

Quinton:2001:SNH

[Qui01b] Reg Quinton. Solaris network hardening: First steps. Sys Admin: The Journal for UNIX Systems Administrators, 10(4): 59–61, April 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Raab:2001:CGF

[Raa01] Rainer Raab. Creating global functions with the Korn shell. Sys Admin: The Journal for UNIX Systems Administrators, 10(3):42, 44–47, March 2001. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Racheneur:2000:SRS

[Rac00] Didier Racheneur. Setting root SUIDed programs at work. Sys Admin: The Journal for UNIX Systems Administrators, 9(5): 63, 65–71, May 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Raja:2001:MSR

[Raj01] Moazam Raja. Managing Sun Rays in technical environments. Sys Admin: The Journal for UNIX Systems Administrators, 10(10):??, October 2001. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Web only.

Ramirez:2000:MWS

[Ram00] Carlos Ramirez. Monitoring Web server reliability. Sys Admin: The Journal for UNIX Systems Administrators, 9(11): 49–50, 53, November 2000. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Ramirez:2001:GYM

[Ram01] Carlos Ramirez. Getting your message across with Apache. Sys Admin: The Journal for UNIX Systems Administrators, 10(3): 8, 10, 12, March 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Rauschkolb:1999:WWB

[Rau99] Mark Rauschkolb. Writing a Web-based admin tool. Sys Admin: The Journal for UNIX Systems Administrators, 8(1): 55–58, January 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Rothman:2001:WFF

[RB01a] Jeffrey Rothman and John Buckman. Which OS is fastest – FreeBSD follow-up. Sys Admin: The Journal for UNIX Systems Administrators, 10(8):??, August 2001. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Rothman:2001:WFH

[RB01b] Jeffrey B. Rothman and John Buckman. Which OS is fastest for high-performance network applications? Sys Admin: The Journal for UNIX Systems Administrators, 10(7):8, 10, 12, 14, 16, July 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Reaves:1993:RBB

[Rea93] Ben Reaves. Reading beyond a bad header with tar. Sys Admin: The Journal for UNIX Systems Administrators, 2(6): 89–??, November/December 1993. CODEN SYADE7. ISSN 1061-2688.

Rehman:1997:DNS

[Reh97] Rafeeq Ur Rehman. Distributed network security through Radius. Sys Admin: The Journal for UNIX Systems Administrators, 6(10):63, 65, 66, 68–70, October 1997. CODEN SYADE7. ISSN 1061-2688.

Rehman:1999:LTP

[Reh99] Rafeeq Ur Rehman. Linux transparent proxy. Sys Admin: The Journal for UNIX Systems Administrators, 8(5):29–30, 32–34, May 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Rehman:2000:LVM

[Reh00] Rafeeq Ur Rehman. Logical volume manager for Linux. Sys Admin: The Journal for UNIX Systems Administrators, 9(8): 14, 16, 18, 20, 22, 25, August 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Rehman:2001:CRH

[Reh01a] Rafeeq Ur Rehman. Configuring Red Hat’s high-availability server. Sys Admin: The Journal for UNIX Systems Administrators, 10(4):43–44, 46, 48, April 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Rehman:2001:OSR

[Reh01b] Rafeeq Ur Rehman. Open source real-time operating systems. Sys Admin: The Journal for UNIX Systems Administrators, 10(1): 91–95, January 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Reznick:1992:OOC

[Rez92a] Larry Reznick. An older option in C for find. Sys Admin: The Journal for UNIX Systems Administrators, 1(2):??, July/ August 1992. CODEN SYADE7. ISSN 1061-2688.

Reznick:1992:SPT

[Rez92b] Larry Reznick. Setting up a printer’s terminfo definition. Sys Admin: The Journal for UNIX Systems Administrators, 1(4):5–??, November/December 1992. CODEN SYADE7. ISSN 1061-2688.

Reznick:1992:URE

[Rez92c] Larry Reznick. Using regular expressions. Sys Admin: The Journal for UNIX Systems Administrators, 1(3):59– ??, September/October 1992.

CODEN SYADE7. ISSN 1061-2688.

Reznick:1993:SHM

[Rez93a] Larry Reznick. Setting up a high-speed modem. Sys Admin: The Journal for UNIX Systems Administrators, 2(5):95– ??, September/October 1993. CODEN SYADE7. ISSN 1061-2688.

Reznick:1993:SPI

[Rez93b] Larry Reznick. Setting up a printer interface file. Sys Admin: The Journal for UNIX Systems Administrators, 2(1):5–??, January/February 1993. CODEN SYADE7. ISSN 1061-2688.

Reznick:1993:TIU

[Rez93c] Larry Reznick. Timing out idle users. Sys Admin: The Journal for UNIX Systems Administrators, 2(6):43–??, November/December 1993. CODEN SYADE7. ISSN 1061-2688.

Reznick:1993:TNO

[Rez93d] Larry Reznick. Turning numlock on. Sys Admin: The Journal for UNIX Systems Administrators, 2(2):39–??, March/April 1993. CODEN SYADE7. ISSN 1061-2688.

Reznick:1993:UCC

[Rez93e] Larry Reznick. Using cron and crontab. Sys Admin: The Journal for UNIX Systems Administrators, 2(4):29–??, July/August 1993. CODEN SYADE7. ISSN 1061-2688.

Reznick:1993:ULL

[Rez93f] Larry Reznick. Using little languages. Sys Admin: The Journal for UNIX Systems Administrators, 2(3):31–??, May/June 1993. CODEN SYADE7. ISSN 1061-2688.

Reznick:1993:WDW

[Rez93g] Larry Reznick. Which day of the week is this? Sys Admin: The Journal for UNIX Systems Administrators, 2(4):79–??, July/ August 1993. CODEN SYADE7. ISSN 1061-2688.

Reznick:1994:CUS

[Rez94a] Larry Reznick. Checking user security. Sys Admin: The Journal for UNIX Systems Administrators, 3(5):22–??, September/October 1994. CODEN SYADE7. ISSN 1061-2688.

Reznick:1994:MSI

[Rez94b] Larry Reznick. Multiport serial installation. Sys Admin: The Journal for UNIX Systems Administrators, 3(2):101– ??, March/April 1994. CODEN SYADE7. ISSN 1061-2688.

Reznick:1994:SP

[Rez94c] Larry Reznick. Setting priorities. Sys Admin: The Journal for UNIX Systems Administrators, 3(6):59–??, November/December 1994. CODEN SYADE7. ISSN 1061-2688.

Reznick:1994:SMA

[Rez94d] Larry Reznick. Smarter mail addresses. Sys Admin: The Journal for UNIX Systems Administrators, 3(4):70–??, July/August 1994. CODEN SYADE7. ISSN 1061-2688.

Reznick:1994:TU

[Rez94e] Larry Reznick. Time for UNIX.

Sys Admin: The Journal for UNIX Systems Administrators, 3(3):65–??, May/June 1994. CODEN SYADE7. ISSN 1061-2688.

Reznick:1994:WYS

[Rez94f] Larry Reznick. Weeding your system. Sys Admin: The Journal for UNIX Systems Administrators, 3(1):91–??, January/February 1994. CODEN SYADE7. ISSN 1061-2688.

Reznick:1995:CPG

[Rez95a] Larry Reznick. Coordinating password and group files. Sys Admin: The Journal for UNIX Systems Administrators, 4(2): 113–??, March/April 1995. CO-DEN SYADE7. ISSN 1061-2688.

Reznick:1995:CTS

[Rez95b]  Larry Reznick.  Creating

terminfo source from binaries. Sys Admin: The Journal for UNIX Systems Administrators, 4(4):49–??, July/August 1995. CODEN SYADE7. ISSN 1061-2688.

Reznick:1995:FDH

[Rez95c] Larry Reznick. Finding disk hogs. Sys Admin: The Journal for UNIX Systems Administrators, 4(2):79–??, March/April 1995. CODEN SYADE7. ISSN 1061-2688.

Reznick:1995:HUA

[Rez95d] Larry Reznick. Hiding UNIX applications in utility wrappers. Sys Admin: The Journal for UNIX Systems Administrators, 4(5):68–??, September/October 1995. CODEN SYADE7. ISSN 1061-2688.

Reznick:1995:PF

[Rez95e] Larry Reznick. Publisher’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 4(3):6, May/June 1995. CODEN SYADE7. ISSN 1061-2688.

Reznick:1995:RLM

[Rez95f] Larry Reznick. Remote login menus. Sys Admin: The Journal for UNIX Systems Administrators, 4(1):??, January/February 1995. CODEN SYADE7. ISSN 1061-2688.

Reznick:1995:SR

[Rez95g] Larry Reznick. Scheduled rebooting. Sys Admin: The Journal for UNIX Systems Administrators, 4(6):117–??, November/December 1995. CODEN SYADE7. ISSN 1061-2688.

Reznick:1995:SHM

[Rez95h] Larry Reznick. Setting up an HP-UX mail hub. Sys Admin: The Journal for UNIX Systems Administrators, 4(3): 123–??, May/June 1995. CO-DEN SYADE7. ISSN 1061-2688.

Ribak:2002:MLN

[Rib02] Jay Ribak. Managing logging with numerous virtual hosts in Apache. Sys Admin: The Journal for UNIX Systems Administrators, 11(2):16, 18–19, February 2002. CODEN SYADE7. ISSN 1061-2688.

Richter:1994:IER

[Ric94a] Thomas Richter. Improved error reporting in AIX. Sys Admin: The Journal for UNIX Systems Administrators, 3(2): 67–??, March/April 1994. CO-DEN SYADE7. ISSN 1061-2688.

Richter:1994:RMP

[Ric94b] Thomas Richter. A refer macro package. Sys Admin: The Journal for UNIX Systems Administrators, 3(4):107–??, July/ August 1994. CODEN SYADE7. ISSN 1061-2688.

Richter:1995:LSA

[Ric95] Thomas Richter. Login surveillance and AIX. Sys Admin: The Journal for UNIX Systems Administrators, 4(1):??, January/February 1995. CODEN SYADE7. ISSN 1061-2688.

Rice:1996:KTS

[Ric96a] Ernest H. Rice, III. Kernel tuning for system I/O under SCO UNIX. Sys Admin: The Journal for UNIX Systems Administrators, 5(1):33–??, January 1996. CODEN SYADE7. ISSN 1061-2688.

Richter:1996:PVA

[Ric96b] Thomas Richter. Password verification in AIX version 4. Sys Admin: The Journal for UNIX Systems Administrators, 5(2):

69–??, February 1996. CODEN SYADE7. ISSN 1061-2688.

Rice:1997:WBG

[Ric97] Brian Rice. A Web-based graphical front end for sar. Sys Admin: The Journal for UNIX Systems Administrators, 6(2):8– ??, February 1997. CODEN SYADE7. ISSN 1061-2688.

Rich:1998:BDR

[Ric98] Amy Rich. Backups and disaster recovery. Sys Admin: The Journal for UNIX Systems Administrators, 7(11):8, 10, 12, November 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Rich:1999:LCP

[Ric99] Amy Rich. Landmining the cracker’s playing field. Sys Admin: The Journal for UNIX Systems Administrators, 8(9): 43–44, 46, 48–49, September 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Rich:2001:QANa

[Ric01a] Amy Rich. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 10(8):61–64, August 2001. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Rich:2001:QANb

[Ric01b] Amy Rich. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 10(9):59–62, September 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Rich:2001:QAa

[Ric01c] Amy Rich. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 10(10):51–54, October 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Rich:2001:QAb

[Ric01d] Amy Rich. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 10(11):45–46, November 2001. CODEN SYADE7. ISSN 1061-2688.

Rich:2001:QAc

[Ric01e] Amy Rich. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 10(12):??, December 2001. CODEN SYADE7. ISSN 1061-2688. Web only.

Rich:2002:QAa

[Ric02a] Amy Rich. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 11(1):37–39, January 2002. CODEN SYADE7. ISSN 1061-2688.

Rich:2002:QAb

[Ric02b] Amy Rich. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 11(2):39–41, February 2002. CODEN SYADE7. ISSN 1061-2688.

Rich:2002:QAc

[Ric02c] Amy Rich. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 11(3):47–50, March 2002. CODEN SYADE7. ISSN 1061-2688.

Rich:2002:QAd

[Ric02d] Amy Rich. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 11(4):53–56, April 2002. CODEN SYADE7. ISSN 1061-2688.

Rich:2002:QAe

[Ric02e] Amy Rich. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 11(5):57–60, May 2002. CODEN SYADE7. ISSN 1061-2688.

Rieken:1993:USF

[Rie93] Bill Rieken. UNIX security facilities. Sys Admin: The Journal for UNIX Systems Administrators, 2(6):33–??, November/December 1993. CODEN SYADE7. ISSN 1061-2688.

Rieken:1994:UAF

[Rie94] Bill Rieken. UNIX accounting facilities. Sys Admin: The Journal for UNIX Systems Administrators, 3(1):107–??, January/February 1994. CODEN SYADE7. ISSN 1061-2688.

Rissmann:2000:PCE

[Ris00] Pam Rissmann. Pushing corporate email out to wireless devices. Sys Admin: The Journal for UNIX Systems Administrators, 9(7):??, July 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Radermacher:1997:ECC

[RK97] Todd Radermacher and Peter Kocks. Executable content: Controlling the potential revolution. Sys Admin: The Journal for UNIX Systems Administrators, 6(11):61–63, 65– 66, November 1997. CODEN SYADE7. ISSN 1061-2688.

Root:1998:STR

[RL98] Darrell Root and Belinda Liviero. Slow: A tool for reporting and diagnosing performance problems. Sys Admin: The Journal for UNIX Systems Administrators, 7(10):57–63, 65– 66, October 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Ronthal:2001:WJI

[Ron01] Adam Ronthal. Why Java isn’t appropriate for everything. Sys Admin: The Journal for UNIX Systems Administrators, 10(7): 52, 54–55, July 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Ruggeri:2000:SA

[Rug00] Jeff Ruggeri. Starfire administration. Sys Admin: The Journal for UNIX Systems Administrators, 9(11S):19–20, 22, 24– 25, November 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Rush:1996:INA

[Rus96] David Rush. IP networking over ATM. Sys Admin: The Journal for UNIX Systems Administrators, 5(5):22–??, May 1996. CO-DEN SYADE7. ISSN 1061-2688.

Reis:1996:PMS

[RW96] Dan Reis and Bill Welch. Performance management on SCO OpenServer. Sys Admin: The Journal for UNIX Systems Administrators, 5(12):8, 10, 14, 16, 18–21, December 1996. CODEN SYADE7. ISSN 1061-2688.

Ryan:1994:DMN

[Rya94] Patrick M. Ryan. Disk mirroring with NFS. Sys Admin: The Journal for UNIX Systems Administrators, 3(1):79–??, January/February 1994. CODEN SYADE7. ISSN 1061-2688.

Soques:1995:UDF

[SA95] Kimbol Soques and Bruce Altman. UNIX as a DOS file and print server. Sys Admin: The Journal for UNIX Systems Administrators, 4(5):45– ??, September/October 1995. CODEN SYADE7. ISSN 1061-2688.

Salvadori:1996:WSM

[Sal96] Luca Salvadori. Web site management by mail. Sys Admin: The Journal for UNIX Systems Administrators, 5(10): 57–??, October 1996. CODEN SYADE7. ISSN 1061-2688.

Salvadori:1997:GR

[Sal97a] Luca Salvadori. GNATS revisited. Sys Admin: The Journal for UNIX Systems Administrators, 6(8):53–55, August 1997. CODEN SYADE7. ISSN 1061-2688.

Salvadori:1997:HMU

[Sal97b] Luca Salvadori. Human: Managing user accounts with a friendly interface. Sys Admin: The Journal for UNIX Systems Administrators, 6(4): 57–??, April 1997. CODEN SYADE7. ISSN 1061-2688.

Salvadori:1998:ACS

[Sal98] Luca Salvadori. Access control in sensitive WWW environments. Sys Admin: The Journal for UNIX Systems Administrators, 7(9), September 1998. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.

com/. Published only on the World-Wide Web.

Salus:2001:DST

[Sal01] Peter H. Salus. Disruptions of service: Types and effects. Sys Admin: The Journal for UNIX Systems Administrators, 10(1): 81–85, January 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Sammut:2001:LKC

[Sam01] Tim Sammut. Little known Cisco IOS security features. Sys Admin: The Journal for UNIX Systems Administrators, 10(12): 17, 20, 22–23, December 2001.

CODEN SYADE7. ISSN 1061-2688.

Sand:1993:ACT

[San93a] Paul A. Sand. Automatically cleaning temporary directories. Sys Admin: The Journal for UNIX Systems Administrators, 2(2):47–??, March/April 1993. CODEN SYADE7. ISSN 1061-2688.

Sand:1993:FLU

[San93b] Paul A. Sand. A form letter utility for UNIX. Sys Admin: The Journal for UNIX Systems Administrators, 2(3):49–??, May/ June 1993. CODEN SYADE7. ISSN 1061-2688.

Santos:1998:LVA

[San98] James Santos. Logical volumes in AIX. Sys Admin: The Journal for UNIX Systems Administrators, 7(6):46, 48–54, June 1998. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Santos:1999:NTT

[San99] James Santos. Network troubleshooting tips. Sys Admin: The Journal for UNIX Systems Administrators, 8(12):53– 54, 55–58, December 1999. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Sarmiento:2001:SFU

[Sar01]  Evan Sarmiento.  Securing

FreeBSD using Jail. Sys Admin: The Journal for UNIX Systems Administrators, 10(5): 31, 32, 34, 36–37, May 2001.

CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Sarmiento:2002:SAS

[Sar02] Evan Sarmiento. Systems administration with scsh. Sys Admin: The Journal for UNIX Systems Administrators, 11(1): 16, 18–22, January 2002. CO-DEN SYADE7. ISSN 1061-2688.

Satdeva:1992:QAd

[Sat92a] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 1(2):??, July/August 1992. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1992:QAe

[Sat92b] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 1(3):5–??, September/October 1992. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1992:QAf

[Sat92c] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 1(4):23–??, November/December 1992. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1993:QAa

[Sat93a] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 2(1):13–??, January/February 1993. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1993:QAb

[Sat93b] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 2(2):75–??, March/April 1993. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1993:QAc

[Sat93c] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 2(3):63–??, May/June 1993. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1993:QAd

[Sat93d] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 2(5):63–??, September/October 1993. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1993:QAe

[Sat93e] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 2(6):25–??, November/December 1993. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1994:QAa

[Sat94a] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 3(1):101–??, January/February 1994. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1994:QAb

[Sat94b] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 3(2):91–??, March/April 1994. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1994:QAc

[Sat94c] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 3(3):109–??, May/June 1994. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1994:QAd

[Sat94d] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 3(4):91–??, July/August 1994. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1994:QAe

[Sat94e] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 3(5):117–??, September/October 1994. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1994:QAf

[Sat94f] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 3(6):117–??, November/December 1994. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1995:DNG

[Sat95a] Bjorn Satdeva. DHCP: The next generation host configuration scheme. Sys Admin: The Journal for UNIX Systems Administrators, 4(1):??, January/February 1995. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1995:QAa

[Sat95b] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 4(1):??, January/February 1995. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1995:QAb

[Sat95c] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 4(2):33–??, March/April 1995. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1995:QAc

[Sat95d] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 4(3):85–??, May/June 1995. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1995:QAd

[Sat95e] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 4(4):103–??, July/ August 1995. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1995:QAe

[Sat95f] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 4(6):97–??, November/December 1995. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1996:QAa

[Sat96a] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 5(2):65–??, February 1996. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1996:QAb

[Sat96b] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 5(3):??, March 1996. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1996:QAc

[Sat96c] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 5(4):??, April 1996. CO-DEN SYADE7. ISSN 1061-2688.

Satdeva:1996:QAd

[Sat96d] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 5(5):67–??, May 1996. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1996:QAe

[Sat96e] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 5(6):74–??, June 1996. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1996:QAf

[Sat96f] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 5(7):77–??, July 1996. CODEN SYADE7. ISSN 1061-

Satdeva:1996:QAg

[Sat96g] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 5(8):??, August 1996. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1996:QAh

[Sat96h] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 5(11):69–72, November 1996. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1996:QAi

[Sat96i] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 5(11):69–72, November 1996. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1996:QAj

[Sat96j] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 5(12):75–76, December 1996. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1997:QAa

[Sat97a] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 6(1):87–??, January 1997. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1997:QAb

[Sat97b] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 6(2):71–??, February 1997. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1997:QAc

[Sat97c] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 6(3):75–??, March 1997. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1997:QAd

[Sat97d] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 6(4):63–??, April 1997. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1997:QAe

[Sat97e] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 6(5):75–??, May 1997. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1997:QAf

[Sat97f] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 6(6):76, 78, 79, June 1997. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1997:QAh

[Sat97g] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 6(8):65–67, August 1997. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1997:QAi

[Sat97h] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 6(9):63–66, September 1997. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1997:QAj

[Sat97i] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 6(11):73–75, November 1997. CODEN SYADE7. ISSN 1061-2688.

Satdeva:1997:QAk

[Sat97j] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 6(12):65–66, December 1997. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Satdeva:1998:QAa

[Sat98a] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 7(1):61–62, January 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Satdeva:1998:QAb

[Sat98b] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 7(2):61–62, 64, February 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Satdeva:1998:QAc

[Sat98c] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 7(3):61–62, March 1998. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Satdeva:1998:QAd

[Sat98d] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 7(4):61–62, April 1998. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Satdeva:1998:QAe

[Sat98e] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 7(5):67–68, May 1998. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Satdeva:1998:QAf

[Sat98f] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 7(6):65–66, June 1998. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Satdeva:1998:QAg

[Sat98g] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 7(7):79–82, July 1998. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Satdeva:1998:QAh

[Sat98h] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 7(9):65–66, September 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Satdeva:1998:QAi

[Sat98i] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 7(10):73–75, October 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Satdeva:1998:QAj

[Sat98j] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 7(11):65–67, November 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Satdeva:1998:QAk

[Sat98k] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 7(12):77–79, December 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Satdeva:1999:QAa

[Sat99a] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 8(1):63–65, January 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Satdeva:1999:QAb

[Sat99b] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 8(2):65–66, February 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Satdeva:1999:QAc

[Sat99c] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 8(3):59–61, March 1999. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Satdeva:1999:QAd

[Sat99d] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 8(4):67, 66–68, April 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Satdeva:1999:QAe

[Sat99e] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 8(5):57–58, 60–61, May 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Satdeva:1999:QAf

[Sat99f] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 8(6):61–64, June 1999. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Satdeva:1999:QAg

[Sat99g] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 8(7):65–67, July 1999. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Satdeva:1999:QAh

[Sat99h] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 8(8):71–72, 74, August 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Satdeva:1999:QAi

[Sat99i] Bjorn Satdeva. Questions and answers. Sys Admin: The Journal for UNIX Systems Administrators, 8(9):79–80, September 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Schaefer:1996:CUF

[SB96] Ed Schaefer and Fred Brunet. Converting UNIX filenames to MS-DOS. Sys Admin: The Journal for UNIX Systems Administrators, 5(2):81–??, February 1996. CODEN SYADE7. ISSN 1061-2688.

Soriano:1999:SYW

[SB99] Ray Soriano and Gary Bahadur.

Securing your Web server. Sys Admin: The Journal for UNIX Systems Administrators, 8(5): 45–46, 48–52, May 1999. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Schaefer:1993:UUP

[Sch93] Ed Schaefer. Using the UNIX pipe in C. Sys Admin: The Journal for UNIX Systems Administrators, 2(4):39–??, July/August 1993. CODEN SYADE7. ISSN 1061-2688.

Schaefer:1994:CMD

[Sch94] Ed Schaefer. Copying and moving directories. Sys Admin: The Journal for UNIX Systems Administrators, 3(5):36– ??, September/October 1994. CODEN SYADE7. ISSN 1061-2688.

Schaefer:1995:SDT

[Sch95]  Ed Schaefer.  A server-

based date/time synchronizer for TCP/IP. Sys Admin: The Journal for UNIX Systems Administrators, 4(3):28–??, May/ June 1995. CODEN SYADE7. ISSN 1061-2688.

Schaefer:1996:SMI

[Sch96] Ed Schaefer. Sybaccess: A menu-driven interface to Sybase. Sys Admin: The Journal for UNIX Systems Administrators, 5(7):51–??, July 1996. CODEN SYADE7. ISSN 1061-2688.

Schaefer:1997:RSC

[Sch97a] Ed Schaefer. Returning a single character in a UNIX shell script. Sys Admin: The Journal for UNIX Systems Administrators, 6(4):53–??, April 1997. CODEN SYADE7. ISSN 1061-2688.

Schaffer:1997:VEF

[Sch97b] William Schaffer. Validating email forwarding. Sys Admin: The Journal for UNIX Systems Administrators, 6(3): 55–??, March 1997. CODEN SYADE7. ISSN 1061-2688.

Schulz:1997:NSI

[Sch97c] Greg P. Schulz. New storage interfaces. Sys Admin: The Journal for UNIX Systems Administrators, 6(4):8–??, April 1997. CODEN SYADE7. ISSN 1061-2688.

Schulz:1998:IFC

[Sch98] Greg P. Schulz. Integrating Fibre Channel into a SCSI world. Sys Admin: The Journal for UNIX Systems Administrators, 7(2):8, 10, 14–16, 18–19, February 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Schaefer:1999:PES

[Sch99a] Ed Schaefer. Password encryption in shell scripts. Sys Admin: The Journal for UNIX Systems Administrators, 8(1):49–54, January 1999. CODEN SYADE7.

ISSN 1061-2688. URL http://

www.samag.com/.

Schwager:1999:SES

[Sch99b] Michael Schwager. De-spamming the enterprise with Sendmail 8.9. Sys Admin: The Journal for UNIX Systems Administrators, 8(2):8, 10, 12–15, February 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Schwager:1999:HSD

[Sch99c] Michael Schwager. How spammers do it. Sys Admin: The Journal for UNIX Systems Administrators, 8(2):??, February 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www. samag.com/. This article is available only electronically.

Schwager:1999:PSP

[Sch99d] Michael Schwager. Proactive spam prevention. Sys Admin: The Journal for UNIX Systems Administrators, 8(3): 19–22, March 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Schwager:1999:SG

[Sch99e] Michael Schwager. Sendmail as gatekeeper. Sys Admin: The Journal for UNIX Systems Administrators, 8(1):23–28, January 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Schwager:1999:SRD

[Sch99f] Michael Schwager. The sendmail rulesets — A detailed look. Sys Admin: The Journal for UNIX Systems Administrators, 8(2):??, February 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/. This

article is available only electronically.

Schwager:1999:AYS

[Sch99g] Mike Schwager. Automating your site with expect. Sys Admin: The Journal for UNIX Systems Administrators, 8(11): 49–54, November 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Schwartz:1999:CRE

[Sch99h] Randal Schwartz. Compiling regular expressions. Sys Admin: The Journal for UNIX Systems Administrators, 8(10): 39–40, 42, October 1999. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Schwartz:1999:FPW

[Sch99i] Randal L. Schwartz. Figuring phone-y words. Sys Admin: The Journal for UNIX Systems Administrators, 8(12): 43–44, 46, December 1999. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Schwartz:1999:MS

[Sch99j]  Randal L. Schwartz.  The

missing symlink. Sys Admin: The Journal for UNIX Systems Administrators, 8(8):33–34, 36, August 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Schaefer:2000:PCL

[Sch00a]  Ed Schaefer.  Processing

command-line options in the shell. Sys Admin: The Journal for UNIX Systems Administrators, 9(7):??, July 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Schaefer:2000:UVT

[Sch00b] Ed Schaefer. Using vi tag files. Sys Admin: The Journal for UNIX Systems Administrators, 9(11):55–56, November 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Schuweiler:2000:BSB

[Sch00c] Greg Schuweiler. Building a SAN backup solution. Sys Admin: The Journal for UNIX Systems Administrators, 9(4):8, 10, 12, 14, April 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Schuweiler:2000:MSS

[Sch00d] Greg Schuweiler. Migrating Solaris to a storage area network. Sys Admin: The Journal for UNIX Systems Administrators, 9(8):8, 10, 12, August 2000. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Schwartz:2000:DCD

[Sch00e] Randal L. Schwartz. Deep copying, not deep secrets. Sys Admin: The Journal for UNIX Systems Administrators, 9(2): 43–46, February 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Schwartz:2000:GIL

[Sch00f] Randal L. Schwartz. Getting it to look the way you want. Sys Admin: The Journal for UNIX Systems Administrators, 9(6):53–55, June 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Schwartz:2000:LAM

[Sch00g] Randal L. Schwartz. Little acts of magic. Sys Admin: The Journal for UNIX Systems Administrators, 9(10):43–45, October 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Schwartz:2000:RDT

[Sch00h] Randal L. Schwartz. Rainy day template fun. Sys Admin: The Journal for UNIX Systems Administrators, 9(4):35, 37–38, 40, April 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Schwartz:2000:WD

[Sch00i] Randal L. Schwartz. So what’s the difference? Sys Admin: The Journal for UNIX Systems Administrators, 9(12):39–40, 42– 43, December 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Schwartz:2000:TEI

[Sch00j] Randal L. Schwartz. Taint so easy, is it? Sys Admin: The Journal for UNIX Systems Administrators, 9(8):53–54, August 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Schaefer:2001:CIZ

[Sch01a] Ed Schaefer. Communicating with Intermec and Zebra bar code printers. Sys Admin: The Journal for UNIX Systems Administrators, 10(5):??, May 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www. samag.com/. Web only; not in paper issue.

Schwartz:2001:DPR

[Sch01b] Randal L. Schwartz. Developing a Perl routine. Sys Admin: The Journal for UNIX Systems Administrators, 10(8):59–60, August 2001. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Schwartz:2001:AAC

[Sch01c] Randal L. Schwartz. It’s all about context. Sys Admin: The Journal for UNIX Systems Administrators, 10(6):53–55, June 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Schwartz:2001:MTT

[Sch01d] Randal L. Schwartz. A MIME is a terrible thing to waste. Sys Admin: The Journal for UNIX Systems Administrators, 10(4): 41–42, April 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Schwartz:2001:PIT

[Sch01e] Randal L. Schwartz. Parsing interesting things. Sys Admin: The Journal for UNIX Systems Administrators, 10(12):24, 26– 27, December 2001. CODEN SYADE7. ISSN 1061-2688.

Schwartz:2001:WA

[Sch01f] Randal L. Schwartz. What is that, anyway? Sys Admin: The Journal for UNIX Systems Administrators, 10(2): 65–67, February 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Schaefer:2002:BSU

[Sch02a] Ed Schaefer. Backup scripts

                                        from UnixReview.com. Sys Ad-

min: The Journal for UNIX Systems Administrators, 11(4): 27–29, April 2002. CODEN SYADE7. ISSN 1061-2688.

Schaefer:2002:CBB

[Sch02b] Ed Schaefer. Creating black box functions in the Korn and Bash shells. Sys Admin: The Journal for UNIX Systems Administrators, 11(3):43–46, March 2002. CODEN SYADE7. ISSN 1061-2688.

Schaefer:2002:GFC

[Sch02c] Ed Schaefer. Geographic failover for cheapskates. Sys Admin: The Journal for UNIX Systems Administrators, 11(4):30, 32, April 2002. CODEN SYADE7. ISSN 1061-2688.

Schwartz:2002:DMT

[Sch02d] Randal L. Schwartz. Doing many things, like pings. Sys Admin: The Journal for UNIX Systems Administrators, 11(5): 52, 54–55, May 2002. CODEN SYADE7. ISSN 1061-2688.

Steen:1996:ISI

[SD96] William Steen and Emmett Dulaney. Internet security information sources. Sys Admin: The Journal for UNIX Systems Administrators, 5(12):65, 66, 68, 71, 73, 74, December 1996. CO-DEN SYADE7. ISSN 1061-2688.

Sefton:1993:USN

[Sef93] Laurie Sefton. UNIX security in a networked environment. Sys Admin: The Journal for UNIX Systems Administrators, 2(1): 38–??, January/February 1993. CODEN SYADE7. ISSN 1061-2688.

Seifried:2000:C

[Sei00a] Kurt Seifried. Crypto 101. Sys Admin: The Journal for UNIX Systems Administrators, 9(5): 16, 20, 22, 24, 26–27, May 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Seifried:2000:HHI

[Sei00b] Kurt Seifried. How to hack – an introduction. Sys Admin: The Journal for UNIX Systems Administrators, 9(11): 44, 46–47, November 2000. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Seifried:2000:HHU

[Sei00c] Kurt Seifried. How to hack UNIX: Part 2. Sys Admin: The Journal for UNIX Systems Administrators, 9(12):77–80, December 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Seifried:2000:PPA

[Sei00d] Kurt Seifried. PAM — Pluggable Authentication Modules. Sys Admin: The Journal for UNIX Systems Administrators, 9(9):8, 10, 12, 14, September 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Seifried:2001:HHW

[Sei01] Kurt Seifried. How to hack Windows. Sys Admin: The Journal for UNIX Systems Administrators, 10(4):??, April 2001. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.

com/. Web only; not in paper issue.

Shalhoup:1997:WBH

[Sha97] Ron Shalhoup. A Web-based helpdesk. Sys Admin: The Journal for UNIX Systems Administrators, 6(9):41, 42, 44– 46, September 1997. CODEN SYADE7. ISSN 1061-2688.

Sherman:1998:AM

[She98] William Sherman. Advisory management. Sys Admin: The Journal for UNIX Systems Administrators, 7(10):39–43, October 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Shearer:2001:PPM

[She01] John Shearer. A Perl package for monitoring traffic. Sys Admin: The Journal for UNIX Systems Administrators, 10(9): ??, September 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/. Web

only.

Shearer:2002:TMG

[She02] John Shearer. Tracking machines with GetHost. Sys Admin: The Journal for UNIX Systems Administrators, 11(5): 45–51, May 2002. CODEN SYADE7. ISSN 1061-2688.

Sklar:2001:AEL

[Skl01] Sandor W. Sklar. The AIX error logging facility. Sys Admin: The Journal for UNIX Systems Administrators, 10(6s): 8, 10, 12–14, June 2001. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Schaefer:1995:CDB

[SL95a] Ed Schaefer and Charles Leonard.

Converting data between UNIX and windows. Sys Admin: The Journal for UNIX Systems Administrators, 4(6):67– ??, November/December 1995. CODEN SYADE7. ISSN 1061-2688.

Sherman:1995:SLD

[SL95b] Andrew B. Sherman and Yuval Lirov. Surviving large-scale distributed computing powerdowns. Sys Admin: The Journal for UNIX Systems Administrators, 4(3):93–??, May/June 1995. CODEN SYADE7. ISSN 1061-2688.

Smith:1996:PMR

[Smi96] Richard S. Smith. Performance management in an RDBMS environment. Sys Admin: The Journal for UNIX Systems Administrators, 5(12):29, 30, 32, 34, 36, 38, December 1996. CO-DEN SYADE7. ISSN 1061-2688.

Smith:1998:WBR

[Smi98] Richard S. Smith. A Web-based revision control system. Sys Admin: The Journal for UNIX Systems Administrators, 7(12): 37–40, December 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Smith:2001:IK

[Smi01a] David Smith. Implementing Kerberos. Sys Admin: The Journal for UNIX Systems Administrators, 10(12):28, 30, 32, 34–38, December 2001. CODEN SYADE7. ISSN 1061-2688.

Smith:2001:HHA

[Smi01b] Drew Smith. Homebrew high availability: Booting Linux from a RAID-1 device. Sys Admin: The Journal for UNIX Systems Administrators, 10(1):59–60, 62, 64–65, January 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Smith:2002:CBE

[Smi02a] Bryan Smith. CD backups for easy file access. Sys Admin: The Journal for UNIX Systems Administrators, 11(4):22, 24, 26, April 2002. CODEN SYADE7. ISSN 1061-2688.

Smith:2002:CA

[Smi02b] David T. Smith. Configuring Amanda. Sys Admin: The Journal for UNIX Systems Administrators, 11(4):8, 10, 12, 14, 16–17, April 2002. CODEN SYADE7. ISSN 1061-2688.

Soto:1995:CSS

[SN95] Jeffrey Soto and Ravindra Nemlekar. Configurable subscriptionbased scripts for system monitoring. Sys Admin: The Journal for UNIX Systems Administrators, 4(6):12–??, November/December 1995. CODEN SYADE7. ISSN 1061-2688.

Sohan:2001:WHM

[Soh01] Ripduman Sohan. Web hosting:

A migrational case study. Sys Admin: The Journal for UNIX Systems Administrators, 10(3): 53–56, March 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Solin:2002:DMQ

[Sol02] Daniel Solin. Developing a MySQL quota daemon. Sys Admin: The Journal for UNIX Systems Administrators, 11(2): 42, 44–47, 49, February 2002. CODEN SYADE7. ISSN 1061-2688.

Standboge:2002:ENO

[Sta02] James Standboge. Encrypted NFS with OpenSSH and Linux. Sys Admin: The Journal for UNIX Systems Administrators, 11(3):30, 32–34, March 2002. CODEN SYADE7. ISSN 1061-2688.

Stockton:1992:MSS

[Sto92] Mark Stockton. More simple scripts. Sys Admin: The Journal for UNIX Systems Administrators, 1(2):??, July/August 1992. CODEN SYADE7. ISSN 1061-2688.

Stone:1995:UTI

[Sto95] Donald C. Stone. UNIX, TCP/ IP, and Macs. Sys Admin: The Journal for UNIX Systems Administrators, 4(3):63–??, May/ June 1995. CODEN SYADE7. ISSN 1061-2688.

Stone:2001:SBN

[Sto01] Brandon Stone. Samba-based network recycle bin. Sys Admin: The Journal for UNIX Systems Administrators, 10(7): 37–39, July 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Strubinger:2002:HBS

[Str02] Ray Strubinger. A homegrown backup solution utilizing RSA keys, SSH, and tar. Sys Admin: The Journal for UNIX Systems Administrators, 11(4): 37–38, April 2002. CODEN SYADE7. ISSN 1061-2688.

Sultzman:2000:SAN

[Sul00] David Sultzman. Storage area networks. Sys Admin: The Journal for UNIX Systems Administrators, 9(8):33, 35, 36–38, 40, August 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Sustic:2000:URF

[Sus00] Matej Sustic. The use of routers in firewall setup. Sys Admin: The Journal for UNIX Systems Administrators, 9(5):29–30, 32, 35–36, 38, 40–41, May 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Swenson:1999:CSV

[Swe99] Timothy Swenson. Customer support — the view from the other end of the phone. Sys Admin: The Journal for UNIX Systems Administrators, 8(5): 53–56, May 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Sweet:2000:PSM

[Swe00a] David Sweet. A perl of a site map. Sys Admin: The Journal for UNIX Systems Administrators, 9(11):8, 10, 12, 14, 16, 19, 22, November 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Swenson:2000:HRS

[Swe00b] Timothy Swenson. How to read a SCSI bus trace. Sys Admin: The Journal for UNIX Systems Administrators, 9(8):43–44, 46– 47, August 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Swindler:1998:TDT

[Swi98] Dale Swindler. Tape drive technology comparison: Sony AIT, Exabyte Mammoth, and Quantum DLT. Sys Admin: The Journal for UNIX Systems Administrators, 7(2):20, 22, 24, 26, 29–30, February 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Tanaka:1998:AIS

[Tan98a] Brian Tanaka. Acquiring and installing a server certificate. Sys Admin: The Journal for UNIX Systems Administrators, 7(4):55–58, April 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Tanaka:1998:WMP

[Tan98b] Brian Tanaka. Writing maintainable Perl programs and shell scripts. Sys Admin: The Journal for UNIX Systems Administrators, 7(12):49–50, 52, December 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Tarvainen:1996:SAL

[Tar96] Scott Tarvainen. A system administration log program using Tcl/Tk. Sys Admin: The Journal for UNIX Systems Administrators, 5(8):??, August 1996. CODEN SYADE7. ISSN 1061-2688.

Taylor:1999:PD

[Tay99] Anthony Taylor. Perl for the DBA. Sys Admin: The Journal for UNIX Systems Administrators, 8(7):20, 23–24, 26– 27, 29–30, July 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Taylor:2001:SWM

[Tay01] Anthony Taylor. Simplifying WindowMaker menus in a distributed environment. Sys Admin: The Journal for UNIX Systems Administrators, 10(7): 31–34, July 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Teer:2000:NAM

[Tee00] Rich Teer. New approaches to making Solaris more secure. Sys Admin: The Journal for UNIX Systems Administrators, 9(11S): 8, 10, 12, 14, 16, 18, November 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Tessier:1996:SSW

[Tes96] Tom Tessier. Shell scripting and the WWW: Creating a searchable Web database. Sys Admin: The Journal for UNIX Systems Administrators, 5(4):??, April 1996. CODEN SYADE7. ISSN 1061-2688.

Thomassen:1995:TTD

[Tho95] Hendrik-Jan Thomassen. A

termcap/terminfo description

for xterm. Sys Admin: The Journal for UNIX Systems Administrators, 4(6):44– ??, November/December 1995. CODEN SYADE7. ISSN 1061-2688.

Thomas:1998:DRF

[Tho98a] Robert Owen Thomas. The downside of routers as firewalls. Sys Admin: The Journal for UNIX Systems Administrators, 7(6):29–31, June 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Thomas:1998:HGS

[Tho98b] Robert Owen Thomas. Haunted by the ghost of smurfing. Sys Admin: The Journal for UNIX Systems Administrators, 7(9): 63–64, September 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Thomas:1998:VPN

[Tho98c] Robert Owen Thomas. Virtual private networks: Help or hype? Sys Admin: The Journal for UNIX Systems Administrators, 7(10):67–68, October 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Thomas:1999:HSM

[Tho99] Jeffrey R. Thomas. Hierarchical storage management. Sys Admin: The Journal for UNIX Systems Administrators, 8(8): 45–48, August 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Tiso:1998:MTI

[Tis98] John Tiso. Managing TCP/ IP with Windows NT 4.0. Sys Admin: The Journal for UNIX Systems Administrators, 7(5): 45–49, May 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Tiso:2000:ASS

[Tis00] John Tiso. Automated security scanning. Sys Admin: The Journal for UNIX Systems Administrators, 9(10):73–74, 76, 78, October 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Tiso:2001:SYC

[Tis01] John Tiso. Securing your Cisco router. Sys Admin: The Journal for UNIX Systems Administrators, 10(7):41–42, 44–45, July 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Totsch:2000:MSS

[Tot00] David Totsch. Managing SUID/ SGID files. Sys Admin: The Journal for UNIX Systems Administrators, 9(9):60, 62–63, 65, September 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Tuszynski:1995:AOS

[Tus95] Jack Tuszynski. Automating operating system configuration. Sys Admin: The Journal for UNIX Systems Administrators, 4(4):78–??, July/August 1995. CODEN SYADE7. ISSN 1061-2688.

Vadala:1998:LSR

[Vad98]  Derek Vadala.  Linux soft-

ware RAID — high performance through low-cost alternatives. Sys Admin: The Journal for UNIX Systems Administrators, 7(6):19–20, 22, 24, 26, June 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

vanLicht:1997:PRA

[van97a] Ernst van Licht. Product review: Artecon’s LynxStak and LynxArray RAID systems. Sys Admin: The Journal for UNIX Systems Administrators, 6(10): 41, 43–45, 47, 48, October 1997. CODEN SYADE7. ISSN 1061-2688.

vanLicht:1997:PRR

[van97b] Ernst van Licht. Product review: RoboMon UNIX v6.1. Sys Admin: The Journal for UNIX Systems Administrators, 6(5):55–??, May 1997. CODEN SYADE7. ISSN 1061-2688.

Weerd:2000:BJS

[vdW00] Peter van der Weerd. Building a Jumpstart server for Solaris. Sys Admin: The Journal for UNIX Systems Administrators, 9(5):8, 10, 12, 14, May 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Velleca:1994:MVS

[Vel94a] Packey Velleca. A method for verifying system integrity. Sys Admin: The Journal for UNIX Systems Administrators, 3(5):89–??, September/October 1994. CODEN SYADE7. ISSN 1061-2688.

Velleca:1994:RBF

[Vel94b] Packey Velleca. Reproducing boot-disk filesystems. Sys Admin: The Journal for UNIX Systems Administrators, 3(1): 71–??, January/February 1994. CODEN SYADE7. ISSN 1061-2688.

Velleca:1994:SRS

[Vel94c] Packey Velleca. A shell-based resource scheduler. Sys Admin: The Journal for UNIX Systems Administrators, 3(6):77– ??, November/December 1994. CODEN SYADE7. ISSN 1061-2688.

Velleca:1995:CTL

[Vel95a] Packey P. Velleca. Comparing technologies for long-term, highcapacity archives. Sys Admin: The Journal for UNIX Systems Administrators, 4(6):103– ??, November/December 1995. CODEN SYADE7. ISSN 1061-

Velleca:1995:MTC

[Vel95b] Packey P. Velleca. Making tapeto-tape copies. Sys Admin: The Journal for UNIX Systems Administrators, 4(3):51–??, May/ June 1995. CODEN SYADE7. ISSN 1061-2688.

Velleca:1998:RAS

[Vel98a] Packey Velleca. Reducing asynchronous serial interrupts. Sys Admin: The Journal for UNIX Systems Administrators, 7(12): 65–68, December 1998. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Velleca:1998:TS

[Vel98b] Packey Velleca. Time synchronization. Sys Admin: The Journal for UNIX Systems Administrators, 7(10):45–46, 48–53, October 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Vincent:2000:RBA

[Vin00] Thomas Vincent. Role-based access control. Sys Admin: The Journal for UNIX Systems Administrators, 9(12):35– 36, 38, December 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Voglmaier:1999:AA

[Vog99] Reinhard Erich Voglmaier. Admin agent. Sys Admin: The Journal for UNIX Systems Administrators, 8(11):39– 43, November 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Voglmaier:2000:SWP

[Vog00a] Reinhard Voglmaier. Simplifying Web production. Sys Admin: The Journal for UNIX Systems Administrators, 9(5):73–76, May 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Voglmaier:2000:LMS

[Vog00b] Reinhard Erich Voglmaier. Link management system. Sys Admin: The Journal for UNIX Systems Administrators, 9(6): 57–59, June 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Voglmaier:2000:MLD

[Vog00c] Reinhard Erich Voglmaier. Managing an LDAP database with Perl and Apache. Sys Admin: The Journal for UNIX Systems Administrators, 9(12): 55–56, 58–59, December 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Voglmaier:2001:AS

[Vog01] Reinhard Erich Voglmaier. The art of spidering. Sys Admin: The Journal for UNIX Systems Administrators, 10(6):37–39, June 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Voglmaier:2002:WPP

[Vog02] Reinhard Voglmaier. Web publishing with Perl objects. Sys Admin: The Journal for UNIX Systems Administrators, 11(2): 8, 10, 12, 14–15, February 2002. CODEN SYADE7. ISSN 1061-2688.

Valian:2000:NAD

[VW00] Peter Valian and Todd K. Watson. NetReg: An automated DHCP registration system. Sys Admin: The Journal for UNIX Systems Administrators, 9(12): 26, 28, 30–32, December 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Wyatt:1999:NM

[WA99] Shannon Wyatt and Bryan C. Andregg. New messages. Sys Admin: The Journal for UNIX Systems Administrators, 8(9): 87, September 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Wade:1996:RMA

[Wad96] Greg A. Wade. Restricting machine access with the rma ‘shell’. Sys Admin: The Journal for UNIX Systems Administrators, 5(9):55–??, September 1996. CODEN SYADE7. ISSN 1061-2688.

Wakefield:2002:MPB

[Wak02]  Robin Wakefield.  Multi-

platform backups. Sys Admin: The Journal for UNIX Systems Administrators, 11(4):18, 20–21, April 2002. CODEN SYADE7. ISSN 1061-2688.

Ward:1992:PFa

[War92a] Robert Ward. Publisher’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 1(1):2, May/June 1992. CODEN SYADE7. ISSN 1061-2688.

Ward:1992:PFb

[War92b] Robert Ward. Publisher’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 1(2):??, July/August 1992. CODEN SYADE7. ISSN 1061-2688.

Ward:1992:PFc

[War92c] Robert Ward. Publisher’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 1(3):2, September/October 1992. CODEN SYADE7. ISSN 1061-2688.

Ward:1992:PFd

[War92d] Robert Ward. Publisher’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 1(4):3, November/December 1992. CODEN SYADE7. ISSN 1061-2688.

Ward:1993:PFa

[War93a] Robert Ward. Publisher’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 2(1):2, January/February 1993. CODEN SYADE7. ISSN 1061-2688.

Ward:1993:PFb

[War93b] Robert Ward. Publisher’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 2(2):2, March/April 1993. CODEN SYADE7. ISSN 1061-2688.

Ward:1993:PFc

[War93c] Robert Ward. Publisher’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 2(3):2, May/June 1993. CODEN SYADE7. ISSN 1061-2688.

Ward:1993:PFd

[War93d] Robert Ward. Publisher’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 2(4):2, July/August 1993. CODEN SYADE7. ISSN 1061-2688.

Ward:1993:PFe

[War93e] Robert Ward. Publisher’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 2(5):4, September/October 1993. CODEN SYADE7. ISSN 1061-2688.

Ward:1993:PFf

[War93f] Robert Ward. Publisher’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 2(6):4, November/December 1993. CODEN SYADE7. ISSN 1061-2688.

Ward:1994:EFa

[War94a] Robert Ward. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 3(4):4, July/August 1994. CO-DEN SYADE7. ISSN 1061-2688.

Ward:1994:EFb

[War94b]  Robert Ward.  Editor’s fo-

rum. Sys Admin: The Journal for UNIX Systems Administrators, 3(6):6, November/December 1994. CODEN SYADE7. ISSN 1061-2688.

Ward:1994:PFa

[War94c] Robert Ward. Publisher’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 3(1):4, January/February 1994. CODEN SYADE7. ISSN 1061-2688.

Ward:1994:PFb

[War94d] Robert Ward. Publisher’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 3(2):4, March/April 1994. CODEN SYADE7. ISSN 1061-2688.

Ward:1994:PFc

[War94e] Robert Ward. Publisher’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 3(3):4, May/June 1994. CODEN SYADE7. ISSN 1061-2688.

Ward:1994:PFd

[War94f] Robert Ward. Publisher’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 3(5):4, September/October 1994. CODEN SYADE7. ISSN 1061-2688.

Ward:1995:EFa

[War95a] Robert Ward. Editor’s forum.

Sys Admin: The Journal for UNIX Systems Administrators, 4(2):6, March/April 1995. CO-DEN SYADE7. ISSN 1061-2688.

Ward:1995:EFb

[War95b] Robert Ward. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 4(4):6, July/August 1995. CO-DEN SYADE7. ISSN 1061-2688.

Ward:1995:EFc

[War95c]  Robert Ward.  Editor’s fo-

rum. Sys Admin: The Journal for UNIX Systems Administrators, 4(6):6, November/December 1995. CODEN SYADE7. ISSN 1061-2688.

Ward:1996:EFa

[War96a] Robert Ward. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 5(1):6, January 1996. CODEN SYADE7. ISSN 1061-2688.

Ward:1996:EFb

[War96b] Robert Ward. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 5(2):6, February 1996. CODEN SYADE7. ISSN 1061-2688.

Ward:1996:EFc

[War96c] Robert Ward. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 5(5):6, May 1996. CODEN SYADE7. ISSN 1061-2688.

Ward:1996:EFd

[War96d] Robert Ward. Editor’s forum. Sys Admin: The Journal for UNIX Systems Administrators, 5(6):6, June 1996. CODEN SYADE7. ISSN 1061-2688.

Ward:1996:EFe

[War96e] Robert Ward. Editor’s forum.

Sys Admin: The Journal for UNIX Systems Administrators, 5(7):7, July 1996. CODEN SYADE7. ISSN 1061-2688.

Ward:1996:EFf

[War96f] Robert Ward. Editor’s forum.

Sys Admin: The Journal for UNIX Systems Administrators, 5(9):6, September 1996. CO-DEN SYADE7. ISSN 1061-2688.

Wartell:1999:CS

[War99]  David Wartell.  Controlling

spam. Sys Admin: The Journal for UNIX Systems Administrators, 8(11):8, 10, 12, November 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Watson:2001:CCS

[Wat01a] Michael Watson. Creating a cold standby E250. Sys Admin: The Journal for UNIX Systems Administrators, 10(9): 34, 36–37, September 2001. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Watson:2001:RRF

[Wat01b] Michael Watson. Replacing

rdist and ftp with scp and

associated utilities. Sys Admin: The Journal for UNIX Systems Administrators, 10(2): 52, 54, February 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Walker:2001:SCS

[WB01] Charles Walker and Larry Bennett. Safer CGI scripting. Sys Admin: The Journal for UNIX Systems Administrators, 10(2): 8, 10, 12, 14, 16, 18, 20, 22, February 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Wynn:1997:CTS

[WC97] Bruce Alan Wynn and Michael Carpenter. Cryptography tools for the systems administrator. Sys Admin: The Journal for UNIX Systems Administrators, 6(6):45, 47–49, 51, June 1997. CODEN SYADE7. ISSN 1061-2688.

Weeks:1999:WUC

[Wee99] John Weeks. Webifying UNIX commands. Sys Admin: The Journal for UNIX Systems Administrators, 8(10):49–50, 52– 54, 56–57, 59–60, October 1999. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Weinstein:1992:JU

[Wei92a] Sydney Weinstein. Joining USE-NET. Sys Admin: The Journal for UNIX Systems Administrators, 1(2):??, July/August 1992. CODEN SYADE7. ISSN 1061-2688.

Weinstein:1992:UEC

[Wei92b] Sydney Weinstein. USENET ELM: A case study in portability between UNIX systems. Sys Admin: The Journal for UNIX Systems Administrators, 1(1): 32–??, May/June 1992. CODEN SYADE7. ISSN 1061-2688.

Weinstein:1993:ZP

[Wei93] Sydney S. Weinstein. Zombie processes. Sys Admin: The Journal for UNIX Systems Administrators, 2(1):49–??, January/February 1993. CODEN SYADE7. ISSN 1061-2688.

Weinert:1995:PSX

[Wei95] Brad Weinert. PCs and serial X. Sys Admin: The Journal for UNIX Systems Administrators, 4(3):111–??, May/June 1995. CODEN SYADE7. ISSN 1061-2688.

Weinberg:1998:CLS

[Wei98] Joshua Weinberg. Centralized logging with syslog. Sys Admin: The Journal for UNIX Systems Administrators, 7(10): 27, 30–34, October 1998. CO-DEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Weidman:2001:SSS

[Wei01a] Randy Weidman. Simple security scripts from simple commands. Sys Admin: The Journal for UNIX Systems Administrators, 10(8):31–32, August 2001. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Weinzettel:2001:ILN

[Wei01b] A. J. Weinzettel. Interoperating Linux with NetWare. Sys Admin: The Journal for UNIX Systems Administrators, 10(7):

24, 26, 28, 30, July 2001. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.

Welter:1998:WMW

[Wel98] Pete Welter. Why is my Web site down? Sys Admin: The Journal for UNIX Systems Administrators, 7(4):18, 20, 22, 24–25, April 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Wenchel:2001:IIS

[Wen01] Kevin Wenchel. Implementing IPsec in the Solaris 8 environment. Sys Admin: The Journal for UNIX Systems Administrators, 10(10):24, 26–29, October 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Westphal:2000:SLI

[Wes00a] Kristy Westphal. Snort — A look inside an intrusion detection system. Sys Admin: The Journal for UNIX Systems Administrators, 9(9):46, 48, 50, 52– 53, September 2000. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Westphal:2000:T

[Wes00b] Kristy Westphal. Tiger. Sys Admin: The Journal for UNIX Systems Administrators, 9(6): 61–62, 64, 66–67, June 2000. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Westphal:2001:TBS

[Wes01] Kristy Westphal. Tools built to sort through Snort. Sys Admin: The Journal for UNIX Systems Administrators, 10(6):62, 64–67, June 2001. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Wetsch:1995:MS

[Wet95] John Wetsch. Multilayered security. Sys Admin: The Journal for UNIX Systems Administrators, 4(1):??, January/February 1995. CODEN SYADE7. ISSN 1061-2688.

Widdowson:2001:TAU

[WF01] Liam Widdowson and John Ferlito. Tales from the abyss: UNIX file recovery. Sys Admin: The Journal for UNIX Systems Administrators, 10(11):16, 18, 20, 22, November 2001. CODEN SYADE7. ISSN 1061-2688.

Ward:1993:FLF

[WG93] Robert Ward and Scott Graham. A form letter facility for E-mail. Sys Admin: The Journal for UNIX Systems Administrators, 2(6):59–??, November/December 1993. CODEN SYADE7. ISSN 1061-2688.

Whalin:1998:VPN

[Wha98]  Greg Whalin.  Virtual pri-

vate networks. Sys Admin: The Journal for UNIX Systems Administrators, 7(9):21–22, 25– 26, September 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

White:1996:AMP

[Whi96] Scott White. The administrator’s menu program. Sys Admin: The Journal for UNIX Systems Administrators, 5(3):??, March 1996. CODEN SYADE7. ISSN 1061-2688.

Widdowson:2000:DMF

[Wid00] Liam Widdowson. Deploying Microsoft FrontPage 2000 UNIX Web services. Sys Admin: The Journal for UNIX Systems Administrators, 9(2):16, 18, 21–22, 24, 27, 29–30, 32, February 2000. CODEN SYADE7. ISSN 1061-

                                        2688. URL http://www.samag.
                                        com/.

Widdowson:2001:JIS

[Wid01] Liam Widdowson. Jailed Internet services. Sys Admin: The Journal for UNIX Systems Administrators, 10(8):39, 41–42, 44–45, August 2001. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Willems:1993:DUR

[Wil93] Leo Willems. A disk usage report generator. Sys Admin: The Journal for UNIX Systems Administrators, 2(4):75–??, July/ August 1993. CODEN SYADE7. ISSN 1061-2688.

Wilcox:2001:IIN

[Wil01a] Mike Wilcox. An introduction to IBM’s Network Dispatcher. Sys Admin: The Journal for UNIX Systems Administrators, 10(6s): 38, 40–41, June 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Wilson:2001:UVD

[Wil01b] Brian Wilson. Using VMWare as a development tool. Sys Admin: The Journal for UNIX Systems Administrators, 10(3): 57–59, March 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Winter:1999:ABD

[Win99] George Winter. Automated boot disk fail over. Sys Admin: The Journal for UNIX Systems Administrators, 8(11):14, 16, 18, 20, 22, 24, 26, November 1999. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Withers:2000:OVS

[Wit00] Alex Withers. OpenBSD as a VPN solution. Sys Admin: The Journal for UNIX Systems Administrators, 9(9):39–40, 42, 44, 45, September 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Withers:2001:IWU

[Wit01] Alex Withers. Integrating Windows 2000 and UNIX using Kerberos. Sys Admin: The Journal for UNIX Systems Administrators, 10(12):39, 41–42, 44, December 2001. CODEN SYADE7. ISSN 1061-2688.

Welter:1997:SSJ

[WM97] Pete Welter and John Meier.

Server-side Java. Sys Admin:

The Journal for UNIX Systems Administrators, 6(8):41, 43, 45– 47, 49–51, August 1997. CO-DEN SYADE7. ISSN 1061-2688.

Wenchel:2000:ICA

[WM00] Kevin Wenchel and Stephen Michaels. Implementing C2 auditing in the Solaris environment. Sys Admin: The Journal for UNIX Systems Administrators, 9(11S):36, 38–39, 41– 44, November 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Wolf:2000:WWB

[Wol00] Gunnar Wolf. WPM — Webbased proxy manager. Sys Admin: The Journal for UNIX Systems Administrators, 9(7):??, July 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Wonchoba:1997:SSP

[Won97] Wayne Wonchoba. Spatchula — A script to patch Solaris LANs. Sys Admin: The Journal for UNIX Systems Administrators, 6(4):47–??, April 1997. CODEN SYADE7. ISSN 1061-2688.

Wong:2002:HWC

[Won02] Stanley Wong. How Web cache proxies do and don’t save on Internet costs. Sys Admin: The Journal for UNIX Systems Administrators, 11(2):34, 36– 38, February 2002. CODEN SYADE7. ISSN 1061-2688.

Woodgate:1994:US

[Woo94] John Woodgate. Using syslog.

Sys Admin: The Journal for UNIX Systems Administrators, 3(6):50–??, November/December 1994. CODEN SYADE7. ISSN 1061-2688.

Woodgate:1996:UFR

[Woo96] John Woodgate. Using fsck to repair a filesystem. Sys Admin: The Journal for UNIX Systems Administrators, 5(1):73–??, January 1996. CODEN SYADE7. ISSN 1061-2688.

Wood:2001:ANI

[Woo01] Bill Wood. AIX network install manager. Sys Admin: The Journal for UNIX Systems Administrators, 10(6s):29–31, June 2001. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Weeks:2000:BPC

[WT00] Robert Weeks and Marc Tamres.

The BigAdmin protocol: Combining tools and community. Sys Admin: The Journal for UNIX Systems Administrators, 9(11S): 45–47, November 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Wynkoop:1994:NA

[Wyn94] Brett Wynkoop. The netgroup advantage. Sys Admin: The Journal for UNIX Systems Administrators, 3(3):50–??, May/ June 1994. CODEN SYADE7. ISSN 1061-2688.

Wynn:1997:ABS

[Wyn97] Bruce Alan Wynn. Automating basic system activity monitoring. Sys Admin: The Journal for UNIX Systems Administrators, 6(2):31–??, February 1997. CODEN SYADE7. ISSN 1061-2688.

Zhang:1998:CCR

[Zha98] Randy Zhang. Configuring Cisco routers and switches for management. Sys Admin: The Journal for UNIX Systems Administrators, 7(9):8, 10, 12, 14, 16, 18– 19, September 1998. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Zhang:1999:DTI

[Zha99a] Randy Zhang. DSL: Technology and implementation. Sys Admin: The Journal for UNIX Systems Administrators, 8(10): 8, 10, 12, 14, 16, 18, 20–21, October 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Zhang:1999:ILS

[Zha99b] Randy Zhang. Implementing LAN switching. Sys Admin: The Journal for UNIX Systems Administrators, 8(6):8, 10, 12–14, 16, 18–20, June 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Zhang:2000:MDN

[Zha00] Randy Zhang. MPLS: Delivering next generation networking services. Sys Admin: The Journal for UNIX Systems Administrators, 9(12):8, 10, 12, 14, 16, 18, December 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Zinkann:1992:BURa

[Zin92a] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 1(2):??, July/ August 1992. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1992:BURb

[Zin92b] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 1(3):53– ??, September/October 1992. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1992:BURc

[Zin92c] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 1(4):14– ??, November/December 1992. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1993:BURa

[Zin93a] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 2(1):29–??, January/February 1993. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1993:BURb

[Zin93b] Elizabeth Zinkann. Books: A user’s report. Sys Admin:

The Journal for UNIX Systems Administrators, 2(2):57– ??, March/April 1993. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1993:BURc

[Zin93c] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 2(3):55–??, May/ June 1993. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1993:BURd

[Zin93d] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 2(4):87–??, July/ August 1993. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1993:BURe

[Zin93e] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 2(5):75– ??, September/October 1993. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1993:BURf

[Zin93f] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 2(6):97– ??, November/December 1993. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1994:BURa

[Zin94a] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 3(1):86–??, January/February 1994. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1994:BURb

[Zin94b] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 3(2):85– ??, March/April 1994. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1994:BURc

[Zin94c] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 3(3):73–??, May/ June 1994. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1994:BURd

[Zin94d] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 3(4):102–??, July/ August 1994. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1994:BURe

[Zin94e] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 3(5):109– ??, September/October 1994. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1994:BURf

[Zin94f] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 3(6):99– ??, November/December 1994. CODEN SYADE7. ISSN 1061-

Zinkann:1995:BURa

[Zin95a] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 4(1):??, January/February 1995. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1995:BURb

[Zin95b] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 4(2):115– ??, March/April 1995. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1995:BURc

[Zin95c] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 4(3):101–??, May/ June 1995. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1995:BURd

[Zin95d] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 4(4):125–??, July/ August 1995. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1995:BURe

[Zin95e] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 4(5):99– ??, September/October 1995. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1995:BURf

[Zin95f] Elizabeth Zinkann. Books: A user’s report. Sys Admin:

The Journal for UNIX Systems Administrators, 4(6):113– ??, November/December 1995. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1996:BURa

[Zin96a] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 5(1):79–??, January 1996. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1996:BURb

[Zin96b] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 5(2):87–??, February 1996. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1996:BURc

[Zin96c] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 5(3):??, March 1996. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1996:BURd

[Zin96d] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 5(4):??, April 1996. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1996:BURe

[Zin96e] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 5(5):75–??, May 1996. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1996:BURf

[Zin96f] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 5(6):79–??, June 1996. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1996:BURg

[Zin96g] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 5(7):67–??, July 1996. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1996:BURh

[Zin96h] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 5(8):??, August 1996. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1996:BURi

[Zin96i] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 5(9):79– ??, September 1996. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1996:BURj

[Zin96j] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 5(10):85–??, October 1996. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1996:BURk

[Zin96k] Elizabeth Zinkann. Books: A user’s report. Sys Admin:

The Journal for UNIX Systems Administrators, 5(11):65– 68, November 1996. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1996:BURl

[Zin96l] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 5(12):77, 78, 80– 82, December 1996. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1997:BURa

[Zin97a] Elizabeth Zinkann. Book: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 6(1):82–??, January 1997. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1997:BURb

[Zin97b] Elizabeth Zinkann. Book: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 6(2):67–??, February 1997. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1997:BURc

[Zin97c] Elizabeth Zinkann. Book: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 6(3):79–??, March 1997. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1997:BURd

[Zin97d] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 6(4):59–??, April 1997. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1997:BURe

[Zin97e] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 6(5):77–??, May 1997. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1997:BURf

[Zin97f] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 6(6):73–75, June 1997. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1997:BURh

[Zin97g] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 6(8):61–63, August 1997. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1997:BURi

[Zin97h] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 6(10):73–75, October 1997. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1997:BURj

[Zin97i] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 6(11):69– 70, November 1997. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1997:BURk

[Zin97j] Elizabeth Zinkann. Books: A user’s report. Sys Admin:

The Journal for UNIX Systems Administrators, 6(12):60, 62– 63, December 1997. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Zinkann:1997:UR

[Zin97k] Elizabeth Zinkann. A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 6(9):59–61, September 1997. CODEN SYADE7. ISSN 1061-2688.

Zinkann:1998:BURa

[Zin98a] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 7(1):63–65, January 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Zinkann:1998:BURb

[Zin98b] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 7(2):57, 59–60, February 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Zinkann:1998:BURc

[Zin98c] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 7(3):63–65, March 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Zinkann:1998:BURd

[Zin98d] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 7(4):63–65, April 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Zinkann:1998:BURe

[Zin98e] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 7(5):61–63, May 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Zinkann:1998:BURf

[Zin98f] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 7(6):61–63, June 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Zinkann:1998:BURg

[Zin98g] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 7(7):73–74, July 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Zinkann:1998:BURh

[Zin98h] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 7(8):63–65, August 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Zinkann:1998:BURi

[Zin98i] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 7(9):67– 70, September 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Zinkann:1998:BURj

[Zin98j] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 7(10):69–71, October 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Zinkann:1998:BURk

[Zin98k] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 7(11):69– 72, November 1998. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Zinkann:1998:BURl

[Zin98l] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 7(12):69–71, December 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Zinkann:1999:BURa

[Zin99a] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 8(1):59– 60, 62, January 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Zinkann:1999:BURb

[Zin99b] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 8(2):59–63, February 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Zinkann:1999:BURc

[Zin99c] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 8(3):63–66, March 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Zinkann:1999:BURd

[Zin99d] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 8(4):57–60, April 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Zinkann:1999:BURe

[Zin99e] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 8(5):62, 64–66, May 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Zinkann:1999:BURf

[Zin99f] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 8(6):65–67, June 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Zinkann:1999:BURg

[Zin99g] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 8(7):61–63, July 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Zinkann:1999:BURh

[Zin99h] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 8(8):67–69, August 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Zinkann:1999:BURi

[Zin99i] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 8(9):75– 78, September 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Zinkann:1999:BURj

[Zin99j] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 8(10):71–74, October 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Zinkann:1999:BURk

[Zin99k] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 8(11):71, 73, 74, November 1999. CODEN SYADE7. ISSN 1061-2688. URL

                                        http://www.samag.com/.

Zinkann:1999:BURl

[Zin99l] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 8(12):79–81, December 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Zinkann:2000:BURa

[Zin00a] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 9(1):67–70, January 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Zinkann:2000:BURb

[Zin00b] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 9(2):71–74, February 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Zinkann:2000:BURc

[Zin00c] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 9(3):71–72, 74–75, March 2000. CODEN SYADE7.

ISSN 1061-2688. URL http://

www.samag.com/.

Zinkann:2000:BURd

[Zin00d] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 9(4):73–75, April 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Zinkann:2000:BURe

[Zin00e] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 9(5):85–87, May 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Zinkann:2000:BURf

[Zin00f] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 9(6):85–87, June 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Zinkann:2000:BURg

[Zin00g] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 9(7):??, July 2000. CODEN SYADE7. ISSN 1061-

2688. URL http://www.samag.
com/.

Zinkann:2000:BURh

[Zin00h] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 9(8):81–83, August 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Zinkann:2000:BURi

[Zin00i] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 9(9):73– 76, September 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Zinkann:2000:BURj

[Zin00j] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 9(10):79–82, October 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Zinkann:2000:BURk

[Zin00k] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 9(11):87– 89, November 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Zinkann:2000:BURl

[Zin00l] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 9(11S):61– 63, November 2000. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Zinkann:2000:BURm

[Zin00m] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 9(12):89–91, December 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Zinkann:2001:BURa

[Zin01a] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 10(1):102, 104– 105, January 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Zinkann:2001:BURb

[Zin01b] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 10(2):59– 60, 62, February 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Zinkann:2001:BURc

[Zin01c]  Elizabeth Zinkann.  Books:

A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 10(3): 65–67, March 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Zinkann:2001:BURd

[Zin01d] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 10(4):65–67, April 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Zinkann:2001:BURe

[Zin01e] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 10(5):71, 73–74, May 2001. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Zinkann:2001:BURf

[Zin01f] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 10(6):73–74, June 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Zinkann:2001:BURg

[Zin01g] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 10(6s):49–51, June 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Zinkann:2001:BURh

[Zin01h] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 10(8):57–58, August 2001. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Zinkann:2001:BURi

[Zin01i] Elizabeth Zinkann. Books: A user’s report. Sys Admin:

The Journal for UNIX Systems Administrators, 10(9):51– 52, September 2001. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Zinkann:2001:BURj

[Zin01j] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 10(10):55–57, October 2001. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Zinkann:2001:BURk

[Zin01k] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 10(11):51– 52, November 2001. CODEN SYADE7. ISSN 1061-2688.

Zinkann:2001:BURl

[Zin01l] Elizabeth Zinkann. Books: A user’s report. Sys Admin: The Journal for UNIX Systems Administrators, 10(12):??, December 2001. CODEN SYADE7. ISSN 1061-2688. Web only.

Zolman:1992:BJL

[Zol92a] Leor Zolman. A background job launcher. Sys Admin: The Journal for UNIX Systems Administrators, 1(2):??, July/August 1992. CODEN SYADE7. ISSN 1061-2688.

Zolman:1992:COJ

[Zol92b] Leor Zolman. A communitystyle overnight job spooler. Sys Admin: The Journal for UNIX Systems Administrators, 1(1): 17–??, May/June 1992. CODEN SYADE7. ISSN 1061-2688.

Zolman:1992:IIB

[Zol92c] Leor Zolman. In-line input or bust. Sys Admin: The Journal for UNIX Systems Administrators, 1(4):47–??, November/December 1992. CODEN SYADE7. ISSN 1061-2688.

Zolman:1992:SLM

[Zol92d] Leor Zolman. A system load monitoring trilogy. Sys Admin: The Journal for UNIX Systems Administrators, 1(3):40– ??, September/October 1992. CODEN SYADE7. ISSN 1061-2688.

Zolman:1993:BSJ

[Zol93a] Leor Zolman. Building a secure journal/logging utility with encryption. Sys Admin: The Journal for UNIX Systems Administrators, 2(6):75–??, November/December 1993. CODEN SYADE7. ISSN 1061-2688.

Zolman:1993:LE

[Zol93b] Leor Zolman. An lp enhancement. Sys Admin: The Journal for UNIX Systems Administrators, 2(2):5–??, March/April 1993. CODEN SYADE7. ISSN 1061-2688.

Zolman:1993:PDM

[Zol93c] Leor Zolman. A project directory management facility. Sys Admin: The Journal for UNIX Systems Administrators, 2(1): 19–??, January/February 1993.

CODEN SYADE7. ISSN 1061-2688.

Zolman:1993:URC

[Zol93d] Leor Zolman. User report: CTAR and company. Sys Admin: The Journal for UNIX Systems Administrators, 2(4):53–??, July/ August 1993. CODEN SYADE7. ISSN 1061-2688.

Zwieback:1998:WCP

[Zwi98] Dave Zwieback. Web capacity planning: How to plan for server growth. Sys Admin: The Journal for UNIX Systems Administrators, 7(4):35–36, 38, 40–41, April 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Zwieback:1999:USP

[Zwi99a] Dave Zwieback. Using Solaris packages. Sys Admin: The Journal for UNIX Systems Administrators, 8(12):29–30, 32, 34, 36, 38, 40, December 1999. CODEN SYADE7. ISSN 1061-2688. URL

http://www.samag.com/.

Zwieback:1999:HH

[Zwi99b] Dave D. Zwieback. Hardening a host. Sys Admin: The Journal for UNIX Systems Administrators, 8(9):51–56, September 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Zwieback:1999:SCP

[Zwi99c] Dave D. Zwieback. Storage capacity planning. Sys Admin: The Journal for UNIX Systems Administrators, 8(8):19–20, 22, 24–26, 28, 30, August 1999. CO-DEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/.